A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.02

Title word cross-reference

2 × 2 [614, 412]. D^{2} [1105]. f(z) = ∫_{0}^{z} dx [795]. ln Γ(Z) [235]. p [1105]. π [159]. \sqrt{2} [954, 1165]. T [284]. w(z) = e^{−z^{2}}(1 + \frac{2i}{\sqrt{π}} ∫_{0}^{z} e^{t^{2}} dt) [236].

w(z) = e^{−z^{2}} ∫_{0}^{z} e^{x^{2}} dx [795].

-Statistics [1105, 1105]. -Year [284].

1 [226, 1251]. 131 [249, 428]. 1800 [984]. 1808 [306]. 1860 [42]. 1861 [397]. 1880 [111]. 1900 [239, 728]. 1914 [304]. 1935 [676]. 1938 [1049]. 1939 [1114, 1176, 730]. 1949 [876]. 1953 [840, 148, 510]. 1954 [110, 770]. 1955 [880]. 1956 [109]. 1957 [49, 485, 142, 238]. 1958 [468, 97, 388, 72].

1959 [342, 708, 1076, 144]. 1960

[609, 540, 368, 137, 772, 576, 879, 393, 391]. 1961 [1337, 435, 1269]. 1962 [401, 803, 601, 1081, 646, 681, 515]. 1963 [1050, 682, 987, 734]. 1964 [942, 847]. 1965 [1121, 1120]. 1968 [1333]. 1975 [864].

2 [830, 640, 1328]. 25 [305].

30th [847]. 39 [883]. 3ème [739].

55 [731]. 56 [106, 396]. 57 [477]. 58

[113, 581].

60 [670]. 61 [619].

A-Level [1197]. Ability [1061, 1131]. Abraham [583]. Abramov [235].

Abramovitz [787, 878]. Abramson [636]. Abridged [576]. Abstract [176, 646]. Abuse [373]. Abuses [430]. Acceptance [750]. Accessions [442, 488, 519, 553, 589, 624, 650, 687, 718, 746, 775, 812, 851, 890, 911, 947, 990, 1025, 1056, 1088, 1124, 1155, 1191, 1226, 1255, 1280, 1312, 1343]. Accident [1129, 359]. Accidents [252, 633, 1275]. Account [194]. Accountancy [1040, 683]. Accounting [835, 540, 38, 896, 475, 1044]. Accounts [71, 496, 476, 477, 505, 977, 20, 644, 937]. Accumulating [1263]. Accuracy [837]. Acher [726]. Acheson [874, 1109]. Achievement [395, 509]. Ackoff [265]. Actes [739]. Action [807]. Activities [219]. Activity [438, 1303, 1113]. Acton [13]. Acts [146, 147]. Actual [599]. Actuaries [208]. Actuary [146, 147]. Adam [484]. Adaptive [264]. Addendum [1224]. Additional [863]. Additions [28, 54, 80, 118, 152]. Address [121]. Adjustment [782, 852, 863, 1144]. Adler [607]. Administering [932]. Administration [480]. Administrative [138]. Adolescent [679]. Adult [1084]. Advanced [285, 1135, 683]. Advancement [139]. Ady [476, 542]. Affairs [1115, 1184, 535]. Affect [414]. Affecting [988, 940, 820]. Africa [881, 476, 509]. African [544]. after [984]. Again [1142]. Against [343]. Agarwala [1249]. Age [1249, 414, 336, 1200, 548, 1301]. Aged [886]. Ages [1164]. Aggregate [878]. Agricultural [629, 43, 704]. Agriculture [424, 300, 43, 1148]. Ahmed [460]. Aid [452]. Airy [235]. Aitken [1154, 343]. Alan [540, 539, 567, 957, 302, 646, 1276, 875]. Albert [304, 76, 1275]. Alcohol [1084]. Alder [133, 370]. Ale [11]. aléatoires [869]. Alexander [585, 1022, 327, 1154]. Alford [680]. Alfred [268, 538, 234]. Algebra [971, 377]. Algebraic [612]. Algèbre [726]. ALGOL [611]. Ali [237, 237]. Alison [467, 468]. Alive [577]. Allen [925, 993, 675, 642, 470, 786]. Allied [1009]. Allocation [776, 354, 525, 523]. Almarin [266]. Alternative [494, 760, 1037]. Alvaro [546]. Alvin [1005]. Amalendu [241]. Amalgamation [40]. American [680, 342, 109, 483, 346, 801, 243, 1305]. Amir [237]. Amtlichen [229]. Analyses [981, 505]. Analysing [820, 291]. Analysis [1274, 922, 605, 160, 826, 1139, 532, 89, 756, 1158, 501, 347, 20, 1100, 1232, 461, 1214, 1265, 565, 189, 601, 127, 250, 1170, 1238, 514, 1298, 125, 635, 61, 724, 1015, 289, 1042, 957, 475, 675, 158, 780, 940, 2, 295, 292, 901, 93, 1272, 1208, 1323, 165, 667, 882, 1212, 168, 538, 973, 163, 13]. Analytical [1173, 976]. Anderson [115, 1235]. Andic [909]. André [667, 1215]. Andrew [1324]. Andrews [128, 824]. Anglo [680]. Anglo-American [680]. Annotated [15]. Annual [1187, 87, 223, 364, 527, 657, 784, 919, 1065, 1203, 1322, 88, 224, 366, 528, 658, 785, 920, 1066, 1204, 1321, 388, 136, 742, 1120]. Answers [1067]. Anthropology [337]. Anticipations [170, 765]. Aoki [1207]. Application [289, 731, 894]. Applications [222, 902, 267, 424, 1143, 971, 788, 917, 758, 663, 467, 707, 1108, 164, 800, 736, 425, 531, 255, 504, 956, 294, 129, 1306, 1035, 263, 755, 960, 193, 699, 475]. Applied [1129, 1130, 1293, 996, 421, 203, 789, 534, 1043]. Appraisal [475, 99]. Appreciation [208]. Apprentices [631]. Approach [1079, 451, 1232, 523, 1015, 475, 1070, 481, 465, 744, 1134]. Approaches [506]. Approximate [998]. Arbroath [1220]. Area [207, 127]. Areas [883, 581, 1117, 1078]. Argentina [1030]. Argument [235, 236]. Armitage [290]. Armour [172]. Army [76]. Arnold [336, 383, 610]. Arrow [421]. Art [641]. Arthur [881, 498, 101, 1054, 743, 514]. Articles [27, 53, 79, 117, 152]. Artificer [631]. Artificial [31, 32]. Ash [1003].

Ashok [241]. Asia [669]. Asian [629]. Aspects [73, 1158, 273, 1119, 628, 557, 1333, 1042]. Aspirations [512]. Assay [702]. Assessing [1016, 1217]. Assessment [237, 738]. Assistance [339]. Associated [10, 1105]. Association [799, 1164, 543, 412, 1074, 630, 139, 1202, 678]. Assurances [387]. Asymptotic [235]. Atindra [1040]. Attachments [512]. Attainment [1061, 1131]. Attendance [643]. Attributes [428]. Audience [44]. Audiences [951]. August [97, 1248]. Austin [459]. Australia [978]. Australian [668, 728]. Authorship [813, 721]. Autocodes [661]. Automatic [293, 136, 742, 1151]. Average [998, 821, 345]. Aviation [1223]. Awaiting [112]. Awarded [493]. Axiomatics [256]. Aylmer [401, 441].

B [614, 469, 870, 942, 966, 1167, 254, 1337, 162, 308, 395, 336, 829, 1074, 293, 255, 210, 209, 1003, 831, 1295, 1266, 1078, 133, 370, 1250, 1121, 838, 1105, 545, 617, 643, 1327, 634, 988, 646, 385, 603, 332, 965, 926, 1242, 378]. Babbage [329, 221]. Bachelor [15]. Back [119, 153, 184, 215, 246, 277, 313, 351, 406, 444, 489, 520, 554, 590, 625, 651, 688, 719, 747, 777, 814, 853, 891, 912, 948, 991, 1026, 1057, 1089, 1125, 1156, 1192, 1227, 1256, 1281, 1313, 1344]. Background [31]. Backwardness [585]. Bacteriophage [1071]. Baggaley [1324]. Bailey [334, 788, 1070]. Bakewell [617]. Bakst [1220]. Balance [496, 1269, 639, 1148]. Balances [111]. Balestra [1079]. Banbury [179]. Band [802]. Bank [188]. Banker [545]. Barbara [1216, 733]. Barer [733]. Bargaining [1048]. Barlow [898]. Barna [583]. Barrie [1004]. Barrington [178]. Bartholomew [549]. Bartlett [164, 325, 1035, 794]. Barton [1277, 376, 1009]. Based [953]. Basic [1253, 609, 1007, 347, 1135, 266, 638, 727, 470, 567, 923, 1040, 665, 1077, 927, 1292, 94]. Basis [1245, 174]. Bass [1033]. Bates [803]. Bayes [322, 443]. Bayesian [1296, 828, 1237]. Beale [1240]. Beatty [576, 234]. Beaver [1087]. Beckenbach [789]. Beds [557]. Beer [14]. Beesley [862, 861]. Behavior [93, 679, 1084]. Behavioral [1139]. Behaviour [680, 893, 932, 1259, 1049, 760, 241, 1179]. Behrend [1178]. Being [359, 1032]. Belgrade [1248]. Bellman [534, 264, 1241]. Bená [1151]. Benefit [693, 410, 861]. Benefits [862, 1287]. Benes [873]. Bengal [237, 145]. Benjamin [336, 1272, 1250, 988]. Bennett [614, 342, 1054]. Berge [467, 972]. Bergstrom [1175]. Berkeley [368]. Bernard [1272, 460]. Bernau [674]. Berners [833]. Berners-Lèe [833]. Berrill [669]. Berry [1043]. Bethlem [49]. Betterment [1183]. Betting [1319]. Betty [196]. Between [1164, 1078, 656, 446, 187, 82, 1095]. Beveridge [551]. Bevolkerungs [230]. Bevolkerungs-Wirtschafts-Sozialstatistik [230]. Beyer [1302, 1008]. Bhagabat [1040]. Bharati [237, 237]. Bharucha [129]. Bharucha-Reid [129]. Bhattacharya [233]. Bhattacharyya [395]. Bibliographic [738]. Bibliography [454, 1231, 1346, 15, 609, 468, 876, 872, 106, 374]. Bids [40]. Bidwell [811]. Billingsley [899]. Bina [241]. Binary [820, 954, 1165]. Binomial [378, 637]. Biological [702, 1040, 971, 409, 704]. Biologists [756, 1147]. Biology [1205, 1219, 931, 1070, 662]. Biomedical [727]. Biometric [337]. Biometrical [97]. Biometrics [1243]. Biometrika [1101, 1010]. Biometry [536]. Biostatistics [1005]. Birger [974]. Birmingham [646]. Birthday [167]. Bivariate [95]. Black [928]. Blaug [1185]. Bliss [1205]. Block [1136]. Blyth [378, 1014, 200]. Board [395, 150, 394, 643, 385]. Boehm [942]. Bogue [1338]. Bombach [104]. Bond [239]. Bonini [1170]. Boody [306]. Book [680, 681, 1187, 968, 22, 206, 393, 1109, 1048, 683, 835, 1297, 113, 237, 270, 676, 459, 609, 614, 662, 702, 507, 241, 342, 936, 1046, 104, 242, 540, 678, 799, 343, 368, 381, 379, 469, 1036, 388, 1220, 1274, 287, 419, 663, 900, 922, 605, 645, 741, 257, 378, 467, 468, 572, 787, 789, 790, 791, 870, 876, 972, 1106, 1137, 1079, 534, 902, 963, 1073, 1207, 269, 484, 131, 502, 707, 740, 1150, 369, 464, 530, 638, 907, 1038, 864, 881, 1050, 45, 46, 99, 137, 141, 208, 296, 297, 298, 341, 547]. Book [615, 1, 71, 305, 942, 97, 329, 267, 397, 1069, 727, 888, 966, 1083, 1249, 715, 826, 901, 1012, 1108, 1300, 199, 840, 1047, 1177, 986, 1301, 164, 196, 481, 482, 497, 498, 1245, 1302, 642, 1139, 93, 19, 47, 390, 391, 432, 515, 73, 974, 1068, 268, 470, 479, 532, 535, 543, 1114, 1006, 436, 539, 668, 669, 670, 841, 883, 424, 567, 238, 923, 1005, 1167, 705, 761, 797, 800, 1040, 1216, 884, 1014, 1115, 980, 1217, 38, 205, 480, 582, 866, 1205, 239, 89, 254, 756, 905, 1271, 1272, 1337, 301]. Book [604, 1218, 49, 958, 1335, 1143, 64, 162, 736, 1034, 1209, 74, 665, 671, 40, 108, 109, 176, 304, 585, 878, 887, 732, 673, 148, 824, 1002, 274, 308, 476, 477, 48, 256, 335, 425, 738, 1007, 1208, 60, 1206, 1296, 772, 68, 195, 330, 462, 501, 537, 759, 1067, 347, 439, 803, 1112, 288, 395, 1147, 1219, 666, 982, 1051, 20, 483, 21, 429, 867, 92, 262, 438, 616, 620, 932, 465, 1072, 930, 1323, 583, 44, 376, 607, 929, 1169, 485, 231, 326, 380, 636, 868, 869, 143, 336, 102, 167, 763, 260, 1267, 144, 271]. Book [339, 431, 546, 577, 579, 846, 1019, 1149, 273, 428, 63, 66, 194, 757, 203, 435, 672, 39, 207, 75, 430, 466, 516, 461, 474, 602, 611, 829, 967, 971, 1000, 1053, 1214, 1221, 712, 307, 348, 478, 232, 1168, 1291, 142, 1041, 786, 135, 338, 975, 909, 843, 934, 1017, 1119, 578, 637, 1074, 1171, 286, 737, 798, 925, 565, 146, 147, 1210, 1186, 346, 587, 1113, 1184, 1018, 531, 608, 970, 1324, 325, 633, 679, 420, 564, 165, 293, 606, 697, 13, 197, 255, 101, 1084, 568, 706, 886, 309, 644, 17, 177, 210]. Book [1054, 1004, 581, 682, 734, 90, 166, 227, 328, 417, 601, 659, 924, 1001, 504, 566, 792, 827, 872, 873, 899, 903, 956, 1236, 1152, 434, 209, 272, 844, 333, 1011, 1334, 1003, 1107, 513, 667, 677, 709, 743, 744, 831, 832, 882, 981, 1049, 1080, 661, 735, 769, 985, 1293, 253, 472, 1170, 1213, 1238, 1295, 1118, 327, 514, 877, 928, 149, 150, 569, 1309, 418, 612, 699, 904, 1211, 767, 806, 979, 111, 306, 1172, 386, 112, 1298, 340, 576, 580, 733, 807, 41, 35, 373, 937, 258, 635, 660, 796, 1212, 1266, 711, 714, 1078]. Book [1179, 1185, 1117, 310, 437, 1181, 387, 294, 23, 377, 571, 825, 959, 962, 506, 762, 833, 834, 1016, 1111, 1329, 976, 801, 1294, 61, 225, 1166, 129, 145, 839, 1033, 211, 16, 100, 103, 228, 37, 847, 536, 880, 987, 133, 370, 931, 106, 169, 1082, 1215, 1222, 1332, 908, 753, 1104, 229, 230, 300, 427, 1304, 175, 1077, 1173, 140, 259, 708, 927, 983, 1013, 1076, 1182, 1247, 1326, 1081, 1015, 1250, 1333, 334, 136, 174, 264, 289, 664, 742, 964, 107, 110, 204, 505, 398, 138, 399, 65, 1042, 549, 72, 201, 303, 760, 765, 957]. Book [961, 1121, 200, 710, 1183, 1306, 226, 285, 460, 1330, 1141, 845, 1136, 544, 1180, 132, 503, 906, 1151, 943, 838, 43, 1009, 1105, 1099, 243, 766, 1043, 1269, 396, 475, 508, 639, 202, 1132, 302, 345, 42, 70, 168, 344, 76, 1292, 67, 647, 674, 1044, 389, 574, 575, 171, 375, 1035, 1146, 545, 675, 1145, 770, 1223, 170, 394, 542, 586, 617, 805, 837, 978, 713, 802, 842, 1243, 299, 509, 548, 643, 473, 541, 538, 1176, 935, 1239, 1327, 12, 416, 634, 1037, 1235, 261, 973, 584, 1071, 392, 788, 830, 848, 969, 1070]. Book [1103, 1273, 1142, 1008, 1102, 988, 98, 290, 1251, 804, 1135, 940, 36, 499, 500, 701, 754, 939, 1052, 1134, 618, 619, 768, 34, 374, 422, 793, 1277, 771, 337, 433, 646, 1248, 1270, 1336, 1338, 426, 463, 984, 1308, 728, 730, 938, 1138, 1148, 1188, 1305, 1325, 240, 139, 1303, 18, 94, 95, 96, 233, 234, 235, 236, 533, 511, 1244, 700, 871, 1110, 1276, 14, 15, 69, 128, 172, 173, 198, 263, 265, 266, 295, 382, 383, 384, 385, 421, 471, 573, 603, 610, 640, 641, 698, 726, 739, 755, 758, 764, 933, 1240, 1299, 1174, 292]. Book [372, 729, 960, 1275, 1045, 1175, 570, 731, 178, 62, 163, 423, 563, 1010, 1101, 879, 977, 510, 134, 291, 371, 613, 703, 704, 794, 795, 941, 836, 1237, 91, 130, 331, 332, 415, 828, 874, 875, 885, 921, 965, 1075, 1120, 1133, 1178, 1307, 105, 512, 1268, 1144, 1140, 926, 193, 458, 1241, 1242, 1328, 179, 1116, 1331, 1246, 898, 1039, 464]. Booker [487]. Books [51, 183, 214, 245, 276, 312, 350, 405]. Boot [762]. Booth [210]. Borenstein [242]. Born [359]. Borts [148]. Bose [241, 1105]. Boston [274]. Bowen [267]. Braddock [239]. Bradford [459, 259]. Branard [512]. Branch [172]. Branching [645]. Brech [586]. Bretherton [637]. Brian [20, 927]. Brighton [847]. Brinley [398]. Britain [616, 111, 938, 139, 995, 323, 560, 654, 455, 1127, 980, 273, 201, 939, 586]. British [864, 1216, 108, 1112, 20, 338, 943, 345, 42, 542, 617, 619, 139, 1093, 751, 495, 725, 85, 1287, 562, 298, 942, 732, 1141, 1148]. Brookings [1017, 710]. Brown [601, 795, 332, 540, 539, 74, 602, 795, 1307]. Brownlee [825, 225]. Bruce [1296, 243]. Bruun [1084]. Bryan [130]. Buckland [609, 738]. Buckley [1183, 887]. Budgetary [676]. Budgets [843]. Budin [480]. Building [744, 1160, 55, 1052, 730, 1120, 938]. Building-Essays [744]. Bulletin [1120]. Bulmer [907]. Bureau [271, 1017, 710, 1180, 574, 575, 701, 984, 941, 1116]. Bureaucracy [138]. Burley [1011]. Burnett [661]. Burnett-Hall [661]. Burnside [64]. Bus [633]. Bush [36]. Business [600, 1114, 665, 595, 347, 803, 260, 9, 597, 970, 1170, 1238, 111, 834, 174, 1138, 1325, 1098, 834, 582, 481, 482, 40, 347]. Butlin [728]. Buzzard [643].

C [683, 270, 1046, 343, 572, 870, 972, 707, 298, 1083, 826, 481, 268, 543, 866, 1205, 665, 887, 824, 1002, 477, 1147, 430, 461, 1168, 737, 798, 1113, 568, 1001, 504, 844, 831, 1213, 979, 23, 762, 833, 145, 103, 664, 243, 1142, 1102, 940, 433, 172, 173, 960, 62, 512, 898]. C.B.E [78]. C.B.E. [1021, 1123]. Cabell [266]. Cagan [879]. Calcul [1209, 736, 908]. Calculating [329, 998, 598]. Calculations [60, 72]. Calculus [621, 537, 903, 289]. Calcutta [113]. California [368]. Calorie [629]. Calot [1209]. Cameron [172]. Campbell [1147]. Canada [887]. Canadian [494]. Cancer [284, 894]. Capelle [741]. Capital [670, 627, 1030, 1081, 674, 675, 242, 478, 346, 587, 838, 202]. Capitalism [1181, 673]. Car [455]. Card [8]. Card-Matching [8]. Care [1220, 1053, 712, 1221]. Carla [973]. Carlo [664, 964, 823, 1130]. Carol [512]. Carr [1022, 1334]. Carr-Saunders [1022]. Carré [758]. Carried [581]. Carroll [206]. Cars [1016, 510]. Cartwright [1106]. Case [343, 1218, 909, 576, 1030, 647, 1270]. Catalytic [98]. Causal [744]. Causation [633]. Causes [1187, 1251]. Cecil [25, 810]. Census [46, 84, 85, 574, 575, 781, 1249, 301, 616, 1110]. Censuses [228, 968, 1146]. Center [569]. Central [394, 646, 1029, 1162, 309]. Centroids [1277]. Certain [83, 998, 959]. Ceylon [880, 943, 770]. Chaînes [755]. Chain [744]. Chains [331, 1237]. Chakrabarti [461]. Chakravarti [241, 1293]. Chakraverti [241]. Chamberlain [142]. Chambers [139]. Chance [359, 1032, 1202, 1135, 376, 607, 568]. Chandrasekhar [46]. Change [237, 271, 431, 139, 1244, 179, 1083, 149, 1309, 342, 511]. Changer [512]. Changes [597, 1031, 832, 85, 879, 107]. Character [763]. Characteristic [738, 699, 90]. Characteristics [694]. Charges [983]. Charles [329, 196, 932, 182, 197, 1170, 78, 213, 329, 210]. Charting [1216]. Charts [174]. Chase [1168]. Chaudhuri [270, 1105]. Chayen [580]. Chess [414]. Chester [576]. Childhood [1336]. Children [48]. China [46]. Choice [93]. Choix [667]. Chosen [273]. Christian [232, 509]. Christine [1054, 576]. Christopher [672, 70]. Christy [485]. Chung [331]. Church [394]. Churchill [573]. Churchman [173, 103]. Circumstances [339]. City [113, 646]. Civil [878]. Clare [1083]. Clark [799, 196, 481, 548]. Class [997]. Classical [256, 1036, 1218]. Classification [219, 491, 1251]. Classifying [188]. Claude [467, 972]. Clement [399]. Clerks [188]. Clinical [459, 75, 140]. Co [881, 1184, 172]. Co-operation [1184]. Co-Ordination [881]. Co-Sponsored [172]. Coal [643, 385]. Coalminer [248, 643]. Cochran [1291]. Codex [232]. Coding [636, 380]. Coefficients [83, 18]. Coen [1345]. Coghill [576]. Coinage [139]. Colclough [939]. Cole [339, 1141]. Colin [799, 378]. Collaboration [1095]. Collaborative [847]. Collar [647]. Collected [252, 1235]. Collecting [169]. Collection [60]. Collective [1048]. College [516, 847]. Colonial [176]. Colonies [1218]. Combinatorial [789, 376]. Comment [203, 220]. Commentary [19, 432]. Comments [356, 1345]. Commerce [1040, 139]. Commercial [1079]. Committee [615, 271, 516, 786, 1082, 1180, 941, 921, 1116, 1196, 1082]. Commodities [456, 269, 485, 307, 127, 307]. Common [703]. Commonwealth [1082]. Communautés [1246]. Communications [819, 202, 381]. Community [1253, 7]. Companies [995, 40]. Company [20, 1270]. Comparative [1220, 732, 10, 39, 760].

Comparée [1246]. Comparing [411]. Comparison [104, 123, 560, 599]. Comparisons [447, 604]. Competition [715]. Compilation [269, 41, 541]. Complete [1042]. Complex [1015, 235, 236, 795]. Components [336]. Composition [1337, 429, 220, 620]. Compound [1319]. Computable [539]. Computational [427]. Computations [1130, 33]. Computer [750, 1301, 930, 1214, 931, 158, 973, 661]. Computers [504, 263, 473, 794, 975, 426]. Comrie [572]. Concentration [751, 9, 597, 1288, 108]. Concepts [638, 927, 1135, 866, 1239]. Concluding [321]. Concurrent [159]. Conditions [347, 43]. Conducting [58]. Conference [678, 305, 543, 884, 271, 1017, 677, 710, 1180, 1248, 1188, 172, 198, 941, 1116, 739]. Confidence [378]. Confidential [142]. Configurations [1136]. Conflict [1063, 138, 324]. Confluent [235, 96]. Congestion [262, 1041]. Connected [84]. Connecting [873]. Connor [797, 194, 717]. Conrad [1014, 200]. Conspectus [8]. Constructing [410]. Construction [878, 918, 1024, 939, 1175, 1051]. Consultants [448]. Consumed [187]. Consumer [241, 1265, 1259, 1049, 169, 1182, 509, 510, 981, 708, 983, 760]. Consumer-Demand [1265]. Consumption [485, 430, 584, 429, 509, 771]. Contagious [1036]. Contemporary [1093]. Context [994]. Contingency [614, 190, 411, 250]. Continuing [48]. Continuous [1261, 1233]. Continuous-Time [1233]. Contribution [667, 1148]. Contributions [347, 1235, 870, 162, 330, 232, 460]. Control [450, 902, 1285, 124, 293, 386, 1028, 174, 264, 1151, 1284, 874, 1241, 1242, 1283, 835, 74, 875]. Controlled [140]. Controversy [361, 161]. Convex [792]. Conway [1072, 608]. Cootner [763]. Coppock [310]. Corporate [270, 995, 239]. Corrado [850, 497, 166, 227]. Corrected [1249]. Correction [1024, 1346]. Correlation [61]. Corrigenda [1131, 443]. Corrigendum [51, 180, 1165]. Cost [683, 238, 168]. Costs [73, 918, 1024, 708, 935]. Council [516, 87, 223, 364, 527, 657, 784, 919, 1065, 1203, 1322, 1095, 931]. Counties [455]. Countries [271, 478, 629, 509, 1048, 881, 1119, 310, 671]. Courcier [476]. Cours [1209]. Course [963, 199, 970, 255, 1132, 375, 37]. Court [48]. Coutie [766]. Cox [900, 922, 253, 45, 377, 458]. Craddock [1301]. Craig [1154, 925]. Cramer [63, 1108, 417]. Crane [103]. Crauford [978]. Creamer [242, 1270]. Credit [708, 983, 1182]. Criminals [580]. Criswell [500]. Criteria [1241]. Critical [7, 295]. Criticism [1130, 431]. Critique [676]. Croft [38]. Croft-Murray [38]. Cronbach [844]. Crop [1069]. Crops [883]. Crossley [345]. Crow [135]. Crowe [800]. Crude [656]. Crump [649]. Cruon [294]. Cukor [541]. Cullingworth [209]. Cumulative [759]. Curia [1245]. Current [559, 24, 50, 77, 114, 151, 181, 212, 244, 275, 311, 349, 400, 440, 486, 517, 550, 588, 622, 648, 684, 716, 745, 773, 808, 849, 889, 910, 944, 989, 1020, 1055, 1085, 1122, 1153, 1189, 1225, 1252, 1278, 1310, 1339, 347, 1053, 712, 1286, 859, 43, 389, 574, 575]. Curvilinear [61]. Curwen [549]. Custom [1220]. Cybernetic [503]. Cybernetics [14]. Cycle [347, 619]. Cycles [1114, 148, 439, 111, 938]. Cyclical [347, 935]. Cyrus [37]. Czechoslovakia [309].

D [1109, 1297, 614, 702, 104, 242, 469, 900, 922, 605, 741, 870, 707, 1012, 1300, 470, 543, 980, 1335, 108, 48, 68, 395, 932, 376, 143, 273, 786, 1018, 679, 197, 644, 831, 981, 1049, 661, 253, 35, 834, 1294, 145, 664, 1009, 171, 542, 978, 643, 473, 1142, 940, 1277, 1270, 1138, 1148, 1303, 235, 128, 641, 758, 795, 332, 828, 458]. D.Sc [1154]. D.Sc. [945, 78]. Dacey [204]. Dangers [392]. Daniel [1012, 965]. dans [499]. Dantzig [26, 603]. Darmois [116]. Dasgupta [1040]. Data [1263, 817, 1249, 1139, 347, 190, 953, 85, 300, 1015, 158, 1274, 1335, 430, 13, 169, 1277, 291, 1075, 1044, 170]. Dates [389]. d’Attente [294]. David [378, 376, 66, 286, 962, 1009, 1223, 1277, 933, 604, 429, 26, 463]. Davies [766, 1178]. Davis [135]. Day [683]. Deane [1141, 106, 977]. Death [10, 1251, 656, 1187]. Death-Rates [10]. Deaths [142]. Deb [241]. December [682]. Decimal [139]. Decimalisation [343]. Decimals [703]. Decision [1073, 131, 1296, 1169, 16, 1134, 371, 1237, 1001, 833, 535]. Decisions [171, 844, 1170, 1238, 834, 1327, 195]. Decomposition [257]. Decompositions [1288]. DEF [428, 249, 428]. DEF-131 [428, 249, 428]. Defects [430]. Defence [428, 249, 428, 428]. Defense [1222]. Degree [446, 1164]. Degrees [493]. Delinquent [679, 211]. Demand [456, 1079, 1265, 667, 709, 981, 557, 1016]. Demande [667]. Deming [260]. Democracies [680]. Demographic [31, 297, 1272, 144, 271]. Demography [1271, 1338, 45]. Demonstration [159, 240]. Density [795]. d’Entrées [1246]. Department [1115, 1184, 43]. Dependent [1145, 757]. Depressed [1117]. Derivatives [795]. Derived [1259]. Derman [37]. Desabie [967]. Desai [1337]. Describing [573]. Description [72]. Descriptive [608, 706, 830]. Design [124, 819, 260, 461, 1000, 1231, 1346, 798, 1136, 134, 514]. Designed [159]. Designs [1042, 491, 1136, 55]. Detection [906]. Determinants [995, 879]. Determination [677]. Deutschen [934]. Developed [335, 271]. Developing [618]. Development [73, 669, 884, 620, 932, 188, 909, 1184, 734, 1118, 1030, 1181, 987, 1077, 1015, 42, 1223, 728, 818, 986, 1115, 583, 135, 729]. Developments [834]. Deviance [807]. Deviates [233]. Deviation [573]. Devons [345, 1190]. Devote [44]. Dewhurst [310]. Dhar [241]. Dharm [883]. Diaries [1019]. Dichotomy [630]. Dick [665]. Diet [342]. Dietrich [698]. Diets [188]. Differences [298]. Different [359]. Differential [235, 902]. Differentials [110]. Diffusion [1285]. Digital [1214, 504, 473]. Dimensional [250]. Dimensions [562, 375]. Dimock [138]. Direct [710]. Directed [1106]. Direction [438]. Disagreement [822]. Disciplinary [172]. Discrete [1036, 1207, 602, 997, 1233]. Discrete-Time [1207, 1233]. Discretionary [760]. Discussion [258, 575]. Diseases [1251]. Disorders [1286]. Distributed [196]. Distribution [157, 257, 790, 1047, 60, 431, 1129, 959, 1332, 701, 281, 1127, 95, 1137, 909, 637, 1018, 909]. Distribution-Free [60]. Distributions [997, 1036, 417, 422, 234, 573]. Diversification [483]. Diversified [600]. Divided [194]. Do [449, 705]. Dobb [1181]. Dobrovolsky [242]. Documentation [1074]. Dodge [17]. Does [414]. Doig [454, 467, 468, 876]. Dollar [842]. Domain [795, 795]. Domestic [72]. Donald [1068, 304, 1338]. d’Ordonnancement [758]. Dorfman [514]. Doris [1051]. Dorothy [195, 339, 886, 647]. Dorwin [1106]. Double [17]. Douglas [1104, 618]. Downie [923]. Dresel [661]. Dreyfus [534]. Drift [1179]. Drinking [1084]. Driver [633]. Dropout [679]. Duari [395]. Dubey [395]. DuBois [796]. Duckworth [384]. Dudley [99, 1123]. Duncan [93, 99, 874]. Dunn [727]. Durable [709]. Durables [510]. Duration [457]. During [359, 581, 336, 111]. Dynamic [1079, 534, 1073, 740, 203, 101, 506, 16]. Dynamics [293, 1151]. Dynkin [332].

Eades [262]. Earl [267, 774]. Earliest [463]. Earnings [1127, 345, 1308]. East [881, 669]. Easterlin [1305]. Easy [60]. Econometric [481, 982, 882, 801, 1046, 743, 744]. Econometricians [1145]. Econometrics [831, 1198, 56, 1045, 102]. Economic [559, 51, 27, 53, 79, 117, 152, 113, 237, 1097, 678, 799, 707, 817, 298, 826, 481, 482, 782, 852, 863, 532, 543, 320, 539, 668, 1258, 1014, 1115, 1217, 38, 1143, 673, 274, 20, 438, 583, 271, 339, 39, 909, 1017, 1119, 1184, 644, 985, 514, 979, 1030, 23, 801, 880, 987, 1082, 219, 1304, 1161, 200, 710, 1141, 1180, 1096, 43, 243, 42, 170, 837, 509, 1176, 988, 940, 768, 433, 984, 562, 728, 1305, 1175, 977, 941, 965, 1116, 908, 479, 669, 980, 585, 1113, 1304, 1333, 1121, 978, 729]. Economica [867]. Economics [507, 678, 1040, 480, 582, 595, 1163, 203, 976, 43, 674, 55, 941, 1140, 267, 268, 205, 665, 1054, 1334, 735, 1222, 1326, 1103, 1138, 69, 1268, 1118, 767, 1244]. Economists [1174]. Economy [864, 725, 435, 346, 882, 1076, 398, 243, 542, 731, 507, 1216, 1112, 348, 309, 832, 1331]. Edgar [182]. Edmond [666]. Edmund [25, 810]. Edna [580]. Educated [1185]. Education [660, 1266, 654, 896, 1287, 426, 1306]. Educational [354, 1173, 1276]. Edward [1034, 478, 387, 133, 370, 1336]. Edwards [260, 1183]. Edwin [789, 135, 811]. Effect [1318]. Effectiveness [395]. Effects [819, 879]. Efficient [283]. Effort [523]. Egerton [171]. Ehrenfeld [659]. Eighteen [310]. Eighth [366, 615]. Einaudi [402]. Einführung [698, 1299, 1039]. Einzig [842]. Eisenhart [573]. Ekambaram [174]. Elaine [880]. Elasticities [953]. Elderly [392]. Elderton [403]. Election [692, 942]. Elections [681]. Electoral [1232]. Electronic [1044, 158, 473, 263]. Elementary [254, 1135, 1133, 1139, 1167, 1034, 335, 1168, 286, 1326, 969]. Elements [1159, 145, 261, 788, 707, 369, 1038, 425, 578, 420, 327, 129, 1033]. Eliasberg [243]. Elimination [127]. Elio [679]. Elizabeth [239, 346, 306]. Ellis [927]. Ely [1190, 345]. Emanuel [193]. Emily [329]. Emma [577]. Emphasis [289]. Empirical [760, 1306]. Employees [1127, 711]. Employers [1287]. Employment [670, 148, 485, 300, 838, 548, 940, 619, 984, 207]. Enders [503]. Energy [243]. Engaged [629]. Engel [953]. Engineering [929, 798, 514, 701, 197, 825, 225]. Engineers [932, 37, 379, 1006, 877]. Engines [329]. England [19, 47, 390, 391, 432, 515, 142, 209, 576, 1110, 180, 126, 357, 394, 394]. English [642, 733, 30, 281, 306]. Enquiries [142]. Enquiry [978]. Enrolments [493]. Enterprise [751]. Entries [496]. Entropy [231]. Environment [769]. Environs [237]. Epidemics [1062]. Epidemiological [388, 188]. Epidemiology [98]. Equation [952]. Equations [1145, 235, 1241]. Equilibrium [987, 824, 348]. Equipment [76]. Equitable [387]. Era [463]. Eradication [943, 966]. Ergodic [899]. Eric [637]. Erik [1143]. Erling [1135]. Ernest [78]. Ernie [160, 2]. Errata [813, 776, 852]. Error [124]. Especial [335]. Essays [532, 162, 203, 1053, 712, 744, 801, 585, 325, 831, 1140]. Essence [536]. Essentials [875]. Estes [36]. Estimated [323]. Estimates [953, 159, 1289, 730]. Estimating [83, 283, 410, 190, 629, 1015, 389]. Estimation [823, 60, 330, 284, 635, 318]. etc [141]. Étude [1246]. Eugene [66, 76]. Europe [297, 1120, 127, 842, 310]. European [1119, 492]. Européennes [1246]. Evaluation [905, 354, 1015, 1029]. Evely [108]. Evelyn [112, 340, 962]. Events [922]. Eversley [940]. Everyman [607]. Evidences [878]. Examination [273, 186, 295]. Examinations [395]. Examples [135]. Exchange [802]. Exercises [834, 1171, 286, 1067, 834].

Exotic [509]. Expansion

[884, 954, 1165, 398]. Expectations [1329]. Expenditure [323, 82, 58, 696, 1049, 72, 220, 302, 434]. Experience [1316, 987, 1077, 1305]. Experiences [274]. Experiment [560, 347, 159, 186]. Experimental [83, 737, 798, 928, 1136, 134]. Experimentation [6, 1028]. Experiments [124, 8, 1231, 1346, 58, 461, 1000, 848, 731]. Expert [615]. Explorations [344]. Exploring [836]. Exponential [894]. Exponentially [196]. Export [732]. Exposed [392]. Extensions [603]. External [1030, 300, 880]. Extremes [793]. Eye [336]. Ezekiel [61]. Ezhov [175].

F [789, 1137, 369, 530, 1038, 208, 1300, 1068, 705, 1002, 48, 1112, 20, 930, 376, 435, 207, 611, 286, 420, 644, 17, 667, 709, 962, 708, 1247, 1081, 1333, 765, 961, 1180, 838, 1009, 243, 548, 643, 538, 1277, 463, 700, 563, 704, 1116]. F. [1254]. F.B.A [1022]. F.R.S. [322, 443, 1154]. F.R.S.E [1154]. Factor [125, 635, 598, 1212, 605]. Factors [820, 629, 988, 940, 234, 438]. Factory [174, 252]. Facts [394]. Factual [270, 843]. Faddeyeva [236]. Fair [514]. Fairhead [834]. Falconer [68]. Families [187, 274, 48]. Family [82, 909, 58, 220, 1188, 429, 299]. Fano [381]. Farm [505, 1121]. Farming [43]. Farnsworth [430]. Farrell [1054]. Faut [909]. Feasibility [1117]. Federal [150, 710]. Federation [198]. Feedback [1266]. Feeding [841]. Feller [956]. Female [123, 1290]. Ferber [582, 169]. Ferguson [715, 901, 1134]. Fertility [856, 1188]. Fertilizers [1069]. Festschrift [962]. Few [273]. FG [513]. Fieller [182]. Fifth [1321, 167]. Fifties [542]. Fifty [208]. Figures [394]. Finance [676, 20, 983, 42, 675, 394, 105, 1276, 300]. Finances [693, 862, 843, 861, 303]. Financial [496, 478, 169]. Financing [1115, 239, 1081, 242, 346, 202]. Finely [194]. Finies [755]. Finite [1136, 1]. Finney [614, 702, 1294]. Firm [821]. Firms [936]. First [785, 963, 208, 970, 255, 37, 1082]. Fish [1082]. Fisher [401, 704, 532, 1000, 441, 848]. Fisk [1145]. Fix [962]. Fixed [560, 627]. Flagle [197]. Fleck [1080]. Fletcher [572]. Flieger [1274]. Florence [767]. Florian [1080]. Flow [741, 749, 475, 571]. Flow-of-Funds [475]. Flows [982]. Fluctuations [1329, 42]. Fogel [801]. Fonctions [667]. Food [320, 430, 318, 319, 218, 858]. Football [1202]. Force [1305, 1264]. Forecast [599, 586, 433]. Forecasting [481, 1258, 74, 602, 766, 1245]. Foreign [485, 360, 770, 1017, 842]. Forest [1243]. Forman [13]. Formation [242, 346, 1081, 202]. Forms [10]. Formulae [493, 1102, 1213]. Formulas [787]. Forsyth [220]. FORTRAN [1012, 1174]. Fortschritts [1080]. Forty [1082]. Forty-First [1082]. Foster [862, 861]. Foundation [799, 172, 635]. Foundations [799, 903, 423, 1242]. Four [768]. Fourth [1204, 368, 515]. Fox [1143, 609, 61]. Fraction [952]. Framework [205]. Frances [135]. Francis [485, 258]. Frank [1106, 1150, 665, 824, 41, 571, 228, 110, 538, 898]. Franklin [532, 549, 226]. Fraser [1297, 35]. Fred [798]. Freda [1072, 608]. Frederic [310]. Frederick [255, 552, 1152]. Frederik [404]. Free [60, 127]. Freeman [530]. Frequency [32]. Freund [416, 1133]. Friedman [1148]. Frisch [167, 344]. Fritz [714]. Front [4, 120, 154, 185, 217, 247, 278, 314, 353, 408, 445, 490, 522, 556, 591, 626, 653, 690, 720, 748, 779, 816, 854, 892, 914, 950, 992, 1027, 1059, 1091, 1126, 1157, 1194, 1229, 1257, 1282, 1315]. Fry [235]. Fuchs [1331]. Function [1108, 827, 505, 1009, 95, 236, 795]. Functions [787, 283, 235, 257, 790, 90, 699, 1043, 96, 95].

Fund [713]. Fundamentals [263]. Funds [749, 478, 475]. Furneaux [273]. Further [1201]. Future [320, 1094, 1128, 319, 218, 753].

G [241, 104, 799, 1220, 419, 741, 468, 791, 876, 740, 907, 1, 397, 715, 901, 1177, 481, 470, 535, 436, 1167, 162, 1209, 74, 48, 537, 262, 930, 1169, 336, 194, 602, 1221, 1291, 142, 786, 934, 1113, 706, 17, 1004, 709, 661, 985, 1293, 149, 699, 937, 258, 635, 1266, 959, 762, 1016, 1013, 964, 399, 961, 285, 460, 845, 1009, 766, 1132, 76, 1292, 67, 375, 542, 617, 473, 1176, 973, 392, 830, 969, 1102, 940, 728, 1325, 235, 739, 570, 423, 130]. G.C.E. [446]. Galatin [1244]. Galbraith [729]. Gambling [628, 463]. Games [972, 902, 1139, 1319, 16, 69, 16, 100, 463]. Ganapati [1036]. Ganeshalingam [1277]. Ganguly [241]. Gani [1071]. Gardiner [1243]. Garth [1309]. Gas [1079]. Gass [663]. Gayen [395]. Geary [707]. Geiringer [868]. General [88, 224, 222, 366, 365, 528, 529, 658, 785, 920, 1066, 1204, 1321, 692, 19, 47, 390, 391, 432, 515, 143, 516, 961, 1318]. Generalized [954, 1165, 1129]. Generation [1244]. Generations [211]. Generators [974]. Genesis [338]. Genetic [579, 334]. Genetics [97, 68]. Geneva [884]. Geoffrey [347, 1182]. Geological [501]. Geometrical [791]. Geometry [375]. George [1139, 668, 148, 465, 587, 255, 1238, 802, 603, 130, 1018]. Georges [116]. Georgescu [976]. Georgescu-Roegen [976]. Gerald [1152]. Gerlough [741]. Gerschenkron [585]. Gershenson [547]. Gesamtwirtschaftlichen [1080]. Ghana [1290]. Ghosh [1218, 731]. Ghouila [972]. Ghouila-Houri [972]. Ghurye [162]. Gibbs [335]. Gibson [340, 112]. Gilroy [966]. Gini [850, 166, 227, 497]. Giovanna [38, 434]. Giovanni [867]. Girault [736]. Gisele [1119]. Giuliano [1074]. give [159]. Given [190]. Gladys [512]. Glass [885, 177]. Glenn [234]. Glennerster [1054, 1276]. Gleser [844]. Glossary [805]. Gnedenko [254]. Goals [835]. Gods [463]. Godwin [790]. Goldberg [977, 132]. Goldberger [101, 101, 743]. Goldine [844]. Goldsmith [639, 70, 759, 508]. Goldstein [706, 830]. Gomme [1345]. Goodman [136, 742]. Goods [29, 509, 709]. Goodwin [978, 392]. Goon [1040]. Gordon [1053, 712, 514, 755]. Gort [483]. Gotkin [706, 830]. Gottlieb [730]. Government [676, 594, 1094, 1095, 843, 146, 147, 514]. Governors [150]. Graduate [1054, 1132, 1054]. Graham [617]. Granger [826]. Grants [251]. Graphical [997]. Graphs [467, 787, 1106]. Graunt [524]. Gray [1210]. Graybill [226, 563]. Great [616, 201, 939, 323, 455, 1127]. Greater [921, 1095]. Greece [581, 987]. Greek [1076, 813, 721]. Green [827]. Greenberg [460]. Greene [16]. Greenwood [498]. Gregor [848]. Gregory [269]. Grenander [63, 612]. Griffin [531]. Gros [548]. Grossman [269]. Group [187, 1099, 500, 385, 1099]. Grouped [330]. Groupings [884]. Groups [195]. Growing [841, 48]. Growth [456, 936, 540, 678, 436, 539, 670, 980, 1143, 620, 856, 1064, 336, 39, 348, 1017, 309, 769, 801, 1141, 1233, 943, 302, 770, 447, 584, 1142, 940, 768, 1270, 562, 1305, 836, 539, 938, 977, 1011]. Gruenberger [930]. Grundy [52]. Guardian [345]. Guenther [866]. Guiana [943]. Guide [1012, 498, 668, 428, 384, 738]. Guided [264]. Guillebaud [268, 979, 23]. Gunewardena [880]. Gunnar [330]. Gupta [1040]. Gurland [662]. Guthrie [1103]. Guttman [929]. Guy [973]. Gy [541].

H [1048, 342, 1220, 900, 790, 464, 530, 481, 482, 498, 1302, 797, 89, 756, 604, 49, 162, 148, 60, 759, 347, 395, 429, 262, 465, 763, 194, 637, 1113, 197, 17, 328, 1011, 981, 1080, 928, 1172, 580, 796, 1212, 1117, 1182, 1326, 399, 544, 243, 1132, 542, 1243, 1008, 1277, 337, 18, 234, 15, 128, 1299, 62, 1010, 1307]. H.C [146, 147]. H.M. [252]. Hacking [924]. Hadley [740, 1169, 1325]. Hague [543, 618]. Haight [1137, 571]. Hájek [1211]. Hal [545]. Half [753]. Hall [661, 397, 426]. Haller [1043]. Halm [802]. Hamilton [76]. Hammersley [664]. Hamming [379]. Hammond [580]. Hand [1329]. Handbook [469, 787, 1137, 1302, 1293, 1298, 1008, 422, 793]. Handel [263]. Handscomb [664]. Hanna [41, 110]. Hannan [165]. Hans [1150, 982, 243]. Harald [1108, 63, 417]. Harary [1106]. Harcourt [1113]. Hardwick [194]. Hare [49]. Harley [576]. Harman [1212]. Harold [162, 326, 514, 487, 392, 1103, 1308]. Harriet [1178]. Harris [645]. Harrod [618, 203, 713]. Harry [547, 1212, 1044]. Hart [803, 597]. Hart-Paris [597]. Harter [1277]. Hartkemeier [1044]. Hartley [498, 1010]. Harvard [504]. Harvey [471]. Haseeb [772]. Hashim [237]. Haswell [1118]. Hatanaka [826]. Hathaway [679]. Hauser [99]. Havas [541]. Hazelwood [1]. Hazlewood [881]. Health [1187, 141, 615, 1083, 48, 579, 338, 1162, 1318, 1251, 1250]. Heath [923]. Heilperin [1074]. Held [368, 97, 601, 678, 847, 107, 543]. Helen [430]. Helliwell [1247]. Hellman [76]. Henderson [980, 20, 660]. Henn [1299]. Henry [133, 370, 754]. Herbert [1078, 500, 421, 292, 512]. Herbold [973]. Herdan [194, 537, 845, 67]. Heritage [399]. Herman [1047, 872, 744, 512]. Hernández [1186]. Hickman [239]. Hicks [205]. hiérarchique [829]. Higgledy [1142]. High [1176]. Higher [654]. Hilda [868, 806]. Hilde [1178]. Hill [1173, 1303, 459, 259, 1303]. Hillier [1152]. Hilsenrath [973]. Hints [1067]. Historical [585, 176, 887]. History [137, 243, 463, 358, 801]. Hodges [638]. Hoeffding [162]. Hoel [969]. Hoffmann [934]. Hogg [925, 433]. Hoguet [481]. Holders [107, 396]. Holman [970]. Holz [196]. Home [112, 340]. Homicide [1275]. Hon [1021]. Honor [162, 167]. Hooke [606]. Hope [1323]. Horace [568]. Horie [713]. Hospital [49, 975, 549]. Hospitals [1220]. Hotelling [162]. Houlden [385]. Houri [972]. Hourly [345]. Household [709, 581, 1081]. Householder [280]. Households [180, 126, 786]. Housing [274, 209, 1146, 940, 885, 921, 1120, 855, 274, 209]. Houthakker [981]. Howard [419, 1054, 327, 754, 1276, 1178]. Hsu [614]. Hufschmidt [514]. Huggins [197]. Hugh [623, 1087]. Hughes [42]. Hugo [98]. Huitson [875, 957]. Hultgren [935]. Humaines [499]. Human [535, 722, 1219, 31, 324, 1234]. Hundred [88, 224, 366, 528, 658, 785, 920, 1066, 1204, 1321, 993]. Hunger [316, 218]. Hunter [898]. Hurst [928]. Hutchinson [378]. Hutton [617]. Hybridisation [848]. Hypergeometric [96, 235]. Hypotheses [34]. Hypothesis [961, 1037]. Hypothesis-Testing [1037].

Ian [924, 674]. Ida [76]. Ideas [567, 565, 1013, 463]. I’Entreprise [640]. II [47, 390, 143, 1135, 793, 1188]. III [19, 432, 1298, 1188]. Ijiri [835]. Ilersic [761, 12]. Illinois [172]. Illness [48]. Ilse [111]. im [1080]. Impact [883, 786, 909, 101]. Import [1148]. Imports [494, 513]. Improvement [453, 452]. Incestuous [32]. Including [971, 16]. Income [157, 799, 305, 840, 274, 772, 1265, 909, 434, 677, 106, 110, 72, 544, 770, 1332, 1047, 431, 70]. Incomes [909, 1178]. Incomplete [1136]. Increasing [283]. Index [572, 454, 800]. Indexes [345]. India [270, 676, 841, 883, 1337, 843, 344, 360, 395, 1118, 240]. Indian [1249]. Indians [177]. Indicator [509]. Indicators [347, 347]. Indirect [710]. Individual [1178, 93]. Individuals [188]. Industrial [222, 994, 269, 436, 1119, 147, 523, 1077, 838, 874, 150, 175, 584]. Industries [397, 1117, 587, 1184]. Industry [751, 484, 109, 495, 979, 1185, 23, 85, 300, 345, 42, 939, 30, 619, 915, 108, 483, 1149, 767, 617, 1052]. Inequalities [1037, 790]. Infant [1336]. Inference [315, 1060, 279, 832, 37, 285, 859, 1135, 960, 828, 1297, 530, 866, 462, 606, 924, 377, 423]. Infertility [31]. Infinitely [91]. Infinitesimal [289]. Inflation [1318, 543]. Inflationary [435]. Influence [1163, 770]. Information [3, 81, 216, 352, 407, 521, 555, 652, 689, 778, 815, 913, 949, 1058, 1090, 1193, 1228, 1314, 381, 592, 532, 380, 1107, 172, 899, 131, 636, 1003, 1329]. Inheritance [1084]. Injuries [147, 1251]. Injury [392]. Input [305, 495, 33, 494, 731, 541]. Input-Output [494, 731, 541]. Inquiry [786]. Insemination [31, 32]. Inspection [428, 249, 17]. Inspectorate [252]. Instalment [1182]. Institute [137, 208, 1245, 576, 172, 1052]. Institution [1017, 1017, 710]. Institutions [478]. Insurance [146, 147]. Integral [952, 795]. Integration [483, 492]. Integrity [691]. Intellectual [691]. Intelligence [1084]. Inter [558, 724, 505, 172]. Inter-Disciplinary [172]. Inter-Farm [505]. Inter-Regional [558, 724]. Interaction [250]. Intercorrelations [187]. Interdependence [583, 1052]. Interim [1196]. Intermediate [464]. Internal [1162]. International [22, 157, 678, 137, 97, 1245, 543, 560, 982, 872, 714, 1166, 447, 713, 1251, 198, 1144, 799, 618]. Internationale [739]. Interpolation [18]. Interpretation [1335, 20]. Intervals [378]. Interviewer [696]. Introducing [341]. Introduction [1092, 419, 741, 1106, 1073, 164, 424, 205, 254, 1206, 1296, 1072, 102, 474, 611, 956, 1236, 796, 1117, 1166, 100, 427, 226, 1035, 134, 91, 965, 828, 35, 1294, 132, 1135, 530, 141, 1271, 1002, 68, 462, 262, 1169, 737, 925, 293, 606, 659, 1152, 133, 370, 1013, 334, 1325, 533, 563, 130, 828, 1241]. Introductory [1324, 1145, 929]. Intuitive [465]. Invariance [1160]. Invention [1176]. Inventive [438]. Inventories [1007, 439]. Inventory [99, 74, 1329, 471]. Investigation [83, 194, 643]. Investigations [819, 422]. Investment [560, 803, 506, 1247, 1334, 171, 728]. Involved [629]. IPA [682]. Iraq [772]. Irving [607]. Irwin [929]. Isaacs [902]. Iskrant [1275]. Islam [882]. Isserlis [945]. Issue [167]. Issues [976, 1222]. Italian [694]. Item [292]. IV [1012]. Ivan [906].

J [683, 614, 662, 702, 241, 1220, 572, 790, 369, 638, 1038, 137, 826, 840, 1301, 482, 498, 642, 797, 205, 582, 732, 1002, 48, 335, 501, 288, 21, 380, 339, 75, 611, 967, 142, 578, 1210, 1186, 587, 970, 165, 293, 309, 644, 209, 844, 1334, 1293, 1172, 373, 711, 310, 959, 962, 762, 834, 1016, 1111, 1294, 1033, 1173, 334, 664, 964, 107, 1151, 838, 1043, 396, 202, 302, 345, 42, 168, 675, 805, 973, 1071, 788, 1070, 1142, 940, 939, 1277, 1270, 1338, 426, 938, 96, 233, 383, 421, 726, 570, 62, 703, 1237, 874, 1140]. Jack [1050, 1269, 302]. Jackpot [1320, 1032]. Jackson [940]. Jacob [1023]. Jacobs [1073, 304]. Jacques [903]. Jaffray [930]. Jahrhunderts [934]. Jamaica [207]. James [206, 803, 15, 1019]. Jan [240]. Janet [336]. Jánossy [905]. Jansson [974]. January [682]. January-December [682]. Jaroslav [1211]. Jati [1143]. Jawetz [1121]. Jay [16]. Jean [727, 1178]. Jeanne [1083]. Jeffrey [1001]. Jeffreys [326]. Jenks [346]. Jerzy [368]. Jessop [1341]. Jewkes [338, 338]. Joan [479, 576, 500, 1174, 875]. Job [512]. Johan [404]. John [1050, 1208, 465, 338, 531, 472, 1298, 310, 437, 536, 1247, 416, 422, 793, 729, 570, 885, 1133, 1140, 623, 1123, 524]. Johnson [464, 798, 65, 168]. Joliet [1275]. Jones [509, 1309]. Jordan [213]. Jöreskog [635].

Josef [936, 328]. Joseph [243, 973, 130]. Joyce [576]. Jr. [1336]. Judith [1179]. Juha [1084]. Julian [827]. Julius [347]. June [368, 884, 601]. Juster [1081, 765].

K.B.E [1022]. K.B.E. [1087]. Kahn [501, 806]. Kai [331]. Kalecki [1114]. Kalton [1013]. Kamat [959]. Kapur [62]. Karl [1143]. Karlin [421, 963, 69]. Karmel [1113]. Karpov [795]. Kaufman [484]. Kaufmann [294, 610, 640]. Kaye [178]. Keen [875]. Keeping [462]. Keilson [827]. Keith [1323]. Kellogg [482]. Kemeny [570]. Kempthorne [97]. Kendall [468, 791, 876, 285, 1009, 1345, 454, 375]. Kendall-Doig [454]. Kendrick [437]. Kenessey [583]. Kenneth [669, 421, 729]. Kent [734]. Kerber [1173]. Kettil [1084]. Keyfitz [1274]. Khinchin [254]. Kidder [685]. Killed [359]. Kim [1277]. Kingdom [104, 840, 803, 435, 513, 1049, 731, 749, 627, 918, 1024, 9, 302]. Kingman [1002]. Kingston [207]. Kintchine [128]. Kish [1267]. Klaassen [1117]. Klaman [308]. Klare [1139]. Klaus [942, 958]. Klein [101, 340, 37, 1182]. Klein-Goldberger [101]. Knightsfield [644, 838]. Knowledge [231]. Knox [48]. Knud [505]. Kopal [289]. Kotz [1107]. Kovary [332]. Krelle [1150]. Kreweras [739]. Krickeberg [958]. Krupp [295]. Kuehn [538]. Kulldorf [330]. Kumar [241]. Künzi [1150, 1299]. Kurth [256]. Kuznets [346, 768]. Kyburg [754].

L [741, 572, 1073, 464, 638, 1, 840, 1177, 482, 535, 797, 756, 905, 109, 425, 501, 759, 21, 194, 75, 142, 1041, 135, 798, 420, 293, 328, 1334, 661, 985, 253, 576, 1266, 211, 16, 133, 370, 1215, 175, 544, 766, 1043, 76, 1176, 34, 1277, 1336, 96, 533, 265, 1240, 163, 423, 371, 512, 926, 1242]. L. [717]. Labor [206, 1305]. Laboratory [368, 701]. Labour [22, 723, 629, 1264, 252]. Lady [871]. Lagged [1258]. Lags [1265]. Laha [241, 1293, 699]. Lahiri [870, 831]. Lai [331]. Lamartine [310]. Lambe [1102]. Lampman [107, 396]. L’Analyse [829]. Lancaster [644]. Land [855]. Landsberg [231, 243]. Lange [946]. Larger [655]. Larry [898]. Larsson [846]. Lary [545]. Lasorsa [867, 867]. Lassar [706, 830]. Latin [1136]. Latscha [614]. Laurence [1074]. Laurie [570]. Lavers [1276]. Law [936]. Lawley [605]. Lawlike [1159]. Lawrence [1111]. Laws [607, 1135, 1135]. Leabo [665]. Leadbetter [1108]. Learned [1200]. Learning [279, 36]. Leary [336]. l’Économetrie [666]. Lectures [768]. Lee [844, 1332, 1242, 833]. Legality [628]. Lehmann [638, 34]. Lehrbuch [230]. Leiby [206]. Length [282, 655]. Lengths [1199]. Leo [706, 1117, 830]. Leon [945]. Leonard [970]. Leone [798]. Leopold [1039]. Lerner [243]. Leser [1046]. Leslie [1267, 807]. Lester [981]. Level [1197, 770]. Levin [76, 1150]. Levinson [568]. Levy [1183]. Lewes [1112]. Lewis [760, 922, 1005, 565, 938]. Li [737]. Library [28, 54, 80, 118, 152, 442, 488, 519, 553, 589, 624, 650, 687, 718, 746, 775, 812, 851, 890, 911, 947, 990, 1025, 1056, 1088, 1124, 1155, 1191, 1226, 1255, 1280, 1312, 1343]. Liddell [643]. Lieberman [1152]. Life [738, 577, 1290, 576, 834, 752, 656]. Likelihood [356, 315, 102, 598, 355]. Limit [234]. Limits [802]. Lincoln [613]. Lindgren [1295, 634]. Lindley [828, 828]. Line [13, 693, 862, 861]. Linéaire [726, 382, 726]. Linear [707, 1149, 61, 100, 961, 226, 1241, 663, 1149, 603, 960]. Linguistic [621, 537]. Linguistics [845, 67]. Linkage [1109, 334]. Linnemann [982]. Lipsey [1051, 639]. Lipton [1217]. Lissowski [838]. List [28, 54, 80, 118, 183, 214, 245, 152, 276, 312, 350, 405, 367, 955]. Listeners [44]. Listening [44]. Literature [468, 876, 454]. Littauer [659]. Little [108, 1142]. Lives [548]. Living [299]. Ll.D [1154]. Lloyd [274]. Loadings [598]. Loans [842]. Local [1094, 280, 806]. Location [940]. Loève [566]. Logan [143]. Logarithmic [959]. Logarithms [703]. Logic [1001, 924]. Lomas [940]. London [397, 1095, 1054, 410, 1029, 1264, 921, 177, 921, 885]. Long [270, 878, 918, 1024, 1305, 977, 928]. Long-Term [918, 1024, 977, 928]. Look [1128]. Lopez [546]. Lord [507, 1021, 1140, 551]. Love [76]. Loveday [970, 199]. Low [994]. Low-Temperature [994]. Lowell [877, 1023]. Ltd [1166]. Luce [93]. Luck [568, 871]. Lucy [1174]. Luigi [402]. Lukacs [699, 541, 90]. Lydall [1308]. Lynes [339]. Lynth [1178].

M [614, 241, 381, 468, 787, 791, 876, 972, 707, 907, 99, 826, 1108, 1301, 164, 481, 532, 424, 923, 1040, 480, 736, 108, 887, 476, 477, 48, 60, 1206, 439, 1112, 429, 431, 194, 461, 909, 1171, 325, 309, 210, 513, 1293, 1118, 514, 928, 576, 1185, 833, 61, 1104, 664, 549, 285, 1009, 375, 1035, 1142, 940, 499, 1277, 433, 236, 128, 173, 382, 471, 1240, 563, 794, 1140]. M.A [945, 1154]. M.M.P.I [679]. Maass [514]. MacDonald [1006]. Machines [1166, 510, 1044]. Machlup [714]. Machol [131]. Mack [1329]. Maclean [674]. Macroeconomic [715]. Macroeconomics [1218]. Madow [162]. Mahalanobis [145]. Mail [156]. Main [1121]. Maine [1220]. Maintenance [361, 161]. Mair [1220]. Maizels [584]. Major [510]. Majumdar [145]. Makower [1171]. Malaria [966, 943]. Malaya [752, 561]. Malcolm [1244]. Male [123, 752]. Malinvaud [1045, 666]. Malthus [1019]. Man [655, 393]. Man-Power [393]. Management [5, 592, 267, 725, 448, 1166, 16, 1327, 383, 421, 471, 14, 266, 835, 103, 173]. Managerial [673]. Managers [174]. Manea [908]. Manescu [908]. Mann [162]. Manning [204]. Manpower [1185, 1306, 73]. Mansell [703]. Manual [135, 174, 389, 261, 1251]. Manuals [389]. Manuel [730]. Manufacturer [439]. Manufactures [584]. Manufacturing [600, 751, 242, 304, 148, 495, 587, 41, 85, 619, 915]. Many [91]. March [884]. Marcus [94]. Margaret [135, 179]. Marginal [190]. Marglin [514, 506]. Market [763, 1079, 308]. Marketing [538]. Marking [186]. Markkanen [1084]. Markov [288, 129, 755, 1237, 332, 823, 331]. Markus [1242]. Marlow [1216]. Marquand [1179]. Marriages [32]. Marris [674, 673]. Marshall [268, 138, 172]. Martin [142, 711, 1237]. Marvin [94]. Mary [1074, 512]. Masanao [1207]. Masken [514]. Mason [1138]. Mass [819]. Massy [538]. Master [414]. Matalon [829]. Matching [8]. Materials [1329, 194]. Maternal [142, 1336]. Maternity [557]. Mathai [1206]. Mathematical [368, 419, 572, 787, 902, 502, 369, 1038, 160, 498, 1064, 925, 1130, 327, 334, 67, 1235, 1070, 36, 2, 1241, 868, 873, 903, 333, 571, 416, 500, 1134, 69, 128, 1240, 570, 62, 415, 898]. Mathematicians [1294]. Mathematics [316, 1067, 395, 409, 735, 569, 1136, 67, 789, 470, 1206, 66, 1006, 461, 1324, 931]. Mathématique [287]. Mathématiques [640]. Mathematische [698, 1039, 926]. Mathematisches [328]. Mathur [395]. Matrices [1326]. Matrix [540, 952, 94, 971]. Matter [119, 4, 153, 184, 215, 246, 120, 154, 185, 217, 277, 313, 351, 406, 247, 278, 314, 353, 444, 489, 520, 554, 408, 445, 490, 522, 590, 625, 651, 688, 556, 591, 626, 653, 719, 747, 777, 814, 690, 720, 748, 779, 853, 891, 912, 948, 816, 854, 892, 914, 991, 1026, 1057, 1089, 950, 992, 1027, 1059, 1125, 1156, 1192, 1227, 1091, 1126, 1157, 1194, 1256, 1281, 1313, 1344, 1229, 1257, 1282, 1315]. Matthai [1213]. Mature [309]. Maude [437]. Maudsley [49]. Maunder [207]. Maurice [1181, 387, 1326]. Max [1180, 1116]. Maxfield [135]. Maximum [102, 598]. Maximum-Likelihood [598]. Maxwell [605, 291]. May [847, 172, 392]. Maynard [514]. McArthur [341]. McCarthy [707]. McCracken [1012]. McIntosh [1104]. McKean [1222]. McLachlan [1053, 1221, 712]. McMahon [672]. Mean [573]. Means [83, 307]. Measles [7]. Measure [1002, 412, 597, 821]. Measurement [822, 305, 918, 1024, 832, 1049, 1179, 1188, 452, 532, 927, 547, 1335, 395]. Measurements [480, 905]. Measures [239, 630]. Measuring [951, 1078, 839]. Mechanics [335, 824, 256]. Mechanism [1144]. Mechanized [1074]. Medgyessy [257]. Median [573]. Medians [1277]. Medical [1093, 86, 1220, 1053, 1221, 712, 578, 409, 931, 259, 76, 290, 596, 704, 1109, 75]. Medicine [662, 1335, 459, 1219, 931, 1070]. Medie [227]. Medium [1097]. Medium-Term [1097]. Meeting [122, 365, 529, 88, 224, 366, 528, 658, 785, 920, 1066, 1204, 1321]. Megalithic [282, 655]. Melville [202]. Mendel [848]. Mendelson [639]. Mental [1286]. Menzler [208, 1254]. Merát [509]. Merger [109]. Mergers [40]. Merrett [675]. Merrigton [1277]. Merrill [342]. Merrington [972]. Merton [37]. Messina [973]. Messung [1080]. Mesure [1215]. Method [702, 604, 820, 188, 1130, 6, 998, 964, 605, 659, 337]. Methodes [758, 666, 499, 640]. Methodological [258, 73]. Methodology [225, 344, 574]. Methods [968, 459, 609, 450, 379, 663, 966, 164, 582, 1205, 60, 411, 194, 1074, 531, 1004, 58, 827, 1293, 660, 1317, 228, 1173, 997, 72, 760, 1043, 389, 1035, 1103, 1102, 500, 69, 128, 266, 1045, 358, 1300, 1068, 923, 1167, 102, 1291, 565, 564, 427, 664, 65, 1323, 1117, 61, 610]. Metholodogy [825]. Metodologica [867]. Metric [139]. Metropolitan [1264].

Michael [1217, 483, 336, 637, 1121, 52]. Michail [1243]. Michal [1114, 309]. Michel [566]. Middle [274, 548]. Middle-Income [274]. Midlands [940]. Migration [1186, 1162, 558, 724]. Milan [1040]. Mile [10]. Milk [30, 281]. Mill [1140]. Miller [572, 48, 501, 1223, 900, 1047]. Millikan [1180, 1116]. Millions [841]. Mindel [1083]. Mining [242]. Ministry [252]. Mintz [111]. Miscellaneous [701]. Mises [868]. Mitchell [942]. Mitra [1213]. Mitte [934]. Mode [695]. Model [893, 539, 744, 882, 769, 1160, 1076, 1234, 240, 101]. Model-Building [1160]. Modèles [640]. Models [662, 1201, 1106, 1285, 952, 494, 1306, 226, 1233, 1071, 98, 155, 173, 610, 570, 731, 894, 38, 1175, 677, 833, 1173]. Modern [1223, 1212, 65, 1133, 193]. Moments [422]. Monachesi [679]. Monde [764]. Monetary [714, 713, 1144, 987]. Money [879, 204]. Monographs [609]. Monte [964, 823, 1130, 664]. Monthly [347]. Mood [563]. Moon [40]. Moore [347, 1182, 1292]. Moran [791, 92, 1236]. Morbidity [888, 248, 143, 516]. Morgan [1296, 201]. Morgenstern [837, 698]. Morlat [739]. Morrell [797, 1172]. Morris [733, 639, 262]. Morrison [329, 329, 1300, 1068]. Mors [983]. Mortality [857, 123, 820, 1336, 547, 988, 1188, 888, 846]. Mortgage [308]. Mortimer [1271]. Moser [298]. Moses [878, 613]. Mosteller [255]. Motor [433, 510]. Motor-Cars [510]. Motorcycle [455]. Movement [180, 126]. Movements [883, 109]. Moving [782, 852, 863]. Moyle [1050]. Mr [220]. Much [577]. Muench [98]. Mukherjee [676]. Mulhall [337]. Multi [250, 1289, 644]. Multi-Dimensional [250]. Multi-Sector [644]. Multi-Stage [1289]. Multichannel [1214]. Multiple [820, 674]. Multiple-Shift [674]. Multipliers [101]. Multistage [1073]. Multivariate [776, 1323, 525, 1300, 1068, 756].

Munamarty [65]. Mundell [1144].

Murder [340]. Murray [601, 418, 38, 1333].

N [113, 270, 1220, 605, 464, 1, 1249, 923, 1218, 1007, 1067, 395, 620, 376, 1149, 757, 142, 286, 1184, 1309, 386, 962, 834, 145, 1222, 334, 1009, 542, 802, 1251, 1277, 463, 728, 233, 236, 62, 795, 375]. N.H.S [592]. Naas [328]. Nair [870, 831]. Nanda [395]. Nankin [178]. Narain [883]. Nataf [667]. Natal [820]. Nathan [1274]. Nation [839, 1269]. National [342, 104, 71, 476, 477, 772, 271, 339, 430, 338, 1017, 953, 682, 710, 1180, 1269, 396, 508, 639, 770, 643, 701, 984, 385, 977, 941, 818, 1116, 339, 1162, 1318, 146, 147, 434, 1180, 1116]. Nations [884, 579, 1188]. Natural [1079, 1205, 485, 788, 703]. Nature [1177]. Naval [172]. Navy [631]. Neal [485]. Neale [646]. near [237]. Nearest [1201]. Need [209]. Needle [159]. Needs [1286, 209, 310, 303, 218, 885]. Negative [637]. Neglected [695]. Negotiations [457]. Neighbour [1201]. Neil [1341]. Neild [619]. Nelson [109]. Nemchinov [735, 809]. Nesbitt [1336]. Netherlands [187]. Netschert [243]. Networks [873, 972]. Neveu [903]. Nevin [478]. Newcastle [48]. Newcomers [177]. Newman [943]. Newtonian [463]. Neyman [378, 368, 962]. Neyman-Shortest [378]. Nicholas [976]. Nicholson [840, 1304]. Nigam [270]. Nigeria [337]. Nigerian [477, 433]. Nineteenth [542]. Nineteenth-Fifties [542]. Ninth [528]. Nixon [137, 805]. No [646]. Nombres [869]. Non [1149, 491]. Non-Linear [1149]. Non-Orthogonal [491]. Noncomputer [16]. Nonlinear [740, 1150]. Nonparametric [1298, 374, 422, 793, 901]. Norma [341]. Normal [679, 95, 233, 234, 795, 1277]. Normalised [795]. Norman [1106, 636, 798, 660, 788, 1070, 649]. North [1123]. Norway [560, 996]. Note [221, 1201, 862, 160, 863, 1319, 597, 561, 861, 1200, 360, 1289, 358]. Notes [24, 50, 77, 114, 151, 181, 212, 244, 275, 311, 349, 400, 440, 486, 517, 550, 588, 622, 648, 684, 716, 745, 773, 808, 849, 889, 910, 944, 989, 1020, 1055, 1085, 1122, 1153, 1189, 1225, 1252, 1278, 1310, 1339, 57, 1158, 575]. Nove [735]. Number [974, 159, 998, 996, 1037]. Numbers [800, 413, 996, 196]. Numerical [1335, 289, 973, 836, 379, 289]. Nurkse [348]. Nurul [882]. Nutrients [187]. Nutrition [187]. Nutter [436].

O [1073, 498, 1034, 872, 744, 310, 65, 766, 509, 541, 1010]. Oakford [613]. Objectives [39, 514]. Observational [722]. Observations [621, 537, 837]. Observer [822]. Occasional [701]. Occupancy [413]. Occupation [143]. Occupational [1031]. Octav [869]. Oettinger [504]. Oettli [1150]. off [678]. Offenders [711]. Office [112, 340, 646, 172, 386]. Officers [141]. Official [1096, 805]. Ogborn [387]. Okigbo [477]. Old [339, 1262]. Oldham [1335]. Olive [727]. Oliver [705]. Olkin [162]. OMNITAB [973]. One [88, 224, 366, 528, 658, 785, 920, 1066, 1204, 1321, 288]. One-Server [288]. Onicescu [869]. Onore [497]. Oparin [644]. Opened [1032]. Opening [1320]. Operate [1044]. operation [1184]. Operational [1171, 895, 1292, 453, 1239, 155, 198, 384, 385, 1111]. Opérationnelle [640, 764, 739]. Operations [1152, 265, 266, 610, 197, 1327, 15, 383, 933]. Optimal [1242]. Optimization [902, 1207]. Optimum [776, 525]. Order [535, 460, 235, 1329]. Ordinal [1277]. Ordinary [1050]. Ordination [881]. Ordnance [135]. Organe [142]. Organization [309, 1121, 1188, 295, 1187, 579, 1251, 1188, 39]. Organizational [1309]. Organizations [932]. Organizing [172]. Origin [206, 221]. Origins [1179, 463].

Orth [932]. Orthogonal [491]. Oscar [97]. Osculatory [18]. Oskar [946, 837, 115]. Other [573, 819, 820, 679]. Otis [99]. Ottawa [97]. Our [753]. Outcomes [820, 679]. Output [104, 269, 670, 485, 495, 33, 494, 85, 838, 731, 305, 541, 984]. Overseas [306]. Overspill [209]. Owen [469]. Owners [1050]. Ownership [201, 455, 510]. Oxford [510].

P [614, 242, 1036, 922, 572, 791, 870, 45, 397, 1047, 481, 1006, 841, 980, 582, 1337, 1335, 304, 476, 477, 759, 803, 395, 92, 231, 143, 194, 474, 1113, 697, 644, 1236, 831, 661, 1170, 928, 767, 386, 1266, 711, 310, 959, 1329, 145, 230, 983, 1076, 549, 838, 1105, 1292, 1044, 1145, 542, 541, 290, 804, 499, 1277, 337, 426, 1299, 1275, 130, 512]. Paige [104]. Paired [604, 1259]. Paish [435]. Pakistan [882]. Pal [257]. Palin [403]. Palmer [65, 512, 668]. Panel [169]. Pant [870, 831]. Paper [862, 160, 1345, 861, 220]. Papers [507, 187, 962, 1332, 1235, 1101, 1328, 986, 1181]. Parameter [378, 695, 378]. Pareto [936]. Paris [597]. Parking [1016]. Parliamentary [942]. Parnes [512]. Parry [938]. Part [19, 47, 390, 432, 884, 828]. Partanen [1084]. Partial [952]. Partially [330]. Participation [1264]. Particle [194]. Particular [1290]. Party [680]. Parzen [193]. Pascal [685]. Patients [284, 1318]. Patil [959, 1036]. Patricia [1019]. Patrick [899, 1239, 500, 755, 52]. Patronise [44]. Pattern [40, 916, 295]. Patterns [679, 709, 1136]. Paul [1150, 1139, 763, 708, 1042, 842, 969, 263, 1275]. Pauline [733]. Pavlopoulos [1076]. Pay [1307]. Payments [1148]. Peacock [302, 1276]. Pearce [667]. Pearson [1010, 614, 286, 576, 1101]. Pech [437]. Peck [1]. Peirce [342]. Peller [1219]. Pensions [1333]. Pentonville [733]. People [359, 339]. Pèpe [499]. Percentiles [422]. Performance [732, 1163, 1217]. Peri [820]. Peri-Natal [820]. Perinatal [1336]. Period [508]. Periodicals [27, 53, 79, 117, 152, 1279, 1311, 1342]. Permissible [802]. Permutational [1320]. Permutations [613]. Perrier [499]. Persistent [580]. Person [10]. Person-Mile [10]. Personal [780]. Personality [679, 1084]. Personnel [844]. Persons [629]. Perspective [585]. Peston [1185, 1326]. Peter [886, 943, 1303]. Peters [1238]. Petersen [296]. Pharmaceutical [592]. Pharmacy [847]. Phelps [1307]. Phelps-Brown [1307]. Phénomènes [294]. Philip [99, 329, 796, 973]. Phillip [1182, 879]. Phillips [266]. Philosophy [479]. Phyllis [106, 1141, 977]. Physical [269, 798, 337]. Piercy [1021]. Piero [307]. Pietro [1079]. Piggledy [1142]. Pilot [237, 169]. Place [409]. Plackett [163]. Plan [734, 818]. Planned [1309]. Planning [237, 1097, 817, 1230, 722, 1119, 309, 209, 654, 1077, 1180, 940, 1188, 240, 1116, 1098, 831, 514, 1181, 506, 671, 344]. Plans [57, 714]. Plant [751, 848]. Plowden [1260]. Podbielski [1119]. Poisson [378, 1137]. Poker [1032]. Policies [1048, 1017, 1247]. Policy [676, 799, 1143, 209, 807, 940, 1178, 799]. Political [507]. Politics [507]. Population [1274, 46, 99, 84, 341, 317, 1083, 1337, 856, 1064, 546, 952, 1031, 1290, 219, 300, 1233, 943, 389, 574, 575, 1273, 1248, 1188, 1234, 296, 47, 390, 940, 1305]. Populations [722, 1234]. Portugal [1245]. Position [56]. Post [251, 40, 1287]. Post-Secondary [1287]. Post-War [40, 251]. Postwar [308, 508, 439, 768]. Potter [485]. Power [1244, 393]. Prabhu [1007, 757]. Practical [6, 159, 1042]. Practice [1220, 797, 905, 143, 1195, 516, 1240, 105]. Practitioners [1318, 516]. Praktischen [230]. Pratique [967]. Pre [251, 781]. Pre-Test [781]. Pre-War [251]. Prediction [631, 602, 292, 1208, 697]. Preference [1259]. Pregnancy [820]. Preliminary [237, 1185]. Prepared [586]. Prescribing [592]. Present [56]. Presentation [270]. Presidential [121]. Prest [105, 1268]. Preston [1051]. Preventive [459]. Price [361, 161, 883, 1265, 821, 21]. Prices [59, 192, 362, 526, 632, 783, 865, 999, 485, 935, 763]. Pricing [619]. Primer [727]. Principes [287]. Principle [356, 355]. Principles [335, 1077, 907, 268, 259, 1146, 1338, 700, 163]. Priori [532]. Priority [709]. Prison [733]. Private [995, 1247]. Probabilités [736, 1209]. Probabilities [994, 1320, 331, 612]. Probability [368, 419, 638, 1302, 162, 1067, 1169, 868, 860, 637, 325, 993, 420, 255, 417, 827, 956, 1236, 1033, 133, 370, 780, 1008, 701, 463, 421, 795, 130, 828, 193, 791, 254, 64, 1034, 1002, 326, 903, 1132, 754, 871, 828, 958, 607, 63, 1210, 564, 255, 566, 37, 132]. Probable [377]. Problem [994]. Problèmes [758, 764]. Problems [1048, 84, 320, 288, 413, 1171, 209, 654, 289, 319, 1105, 1037, 701, 793, 492, 1237, 894, 1046, 976, 1304, 1273, 930, 546, 1053, 1221, 712, 1172, 545, 266]. Procedures [83, 428, 1173, 1261, 1168]. Proceedings [88, 122, 224, 366, 365, 528, 529, 658, 785, 920, 1066, 1204, 1321, 678, 97, 1245, 543, 172, 368, 884, 1041, 601, 504, 1248, 198]. Process [933]. Processes [936, 1060, 1108, 164, 288, 872, 129, 264, 1035, 788, 1241, 900, 645, 963, 1073, 131, 1007, 1296, 418, 904, 500, 332]. Processing [1044]. Prodan [1243]. Production [436, 301, 283, 150, 85, 505, 1261, 599, 584, 307]. Productivité [1215]. Productivity [104, 305, 449, 448, 85, 447, 453, 984, 452, 1018, 149, 437, 617]. Professor [1190]. Profit [803, 1177]. Profits [935, 1307]. Program [1044, 575, 973]. Programed [706]. Programmation [726, 382]. Programme [540, 539, 670, 44, 769, 836]. Programmed [1324, 830]. Programmes [819]. Programming [663, 1073, 707, 1149, 661, 100, 136, 69, 603, 1240, 534, 502, 740, 1150, 1012, 474, 611, 333, 762, 742, 838, 972]. Programs [1174, 1214]. Progress [893, 274, 1053, 1221, 712, 265]. Project [237, 872]. Projecting [493]. Projections [1188, 981]. Projects [675]. Properties [750, 1108, 101, 573]. Property [201]. Proposal [802]. Proposals [1029]. Proschan [898]. Prose [813, 721]. Prospect [280, 713]. Prospects [310, 243]. Prosperity [339]. Protection [804]. Prys [392]. Przestepska [838]. Psychiatric [1317]. Psychological [796, 372, 844]. Psychology [1266, 996, 660]. Public [141, 44, 202, 302, 1083, 1247, 303, 105, 1268]. Published [797]. Puerto [909, 1186]. Pugh [834, 834]. Punching [1044]. Pupils [1061, 1131]. Purchases [765]. Purchasing [893]. Pursuit [902]. Pyatt [709].

Quadratic [1241, 762]. Qualifications [616]. Qualitative [291]. Quality [450, 451, 1182, 174, 170, 875, 1075, 269, 386, 874]. Quante [230]. Quantitative [1143, 68, 145, 965, 1219, 845, 538]. Quarter [391, 515]. Quarterly [391, 515]. Quenouille [60, 128]. Queueing [1, 425, 128]. Queues [533, 1007, 253]. Quinquennial [146, 147, 734]. Quoted [1050].

R [609, 614, 241, 379, 1220, 900, 922, 378, 870, 1073, 707, 45, 1, 942, 1108, 1139, 93, 470, 668, 424, 923, 705, 761, 797, 800, 205, 1218, 256, 738, 1208, 501, 759, 395, 1147, 20, 262, 466, 602, 611, 971, 786, 1210, 1113, 1324, 679, 564, 1001, 792, 831, 1293, 253, 1213, 1118, 877, 928, 699, 806, 576, 437, 294, 959, 1111, 717, 16, 103, 1222, 1304, 1333, 345, 42, 171, 1145, 643, 12, 36, 939, 619, 1336, 1138, 1305, 1299, 960, 1175, 163, 704, 105, 1268, 458, 1241, 1331, 1270]. R.S.S. [1196]. Race [145]. Ractliffe [369, 1038, 611]. Radhakrishna [145]. Radiation [579]. Radio [819]. Ragnar [167, 348, 344]. Rahman [1067]. Rahmen [1080]. Rail [881]. Railroads [801]. Railway [410]. Raj [270]. Ralph [109]. Ramanujacharyulu [972]. Ramsbottom [25]. Randolph [266]. Random [196, 413, 763, 792, 904, 505, 1043, 233, 700, 613, 130, 936, 974, 417, 418, 503, 1303]. Randomness [422]. Rank [1211]. Rao [870, 831, 1213, 145, 832, 65, 960]. Rapid [60]. Rasaputram [770]. Rasmussen [505]. Rate [438, 284, 1011, 802]. Rates [123, 10, 786, 587, 280, 345, 1264, 656]. Ratio [1288, 1289]. Ratios [838, 917]. Ray [372, 134]. Raymond [508, 639, 70]. Rayner [1142]. Read [466]. Readings [502]. Readship [682]. Real [304]. Reality [1316]. Really [449]. Reappraisal [1197]. Received [183, 214, 245, 276, 312, 350, 405]. Recession [347]. Recherche [640, 739, 764]. Recognition [916]. Recommendations [1146]. Recommended [1253]. Reconciliation [1064, 952]. Reconstruction [40]. Record [1109]. Recording [822]. Recovery [347]. Recurrent [251]. Redish [427]. Redistribution [799, 840]. Reduce [392]. Reduction [451]. Redundancy [806, 647]. Reed [1023]. Rees [304]. Reference [164, 669, 797, 335, 1290, 629, 987, 505, 943, 1035]. Reform [206, 714, 593]. Refraction [336]. Refrigerators [510]. Reginald [623]. Region [1264]. Regional [237, 986, 884, 723, 558, 724, 1161, 148]. Register [1162]. Registers [1318]. Registrar [19, 47, 390, 391, 432, 515]. Registrar-General [19, 47]. Registration [996]. Regression [83, 1158, 820, 61, 1160, 163]. Regulation [697]. Regulations [86, 222, 363]. Reichmann [373]. Reid [129]. Reinterview [575]. Rejoinder [361, 862]. Related [1108, 959, 1234, 95]. Relating [1051, 682, 514]. Relation [446]. Relations [656, 935]. Relationship [82, 285]. Relationships [1258, 1159]. Reliability [569, 898]. Reluctant [512]. Remark [356, 355]. Remarks [1093, 321]. Remuneration [1318]. Renewal [288, 458]. Repeat [1259]. Repeated [1263]. Repercussions [806]. Replication [1015, 1015]. Report [87, 223, 364, 527, 657, 784, 919, 1065, 1203, 1322, 49, 44, 1017, 897, 847, 1082, 1333, 710, 674, 574, 643, 1196, 146, 147, 112, 340, 1185, 615, 142, 786, 1330, 1248, 921]. Reporting [347]. Reports [549, 301, 1082]. Requirements [817, 1030, 629, 318]. Resale [361, 161]. Research [363, 799, 305, 73, 1205, 1219, 932, 44, 1260, 271, 39, 516, 1053, 1017, 1171, 197, 677, 569, 112, 340, 576, 660, 1111, 895, 1317, 931, 710, 156, 453, 643, 1327, 155, 15, 172, 198, 266, 383, 385, 933, 894, 984, 941, 260, 712, 1152, 807, 1173, 760, 1180, 1292, 1239, 198, 265, 384, 385, 610, 372, 704, 941, 965, 1116, 582, 620, 962]. Research/Management [383]. Reserve [150]. Residential [1079, 308, 730]. Resolution [1063]. Resource [485, 514]. Resources [310]. Respecting [221]. Responses [1069]. Restatement [209]. Restraint [1029]. Restricted [1037]. Results [446, 1164, 942, 823, 757, 1107, 1121, 793, 836]. Retrospect [713]. Return [391, 515, 587, 1186]. Returns [283]. Revell [1050, 1269]. Revenue [710]. Reverend [322, 443]. Review [680, 681, 1187, 968, 22, 206, 393, 1109, 1048, 683, 835, 1297, 183, 214, 245, 276, 312, 350, 405, 113, 237, 270, 676, 459, 609, 614, 662, 702, 507, 241, 342, 936, 1046, 104, 242, 540, 678, 799, 343, 368, 381, 379, 469, 1036, 388, 1220, 1274, 287, 419, 663, 900, 922, 605, 645, 741, 257, 378, 467, 468, 572, 787, 789, 790, 791, 870, 876, 972, 1106, 1137, 1079, 534, 902, 963, 1073, 1207, 269, 484, 131, 502, 707, 740, 1150, 369, 464, 530, 638, 907, 1038, 864, 881, 1050, 45, 46, 99]. Review [137, 141, 208, 296, 297, 298, 341, 547, 615, 1, 71, 305, 942, 97, 329, 267, 397, 1069, 727, 888, 966, 1083, 1249, 715, 826, 901, 1012, 1108, 1300, 199, 840, 1047, 1177, 986, 1301, 164, 196, 481, 482, 497, 498, 1245, 1302, 642, 1139, 93, 19, 47, 390, 391, 432, 515, 73, 974, 1068, 268, 470, 479, 532, 535, 543, 1114, 1006, 436, 539, 668, 669, 670, 841, 883, 424, 567, 238, 923, 1005, 1167, 705, 761, 797, 800, 1040, 1216, 884, 1014, 1115, 980, 1217, 38, 205, 480, 582, 866]. Review [1205, 239, 89, 254, 756, 905, 1271, 1272, 1337, 301, 604, 1218, 49, 958, 1335, 1143, 64, 162, 736, 1034, 1209, 74, 665, 671, 40, 108, 109, 176, 304, 585, 878, 887, 732, 673, 148, 824, 1002, 274, 308, 476, 477, 48, 256, 335, 425, 738, 1007, 1208, 60, 1206, 1296, 772, 68, 195, 330, 462, 501, 537, 759, 1067, 347, 439, 803, 1112, 288, 395, 1147, 1219, 666, 982, 1051, 20, 483, 21, 429, 867, 92, 262, 438, 616, 620, 932, 465, 1072, 930, 1323, 583, 44, 376, 607, 929, 1169, 485, 231, 326, 380, 636, 868]. Review [869, 143, 336, 102, 167, 763, 260, 1267, 144, 271, 339, 431, 546, 577, 579, 846, 1019, 1149, 273, 428, 63, 66, 194, 757, 203, 435, 672, 39, 207, 75, 430, 466, 516, 461, 474, 602, 611, 829, 967, 971, 1000, 1053, 1214, 1221, 712, 307, 348, 478, 232, 1168, 1291, 1231, 1346, 142, 1041, 786, 135, 338, 975, 909, 843, 934, 1017, 1119, 578, 637, 1074, 1171, 286, 737, 798, 925, 565, 146, 147, 1210, 1186, 346, 587, 1113, 1184, 1018, 531, 608, 970, 1324, 325, 633, 679, 420, 564, 165, 293, 606, 697, 13]. Review [197, 255, 101, 1084, 568, 706, 886, 309, 644, 17, 177, 210, 1054, 1004, 581, 682, 734, 90, 166, 227, 328, 417, 601, 659, 924, 1001, 504, 566, 792, 827, 872, 873, 899, 903, 956, 1236, 1152, 434, 209, 272, 844, 333, 1011, 1334, 1003, 1107, 513, 667, 677, 709, 743, 744, 831, 832, 882, 981, 1049, 1080, 661, 735, 769, 985, 1293, 253, 472, 1170, 1213, 1238, 1295, 1118, 327, 514, 877, 928, 149, 150, 569, 1309, 418, 612, 699, 904, 1211, 767, 806, 979, 111, 306, 1172, 386, 112, 1298, 340, 576, 580, 733, 807, 41]. Review [35, 373, 937, 258, 635, 660, 796, 1212, 1266, 711, 714, 1078, 1179, 1185, 1117, 310, 437, 1181, 387, 294, 23, 377, 571, 825, 959, 962, 506, 762, 833, 834, 1016, 1111, 1329, 976, 801, 1294, 61, 225, 1166, 129, 145, 839, 1033, 211, 16, 100, 103, 228, 37, 847, 536, 880, 987, 133, 370, 931, 106, 169, 1082, 1215, 1222, 1332, 908, 753, 1104, 229, 230, 300, 427, 1304, 175, 1077, 1173, 140, 259, 708, 927, 983, 1013, 1076, 1182, 1247, 1326, 1081, 1015, 1250, 1333, 334, 136, 174, 264, 289, 664, 742, 964, 107, 110, 204, 505]. Review [398, 138, 399, 65, 1042, 549, 72, 201, 303, 760, 765, 957, 961, 1121, 200, 710, 1183, 1306, 226, 285, 460, 1330, 1141, 845, 1136, 544, 1180, 132, 503, 906, 1151, 943, 838, 43, 1009, 1105, 1099, 243, 766, 1043, 1269, 396, 475, 508, 639, 202, 1132, 302, 345, 42, 70, 168, 344, 76, 1292, 67, 647, 674, 1044, 389, 574, 575, 171, 375, 1035, 1146, 545, 675, 1145, 770, 1223, 170, 394, 542, 586, 617, 805, 837, 978, 713, 802, 842, 1243, 299, 509, 548, 643, 473, 541, 538, 1176, 935, 1239, 1327, 12, 416]. Review [634, 1037, 1235, 261, 973, 584, 1071, 392, 788, 830, 848, 969, 1070, 1103, 1273, 1142, 1008, 1102, 988, 98, 290, 1251, 804, 1135, 940, 36, 499, 500, 701, 754, 939, 1052, 1134, 618, 619, 768, 34, 374, 422, 793, 1277, 771, 337, 433, 646, 1248, 1270, 1336, 1338, 426, 463, 984, 1308, 728, 730, 938, 1138, 1148, 1188, 1305, 1325, 240, 139, 1303, 18, 94, 95, 96, 233, 234, 235, 236, 533, 511, 1244, 700, 871, 1110, 1276, 14, 15, 69, 128, 172, 173, 198, 263, 265, 266, 295, 382, 383, 384, 385, 421, 471, 573, 603, 610]. Review [640, 641, 698, 726, 739, 755, 758, 764, 933, 1240, 1299, 1174, 292, 372, 729, 960, 1275, 1045, 1175, 570, 731, 178, 62, 163, 423, 563, 1010, 1101, 879, 977, 510, 134, 291, 371, 613, 703, 704, 794, 795, 941, 836, 1237, 91, 130, 331, 332, 415, 828, 874, 875, 885, 921, 965, 1075, 1120, 1133, 1178, 1307, 105, 512, 1268, 1144, 1140, 926, 193, 458, 1241, 1242, 1328, 179, 1116, 1331, 1246, 898, 1039, 1324, 204]. Reviews [51, 367, 955, 1224]. Revised [72]. Revision [150]. Rhetoric [258]. Rhodes [810]. Rice [1340]. Richard [540, 534, 539, 38, 108, 868, 431, 434, 1049, 377, 136, 264, 742, 374, 512, 898]. Richmond [1327]. Rico [909, 1186]. Ridley [1083]. Riordan [472]. Risk [392]. Risks [392]. Rivett [1239]. Road [881, 323, 29, 359, 392]. Robb [1266]. Robbins [1140, 507]. Robert [680, 381, 1106, 131, 199, 582, 74, 1051, 1019, 925, 970, 606, 197, 255, 1003, 514, 801, 169, 107, 399, 396, 639, 973, 36, 1336, 1276, 383, 613, 1144]. Robertson [399]. Robin [673, 674]. Robinson [479, 1214, 503, 91]. Robson [1140]. Robustness [83]. Rodwin [274]. Roegen [976]. Roelandts [433]. Roessler [133, 370]. Roger [16, 103]. Role [478, 710, 155]. Roma [142]. Román [541]. Romig [17]. Ronald [401, 609, 40, 1000, 441, 1223, 586, 538, 848]. Rooms [855]. Rosamund [342]. Rosenblatt [601, 418]. Rosenhead [572]. Rosenthal [383]. Rostow [678]. Roth [1016]. Rourke [255]. Rowe [1049]. Roy [241, 203, 258, 713, 618, 197, 1293]. Royal [49, 88, 224, 366, 365, 528, 658, 785, 920, 1066, 1204, 1321, 631]. Rt. [1021]. Rudra [241]. Rufus [902]. Rumours [1062]. Runnenburg [288]. Russell [265]. Russia [358]. Ruth [177, 1329, 885]. Ryan [21].

S

[968, 1297, 113, 676, 614, 241, 242, 870, 502, 46, 1249, 1177, 164, 482, 535, 162, 732, 1002, 48, 68, 462, 501, 395, 929, 577, 971, 142, 1171, 286, 325, 13, 101, 706, 210, 1152, 333, 1107, 743, 831, 832, 981, 735, 985, 1213, 1238, 340, 35, 834, 100, 174, 399, 65, 1136, 1105, 1035, 1176, 1235, 392, 830, 1134, 1277, 383, 933, 372, 1010, 1101, 134, 794, 415, 1075, 512].

Saaty [425]. Sabine [1243]. Salter [149]. Salzer [18]. Sam [243, 1336]. Samet [661]. Sampford [424]. Sample [1108, 413, 207, 953, 998, 261, 701, 793, 573, 260]. Samples [330]. Sampling [559, 57, 424, 428, 249, 65, 1261, 261, 1289, 573, 358, 567, 1267, 17, 1303, 968, 1072, 17, 228]. Samuel [963, 686, 132, 1327, 69, 421, 518]. Samuels [544]. Sandee [240]. Sargant [767]. Sarhan [460]. Satya [141, 966]. Saul [663, 308]. Saunders [1022, 70]. Savage [374, 423]. Saving [995, 803, 1148]. Savings [510]. Sawers [1223]. Saxena [62]. Scale [283, 1184, 654, 1244, 484]. Scarf [421]. Schaafsma [1037]. Scheffe [89]. Scheme [985]. Schipper [760]. Schlesinger [1336]. Schmetterer [1039]. Schmid [328]. School [1163, 1054, 847]. Schools [1196]. Schumaker [465]. Schumpeter [306]. Schurr [243]. Science [258, 825, 225, 931, 139, 383, 421]. Sciences [971, 499, 173, 727, 1139, 1040, 1205, 501, 798, 1111, 788, 570]. Scientific [1245, 567, 616, 606, 399, 319]. Scientist [210]. Scientists [379, 1006, 932, 877, 1072, 1013]. Scores [1277]. Scotland [1220, 1162, 43, 399]. Scott [172, 298, 513, 487]. Scottish [43, 43]. Seal [756]. Searle [971, 552]. Seasonal [782, 852, 863]. Sebastian [659]. Seber [961]. Second [920, 199, 1053, 146, 147, 1128, 235, 198]. Secondary [1287]. Section [86, 222, 363]. Sector [270, 1268, 644]. Security [238, 272]. Seeger [1042]. seit [934]. Selected [329, 883, 127, 344, 1101, 1328]. Selecting [631, 1117]. Selection [354, 413, 273]. Selectiveness [44]. Self [172]. Self-Organizing [172]. Selin [906]. Sen [113]. Sengupta [1143, 233]. September [1248]. Sequential [57, 189, 998, 1004, 290]. Sergeevich [809]. Serial [954, 1165, 576]. Series [799, 315, 922, 782, 852, 863, 532, 823, 1100, 1053, 1214, 601, 872, 959, 345, 826, 602, 165]. Serious [392]. Server [288]. Service [472, 1331, 1162, 1318, 338, 975]. Services [592, 44]. ses [755]. Sets [510]. Seven [16]. Seventh [224]. Several [793]. Sex [1337]. Shackle [1177, 535, 985, 1176]. Shapiro [1336]. Share [821, 107, 396]. Shares [1050]. Shaul [623]. Sheet [496, 1269, 639]. Sheps [1083]. Sheridan [336]. Sherman [295]. Shewhart [1086]. Shift [674]. Shigeo [713]. Shiskin [347]. Short [270, 1061, 1131, 1100, 882, 261, 977, 766]. Short-Sighted [1061, 1131]. Short-Term [270, 882, 977, 766]. Shortest [378]. Should [139, 593]. Show [705]. Shreider [964]. Shrikhande [870, 831]. Sidák [1211]. Sighted [1061, 1131]. Sigismund [1219]. Signals [347]. Significance [614, 1263, 170]. Signs [822]. Silvey [1045]. Simaika [928]. Simey [210, 210]. Simon [346, 674, 768, 977]. Simonnard [382]. Simple [630]. Simplified [893, 1266]. Simpson [1220, 239]. Simulation [1100, 641]. Sinaiticus [232]. since [676, 397, 239, 878, 111, 345]. Singapore [178]. Single [17, 289]. Sir [401, 1087, 441, 1022, 403]. Sisal [979, 23]. Sisson [16]. Situations [523]. Six [1332]. Sixth [88]. Sixties [672]. Sixty [167]. Sixty-Fifth [167]. Size [751, 7, 82, 1332, 1337]. Skill [1202]. Skorokhod [904]. Slater [96, 1174]. Small [995, 1184, 360, 500, 701, 1270, 573, 484, 194]. Small-Company [1270]. Small-Sample [701, 573]. Small-Scale [484]. Smart [578]. Smirnov [235, 795, 473]. Smith [683, 194, 1041, 1270, 482, 665, 66, 253, 708]. Smokler [754]. Smoothing [602]. Snedecor [1291]. Snell [570]. Snow [78]. Social [540, 693, 862, 141, 298, 238, 1115, 205, 410, 438, 1072, 431, 1184, 210, 861, 807, 937, 1111, 1013, 896, 1330, 475, 940, 570, 38, 272, 807, 988]. Socialist [671]. Societies [1200, 198]. Society [680, 208, 88, 221, 224, 366, 365, 528, 658, 785, 920, 1066, 1204, 1321, 1140]. Socio [113, 237]. Socio-Economic [113, 237]. Sociological [733]. Sociology [767]. Solodovnikov [293]. Solomon [500, 292]. Soltow [1332]. Solution [930]. Solutions [1286, 235, 1067]. Some [1093, 237, 1164, 1345, 1158, 190, 273, 953, 58, 1286, 33, 629, 834, 523, 654, 1042, 43, 1105, 575, 996, 836, 1119, 385]. Sondages [967]. Sorsby [336]. Sorties [1246]. Sorting [1044]. Source [66, 1051]. Sources [1253, 320, 72, 303, 319]. Southern [337]. Soviet [484, 436, 1077, 269]. Sozialstatistik [230]. Space [1016]. Spaces [1136]. Special [365, 529, 164, 669, 797, 629, 987, 505, 943, 1035, 235]. Specification [428, 249]. Spectral [1100, 826]. Spencer [481]. Spent [112]. Spiegelman [1271]. Spitzer [700]. Sponsored [172]. Spurr [482, 1170]. Squares [1136]. Sraffa [307]. St. [549]. Stabilization [1143]. Stable [823, 546]. Stacey [179]. Staff [251, 897]. Stafford [14]. Stage [1289]. Stanback [439]. Standard [573]. Standardisation [22]. Standards [701]. Stanley [536, 686]. Starke [679]. State [886, 110]. States [1079, 176, 148, 485, 981, 508, 545, 984, 730, 7, 1286, 104, 1047, 437, 639, 1336, 297]. Stationary [893, 331, 1108]. Statistica [867, 867]. Statistical [51, 88, 27, 53, 79, 117, 224, 152, 366, 365, 528, 658, 785, 920, 1066, 1204, 1321, 1253, 609, 799, 368, 381, 469, 750, 922, 605, 468, 876, 530, 84, 137, 298, 160, 1300, 19, 47, 390, 432, 1068, 923, 1167, 866, 1205, 756, 49, 695, 176, 824, 256, 335, 60, 1206, 1296, 462, 395, 354, 1095, 1169, 194, 454, 232, 1168, 286, 191, 564, 293, 255, 628, 659, 924, 1213, 523, 1317, 654, 37, 557, 1013, 1076, 174, 964, 226, 859, 1044, 394, 973, 780, 1251, 1135, 701, 34, 1262, 337, 646, 1270]. Statistical [463, 2, 492, 960, 423, 510, 1075, 1328, 221, 459, 702, 966, 480, 239, 74, 738, 501, 1291, 1074, 565, 832, 1170, 1238, 1295, 569, 635, 660, 825, 225, 300, 1151, 168, 634, 1103, 1102, 1138, 471, 1045, 371, 704]. Statisticians [594, 595, 596, 1010]. Statistics [1093, 1187, 393, 559, 1097, 368, 388, 269, 5, 141, 615, 1230, 727, 1139, 73, 691, 1006, 705, 1040, 162, 665, 671, 1094, 176, 887, 1197, 1067, 1051, 607, 694, 29, 485, 143, 579, 63, 194, 516, 971, 1162, 970, 1324, 409, 6, 706, 1128, 1286, 897, 723, 306, 1298, 1266, 959, 962, 1304, 1161, 285, 1096, 1105, 252, 830, 701, 1134, 422, 793, 324, 646, 915, 573, 828, 1120, 593, 1196, 1195, 1098, 206, 419, 870, 369, 638, 907, 1038, 46, 341, 199, 482, 498, 1302, 668, 797, 1014, 929, 868, 466, 578, 737, 925]. Statistics [608, 970, 325, 420, 568, 1004, 1293, 327, 41, 373, 796, 1166, 133, 370, 175, 259, 1250, 200, 460, 1105, 76, 805, 416, 1235, 969, 1008, 374, 1325, 62, 563, 415, 874, 1133, 22, 267, 1069, 1301, 761, 797, 1205, 1112, 1147, 465, 135, 798, 531, 272, 877, 35, 1294, 1104, 12, 372, 965]. Statistics. [464]. Statistik [230, 698, 926, 1039, 229]. Statistique [287]. Statistiques [666, 499]. Status [1286]. Steel [599]. Steering [1099]. Stefan [102]. Steffensen [404]. Stegun [787]. Stein [356]. Steindl [936]. Stephen [514, 506]. Sterling [672]. Stevens [1074]. Stichproben [229]. Stigler [587]. Stochastic [900, 1060, 963, 1207, 1285, 1108, 164, 1007, 872, 1035, 788, 1234, 1284, 662, 472, 1071]. Stochastically [1145]. Stock [627, 763, 1329, 879]. Stocks [1172]. Stoller [933]. Stone [540, 539, 38, 336, 434, 1049, 434]. Storage [757, 928, 92, 1172]. Straight [13]. Straight-Line [13]. Strategies [1319]. Strategy [933]. Street [178]. Structural [832, 511, 1106, 583]. Structure [1297, 237, 39, 1031, 959, 880, 201, 30, 281, 1308, 1141]. Structures [1200, 612]. Stuart [534, 1140, 1340, 567, 285]. Student [1164, 208]. Students [141, 251, 1040, 897].

Studi [497]. Studies [305, 722, 188, 959, 398, 544, 43, 299, 701, 512, 579, 1117, 1270, 241, 1114, 803, 20, 435, 904, 639, 36, 754, 421, 292, 977]. Study [237, 936, 446, 994, 1220, 99, 298, 888, 631, 40, 274, 48, 1007, 395, 982, 1100, 431, 1163, 39, 232, 909, 843, 1084, 1054, 513, 733, 494, 1329, 85, 169, 760, 243, 42, 548, 1052, 619, 337, 179, 732, 928, 145, 647]. Stuvel [799, 937]. Subject [454]. Subjective [754]. Success [631]. Suggested [1286]. Suggestions [39]. Suits [965]. Sukhatme [841]. Sum [759]. Summary [1333, 1248]. Summer [1245]. Summers [1238]. Superpositions [257]. Suppes [500]. Supplement [358, 176]. Supplementary [190]. Supplies [218, 858]. Supply [320, 430, 319]. Supporting [71]. Surnames [1199]. Survey [113, 237, 1050, 1062, 207, 953, 189, 581, 1198, 979, 23, 895, 1261, 574, 575, 158, 261, 510, 1283, 806, 1267]. Surveys [776, 75, 58, 682, 696, 525, 1015, 156, 228]. Survival [284]. Sustained [678]. Sverdrup [1135, 1135]. Sveshnikov [1043]. Swaroop [966, 141]. Sweden [846]. Swing [1232]. Swings [878, 1305]. Sykes [675]. Sylvain [659]. Sylvia [338]. Symbolic [474]. Symmetric [1009]. Symmetry [793]. Symposium [368, 97, 20, 1041, 601, 504, 847, 299]. Syski [262]. Syskind [838]. System [71, 710, 139, 150, 714]. Systèmes [869]. Systems [1207, 197, 514, 172, 263, 1207, 262, 293, 472, 503, 1151, 471, 476, 937]. Szabady [1273].

T [680, 239, 425, 485, 231, 925, 210, 1011, 576, 807, 377, 129, 211, 334, 302, 42, 1235, 788, 1070, 385, 332]. Table [412, 190, 258, 614]. Tableaux [1246]. Tables [498, 1302, 47, 390, 411, 428, 17, 682, 1290, 250, 1213, 576, 541, 1008, 752, 1010, 704, 656, 235, 795, 469, 572, 787, 1, 1011, 959, 1009, 701, 614, 378, 71, 637, 18, 95, 233, 234, 235, 236, 573, 613, 703, 795]. Tabulating [76]. Tage [846]. Tagore [237]. Takacs [533]. Take [678, 40]. Take-off [678]. Take-Over [40]. taken [135]. Talbot [337]. Tally [11]. Tanganyika [979, 23]. Tariffs [127]. Tastes [44]. Tavistock [1052]. Tax [1017]. Taxes [710]. Taylor [1002, 981]. Teach [1171]. Teacher [1104]. Teaching [6, 593, 1196]. Team [872, 643]. Technical [149, 523, 1223, 1184]. Technique [190, 635, 1015]. Techniques [1046, 1171, 514, 169, 289, 538, 1138, 385, 759, 173]. Technischen [1080]. Technological [616, 1244]. Technology [847, 172, 103]. Tee [964]. Telephone [262, 873]. Television [951, 819, 510]. Temperature [994]. Ten [509]. Tercentenary [524]. Terence [733]. Terent’ev [236]. Terleckyj [620]. Term [270, 1097, 918, 1024, 882, 977, 928, 766]. Terms [805]. Test [954, 1165, 781]. Testing [614, 1259, 60, 1037, 34]. Tests [1263, 750, 844, 998, 847, 1211]. Tetley [464]. Tew [20]. Text [1266, 1145, 464]. Text-book [464]. Textbook [706, 67, 830]. Texts [609]. Th [288]. Their [1164, 269, 1108, 548, 935, 917, 298, 48, 1007, 44, 504, 129, 541]. Theodore [645]. Theorems [380, 94]. Theoretic [1134]. Theoretical [93, 1306]. Théorie [667, 294, 967, 755]. Theories [1064, 571]. Theory [381, 900, 645, 467, 1106, 902, 1073, 1285, 715, 532, 1114, 424, 797, 800, 254, 1143, 673, 425, 330, 92, 262, 636, 868, 8, 1259, 1129, 873, 899, 956, 569, 904, 825, 225, 129, 100, 334, 285, 1043, 996, 1176, 804, 701, 618, 533, 871, 69, 128, 563, 1284, 105, 193, 1241, 1283, 1195, 898, 741, 958, 288, 21, 1169, 380, 546, 757, 1041, 566, 827, 1236, 1003, 1107, 743, 985, 1295, 1033, 961, 906, 634, 36, 94, 371, 458, 1242, 905, 64, 326, 1211, 261, 332]. Thermal [1244]. Thermodynamics [335]. Third [1066, 391]. Thirtieth [658]. Thirty [785, 920, 1066, 1204, 1321]. Thirty-Fifth [1321]. Thirty-First [785]. Thirty-Fourth [1204]. Thirty-Second [920]. Thirty-Third [1066]. Thomas [1220, 986, 439, 262, 1019, 322, 443, 293, 255, 514, 1081, 765, 1134, 1178, 255, 398]. Thompson [703]. Thomson [160]. Thonstad [1306]. Thor [935]. Thorbecke [1143]. Thorp [1034]. Thought [1176]. Three [336]. Throughout [548]. Tibor [583]. Tiers [764]. Tiers-Monde [764]. Time [315, 1207, 826, 782, 852, 863, 532, 535, 823, 288, 44, 602, 1214, 601, 872, 1233, 757, 165, 112]. Time-Dependent [757]. Time-Series [165]. Times [176, 463, 1303]. Tindall [232]. Tisserand [499]. Tisserand-Perrier [499]. Titchen [383]. Titmuss [431]. Tocher [641]. together [345]. Token [67]. Tolerance [234]. Tolerance-Limit [234]. Tome [640]. Tony [339]. Tool [5]. Top [107, 396]. Topical [203]. Topics [1207]. Tore [1306]. Tortrat [287]. Total [389]. Totals [190]. Touko [1084]. Tour [264]. Town [1230]. Towns [298, 1099]. Townsend [886]. Tracts [794]. Trade [884, 982, 485, 127, 306, 880, 300, 360, 42, 770, 584, 804, 618, 619, 111]. Trader [545]. Tradition [1176, 179]. Traffic [741, 571, 433, 1029, 873, 1099]. Training [594, 595, 596]. Transactions [496, 1078, 360]. Transformations [792]. Transition [331]. Transmission [381]. Transport [881, 323, 10, 29, 300]. Transportation [972, 202, 941]. Transvariazione [166]. Travel [1019]. Travis [804]. Treated [284]. Treatment [1135]. Treatments [494]. Trend [901]. Trends [1048, 676, 297, 317, 857, 495, 125, 310, 437, 859, 1141, 584, 1188, 270, 888, 485, 771]. Trial [112]. Trials [75, 964, 290, 140]. Tribute [524]. Triennial [49]. Triple [630]. Tropical [509]. Trustees [1101]. Tucker [419, 1132]. Turner [1048]. Turney [1266]. Tutor [1166]. Twenty [88, 224, 366, 528]. Twenty-Eighth [366]. Twenty-Ninth [528]. Twenty-Seventh [224]. Twenty-Sixth [88]. Twins [1084]. Two [1220, 187, 258, 1288, 793, 1029, 258]. Tyne [48]. Type [67]. Type-Token [67]. Types [1121].

U [1007, 757, 1184, 1251]. U.K [1220, 1269]. U.N [1115]. U.S [111]. U.S. [574, 575]. U.S.A [1220]. Uemura [966]. Ulf [63, 612]. Ulmer [202]. Umarji [564]. Unbiased [378]. Uncertainty [1052, 171]. Underdeveloped [1048, 207, 478]. Undergraduates [1067]. Underground [410]. Undernourishment [561]. Understand [607]. Undeveloped [881]. Unemployment [558, 511]. Unilever [586]. Union [676, 484, 436, 843, 1077]. Unit [112, 340, 394, 282]. United [104, 1079, 297, 840, 1047, 884, 176, 148, 803, 485, 579, 435, 513, 981, 1049, 437, 508, 639, 302, 545, 1336, 984, 730, 1188, 731, 7, 749, 627, 918, 1024, 9, 1286]. Units [655]. Unity [231]. Universities [493, 271, 1180, 941, 1116]. Universities-National [1180, 1116]. University [237, 368, 251, 395, 273, 1163, 601, 897]. Unternehmensforschung [1299]. Upon [1163, 48]. Upper [178]. Urban [581]. Urgent [1286]. Urquhart [887]. Use [1253, 141, 1014, 288, 607, 579, 428, 188, 1084, 735, 491, 200, 541, 1175, 855, 373]. Uses [430]. Utilisation [674]. Utilities [202]. Utilization [1185, 303]. Utting [339].

V [1046, 841, 254, 578, 925, 293, 873, 735, 904, 433, 236, 1275, 795, 828]. Vajda [1136, 502, 333, 100, 1136]. Valavanis [102]. Valerie [940]. Value [360]. Values [195, 236]. Variability [696]. Variable [820, 289]. Variables [417, 792, 130]. Variance [89, 1298, 957, 1015, 505, 1042]. Variates [91]. Variations [1135, 505, 802]. Various [10, 573]. Vasilii [809]. Vehicle [433]. Vera [243]. Verdoorn [582]. Verhulst [173]. Vernon [565]. Vesselo [466]. Veverka [302]. VI [884]. Victor [201, 1331]. Victoria [693, 862, 861]. Victoriana [191]. View [723, 1284]. Viewers [44]. Viewing [44]. Viewpoint [828]. Viewpoint.Part [828]. Views [1273]. VII [106]. Village [1118]. Villages [237]. Vincent [1074, 1215]. Viscount [518]. Visva [237, 237]. Visva-Bharati [237, 237]. Vital [1187, 388, 1274, 46, 579, 1250]. Vitality [138]. Vol [1187, 305, 884, 347, 226, 285, 261, 830, 1135]. Vols [1328]. Volume [3, 81, 216, 352, 407, 521, 555, 652, 689, 778, 815, 913, 949, 1058, 1090, 1193, 1228, 1314, 143, 706, 106, 136, 1251, 1248, 1188, 1241, 63]. Voter [642]. Voting [680]. Vue [736].

W [683, 678, 379, 922, 378, 137, 826, 196, 481, 482, 268, 923, 866, 239, 64, 162, 40, 48, 738, 1296, 21, 143, 260, 273, 194, 435, 207, 1291, 142, 135, 792, 1238, 1295, 327, 149, 979, 580, 373, 437, 23, 834, 801, 204, 549, 1141, 508, 639, 805, 978, 634, 1037, 1235, 1103, 804, 433, 266, 384, 703, 512]. W. [160]. Wachstum [934]. Wachstumsprozesses [1080]. Wadsworth [130]. Waerden [926]. Wage [1048, 345, 457, 1048, 1179]. Wages [304]. Wagner [471]. Wahrscheinlichkeitsrechnung [698]. Waiting [288]. Waiting-Time [288]. Wales [19, 47, 390, 391, 432, 515, 142, 576, 209, 1110]. Walk [700]. Walker [142]. Wallace [983, 1113]. Walley [1277]. Wallis [386]. Walsh [973, 1298, 422, 793]. Walter [800, 1041, 253, 1086]. Walther [934]. Walton [48]. Wani [959]. Want [449]. War [40, 860, 251, 878]. Warfare [902]. Waring [1129]. Warnasena [770]. Warren [436, 871]. Washing [510]. Waste [451]. Water [514]. Water-Resource [514]. Waterville [1220]. Wavelets [503]. Wealth [799, 305, 106, 1332, 107, 1269, 396, 508, 305, 677, 544, 70, 839]. Wealth-Holders [107, 396]. Weather [1028]. Weaver [871]. Wedderburn [886, 647]. Weekly [345]. Wegner [474]. Weinberg [465]. Weiss [371]. Weissberg [234]. Welfare [886]. Wells [732]. Welsh [398, 1121]. Werner [1150]. West [177, 103, 940, 173]. Westergaard [885]. Western [297, 1119]. Wetherill [1167, 1004]. Wheat [357, 771]. Whelpton [685]. which [392]. White [647]. White-Collar [647]. Whittle [697]. Whitworth [993]. Wholesale [59, 192, 362, 526, 632, 783, 865, 999]. Wide [127]. Wilcock [512]. Wilhelm [1274, 1150]. Wilkins [807, 211]. Wilkinson [1041]. Wilks [686, 1235, 929, 415]. Willard [1220, 335]. William [609, 296, 1302, 583, 1041, 197, 956, 1170, 1238, 403, 509, 538, 1008, 36, 372, 134, 993, 552, 1341]. Williams [834, 392, 1149]. Williamson [637, 1171]. Willner [195]. Wilson [986, 811]. Wine [877]. Wirtschaft [934]. Wirtschafts [230]. Wisdom [1262]. Wiseman [302]. within [48]. Witold [838]. Witts [75]. Wladimir [577]. Wm [1270]. Wolberg [1208]. Wold [872, 744]. Wolf [298, 420]. Wolfowitz [380]. Woman [548]. Women [548]. Woodward [759]. Wooldridge [611]. Woolley [1078]. Woolton [774]. Work [577, 1231, 1346, 1054, 1266, 548, 643, 512, 1213, 674]. Work-Text [1266]. Workable [715]. Workers [141]. Working [1031, 1290, 1099, 548, 752]. World [1187, 317, 857, 579, 348, 1078, 545, 584, 1251, 218, 858, 1248, 1188, 618, 1187, 1274, 1273, 1188]. Worswick [542]. Wörterbuch [328]. Woytinski [577, 577]. Wright [206]. Wrigley [142]. Writings [329].

X [799]. Xenophon [987].

Ya [254]. Yates [310, 228, 704]. Year

[19, 47, 390, 432, 167, 284, 724].

Year-by-Year [724]. Yearbook [144].

Years [208, 49, 993, 359, 1176]. Yields [357]. York [271]. Young [48]. Yourself [1171]. Yovitz [172]. Yu [964].

Z [1106, 541]. Zadek [293]. Zarkovich [968, 1075]. Zbynék [1211]. Zdenek [289]. Ziderman [1185]. Ziegler [973]. Zolotas [987]. Zoltan [583]. Zwet [792].

References

Benson:1959:BRB

[1] F. Benson. Book review: Finite Queueing Tables, by L. G. Peck; R. N. Hazelwood. Journal of the Royal Statistical Society. Series A (General), 122(1): 109–110, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343064. See cor-

rigendum [37].

Thomson:1959:EMS

[2] W. E. Thomson. Ernie — a mathematical and statistical analysis. Journal of the Royal Statistical Society. Series A (General), 122(3):301–333, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342795. See note [160].

Anonymous:1960:VIa

[3] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 123(1):??, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343182.

Anonymous:1960:FM

[4] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 123(1):??, ????

1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343183.

Beaver:1960:STM

[5] Hugh Beaver. Statistics as a tool of management. Journal of the Royal Statistical Society. Series A (General), 123 (1):1–9, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343184.

Jowett:1960:PET

[6] G. H. Jowett and Hilda M. Davies. Practical experimentation as a teaching method in statistics. Journal of the Royal Statistical Society. Series A (General), 123(1):10–36, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343185.

Bartlett:1960:CCS

[7] M. S. Bartlett. The critical community size for measles in the United States. Journal of the Royal Statistical Society. Series A (General), 123(1):37–44, ???? 1960. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343186.

Gridgeman:1960:CME

[8] N. T. Gridgeman. Card-matching experiments: A conspectus of theory. Journal of the Royal Statistical Society. Series A (General), 123(1):45–49, ???? 1960. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343187.

Hart:1960:BCU

[9] P. E. Hart. Business concentration in the United Kingdom. Journal of the Royal Statistical Society. Series A (General), 123(1):50–58, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343188.

Garwood:1960:CDR

[10] F. Garwood and Grace O. Jeffcoate. Comparative death-rates per person-mile associated with various forms of transport, 1952–57. Journal of the Royal Statistical Society. Series A (General), 123(1):59–61, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343189.

Myatt-Price:1960:TA

[11] Evelyn M. Myatt-Price. A tally of ale. Journal of the Royal Statistical Society. Series A (General), 123(1):62–67, ???? 1960. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343190.

Silcock:1960:BRB

[12] H. Silcock. Book review: Statistics, by A. R. Ilersic. Journal of the Royal Statistical Society. Series A (General), 123(1):68–70, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343191.

Jessop:1960:BRB

[13] W. N. Jessop. Book review: Analysis of Straight-Line Data, by Forman S. Acton. Journal of the Royal Statistical Society. Series A (General), 123(1):70, ???? 1960. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343192.

Vajda:1960:BRBa

[14] S. Vajda. Book review: Cybernetics and Management, by Stafford Beer. Journal of the Royal Statistical Society. Series A (General), 123(1):70–71, ???? 1960. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2343193.

Vajda:1960:BRBb

[15] S. Vajda. Book review: Operations Research: An Annotated Bibliography, by James H. Bachelor. Journal of the Royal Statistical Society. Series A (General), 123(1):71, ???? 1960. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343194.

Morton:1960:BRBa

[16] G. Morton. Book review: Dynamic Management Decision Games, including Seven Noncomputer Games, by Jay R. Greene; Roger L. Sisson. Journal of the Royal Statistical Society. Series A (General), 123(1):71–72, ???? 1960. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343195.

Keen:1960:BRB

[17] J. E. Keen. Book review: Sampling Inspection Tables: Single and Double Sampling, by H. F. Dodge; H. G. Romig. Journal of the Royal Statistical Society. Series A (General), 123(1):72– 73, ???? 1960. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2343196.

Tocher:1960:BRBa

[18] K. D. Tocher. Book review: Tables of Osculatory Interpolation Coefficients, by H. E. Salzer. Journal of the Royal Statistical Society. Series A (General), 123(1):73, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343197.

Carrier:1960:BRBa

[19] N. H. Carrier. Book review: The Registrar-General’s Statistical Review of England and Wales for the Year 1957, Part III, Commentary. Journal of the Royal Statistical Society. Series A (General), 123(1):73–74, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343198.

Florence:1960:BRB

[20] P. Sargant Florence. Book review: Studies in Company Finance: A Symposium on the Economic Analysis and Interpretation of British Company Accounts, by Brian Tew; R. F. Henderson. Journal of the Royal Statistical Society. Series A (General), 123(1):74–75, ???? 1960. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343199.

Forsyth:1960:BRB

[21] F. G. Forsyth. Book review: Price Theory, by W. J. L. Ryan. Journal of the Royal Statistical Society. Series A (General), 123(1):75–77, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343200.

Ainsworth:1960:BRB

[22] R. B. Ainsworth. Book review: The International Standardisation of Labour Statistics. Journal of the Royal Statistical Society. Series A (General), 123 (1):77–78, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343201.

Middleton:1960:BRB

[23] B. F. Middleton. Book review: An Economic Survey of the Sisal Industry of Tanganyika, by C. W. Guillebaud. Journal of the Royal Statistical Society. Series A (General), 123(1):78–79, ???? 1960. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2343202.

Anonymous:1960:CNa

[24] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 123(1):80–85, ???? 1960. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343203.

Ainsworth:1960:ECR

[25] R. B. Ainsworth and J. W. Nixon. Edmund Cecil Ramsbottom, 1881–1959. Journal of the Royal Statistical Society. Series A (General), 123(1):86–87, ???? 1960. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2343204.

Hemelrijk:1960:DVD

[26] J. Hemelrijk. David van Dantzig, 1900– 1959. Journal of the Royal Statistical Society. Series A (General), 123(1):87– 88, ???? 1960. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2343205.

Anonymous:1960:SEAa

[27] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 123(1):89–98, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343206.

Anonymous:1960:LALa

[28] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 123 (1):99–106, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343207.

Glover:1960:STG

[29] K. F. Glover. Statistics of the transport of goods by road. Journal of the Royal Statistical Society. Series A (General), 123(2):107–139, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342652.

Strauss:1960:SEM

[30] E. Strauss. The structure of the English milk industry. Journal of the Royal Statistical Society. Series A (General), 123 (2):140–173, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342653.

Glass:1960:HIA

[31] D. V. Glass. Human infertility and artificial insemination: The demographic background. Journal of the Royal Statistical Society. Series A (General), 123 (2):174–181, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342654.

Hajnal:1960:AIF

[32] John Hajnal. Artificial insemination and the frequency of incestuous marriages. Journal of the Royal Statistical Society. Series A (General), 123(2):

182–194, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342655.

Matuszewski:1960:SIO

[33] T. I. Matuszewski. On some input– output computations. Journal of the Royal Statistical Society. Series A (General), 123(2):195–199, ???? 1960. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342656.

Stuart:1960:BRB

[34] A. Stuart. Book review: Testing Statistical Hypotheses, by E. L. Lehmann. Journal of the Royal Statistical Society. Series A (General), 123(2):200– 201, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342657.

Maurice:1960:BRB

[35] Rita J. Maurice. Book review: Statistics: An Introduction, by D. A. S. Fraser. Journal of the Royal Statistical Society. Series A (General), 123 (2):201–202, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342658.

Stone:1960:BRB

[36] M. Stone. Book review: Studies in Mathematical Learning Theory, by Robert R. Bush; William K. Estes. Journal of the Royal Statistical Society. Series A (General), 123(2):202– 203, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342659.

Muir:1960:BRB

[37] A. Muir. Book review: Probability and Statistical Inference for Engineers: A First Course, by Cyrus Derman; Merton Klein. Journal of the Royal Statistical Society. Series A (General), 123 (2):203, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342660.

Corlett:1960:BRB

[38] W. J. Corlett. Book review: Social Accounting and Economic Models, by Richard Stone; Giovanna Croft-Murray. Journal of the Royal Statistical Society. Series A (General), 123 (2):204, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342661.

Hazlewood:1960:BRB

[39] A. Hazlewood. Book review: The Comparative Study of Economic Growth and Structure: Suggestions on Research Objectives and Organization. Journal of the Royal Statistical Society. Series A (General), 123(2):205–206, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342662.

Deane:1960:BRBa

[40] Marjorie Deane. Book review: Business Mergers and Take-Over Bids: A Study of the Post-War Pattern of Amalgamation and Reconstruction of Companies, by Ronald W. Moon. Journal of the Royal Statistical Society. Series A (General), 123(2):206, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342663.

Maton:1960:BRB

[41] J. M. Maton. Book review: The Compilation of Manufacturing Statistics, by Frank A. Hanna. Journal of the Royal Statistical Society. Series A (General), 123(2):206–207, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2342664.

Rimmer:1960:BRB

[42] G. Rimmer. Book review: Fluctuations in Trade, Industry, and Finance: A Study of British Economic Development, 1850–1860, by J. R. T. Hughes. Journal of the Royal Statistical Society. Series A (General), 123(2):207– 209, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342665.

Raeburn:1960:BRB

[43] J. R. Raeburn. Book review: Scottish Agricultural Economics: Some Studies of Current Economic Conditions in Scottish Farming, by Department of Agriculture for Scotland. Journal of the Royal Statistical Society. Series A (General), 123(2):209, ???? 1960. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342666.

George:1960:BRB

[44] R. F. George. Book review: The Public and the Programme: An Audience Research Report on Listeners and Viewers, the Time They Devote to Listening and Viewing, the Services They Patronise, their Selectiveness and their Tastes. Journal of the Royal Statistical Society. Series A (General), 123 (2):209–211, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342667.

Benjamin:1960:BRBa

[45] B. Benjamin. Book review: Demography, by P. R. Cox. Journal of the Royal Statistical Society. Series A (General), 123(2):211–212, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342668.

Benjamin:1960:BRBb

[46] B. Benjamin. Book review: China’s Population: Census and Vital Statistics, by S. Chandrasekhar. Journal of the Royal Statistical Society. Series A (General), 123(2):213, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342669.

Carrier:1960:BRBb

[47] N. H. Carrier. Book review: The Registrar-General’s Statistical Review of England and Wales for the Year 1958, Part II, Tables, Population. Journal of the Royal Statistical Society. Series A (General), 123(2):213– 214, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342670.

Douglas:1960:BRB

[48] J. W. B. Douglas. Book review: Growing Up in Newcastle upon Tyne: A Continuing Study of Health and Illness in Young Children within their Families, by F. J. W. Miller; S. D. M. Court; W. S. Walton; E. G. Knox. Journal of the Royal Statistical Society. Series A (General), 123(2):214–215, ???? 1960. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342671.

Curwen:1960:BRB

[49] M. P. Curwen. Book review: The Bethlem Royal Hospital and the Maudsley Hospital Triennial Statistical Report, Years 1955–1957, by E. H. Hare. Journal of the Royal Statistical Society. Series A (General), 123(2):215, ???? 1960. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342672.

Anonymous:1960:CNb

[50] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 123(2):216–220, ???? 1960. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342673.

Anonymous:1960:CRS

[51] Anonymous. Corrigendum: Reviews of statistical and economic books. Journal of the Royal Statistical Society. Series A (General), 123(2):220, ???? 1960. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342674.

Smith:1960:PMG

[52] Cedric A. B. Smith. Patrick Michael Grundy, 1917–1959. Journal of the Royal Statistical Society. Series A (General), 123(2):221–222, ???? 1960. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342675.

Anonymous:1960:SEAb

[53] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 123(2):223–230, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342676.

Anonymous:1960:LALb

[54] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 123 (2):231–235, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342677.

Stone:1960:EBD

[55] P. A. Stone. The economics of building designs. Journal of the Royal Statistical Society. Series A (General), 123 (3):237–273, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342468.

Robinson:1960:PPE

[56] Joan Robinson, J. Downie, and C. B. Winsten. The present position of econometrics. Journal of the Royal Statistical Society. Series A (General), 123 (3):274–296, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342469.

Anscombe:1960:NSS

[57] F. J. Anscombe. Notes on sequential sampling plans. Journal of the Royal Statistical Society. Series A (General), 123(3):297–306, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342470.

Kemsley:1960:SEM

[58] W. F. F. Kemsley and J. L. Nicholson. Some experiments in methods of conducting family expenditure surveys.

Journal of the Royal Statistical Society. Series A (General), 123(3):307– 328, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342471.

Anonymous:1960:WP

[59] Anonymous. Wholesale prices in 1959. Journal of the Royal Statistical Society. Series A (General), 123(3):329– 335, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342472.

Duckworth:1960:BRB

[60] W. E. Duckworth. Book review: Rapid Statistical Calculations: A Collection of Distribution-Free and Easy Methods of Estimation and Testing, by M. H. Quenouille. Journal of the Royal Statistical Society. Series A (General), 123 (3):336–337, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342473.

Moore:1960:BRB

[61] P. G. Moore. Book review: Methods of Correlation and Regression Analysis, Linear and Curvilinear, by M. Ezekiel; K. A. Fox. Journal of the Royal Statistical Society. Series A (General), 123 (3):337–338, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342474.

Welch:1960:BRB

[62] B. L. Welch. Book review: Mathematical Statistics, by J. N. Kapur; H. C. Saxena. Journal of the Royal Statistical Society. Series A (General), 123 (3):338–339, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342475.

Hammersley:1960:BRBa

[63] J. M. Hammersley. Book review: Probability and Statistics: The Harald Cramer Volume, by Ulf Grenander. Journal of the Royal Statistical Society. Series A (General), 123(3):339, ???? 1960. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2342476.

David:1960:BRB

[64] F. N. David. Book review: Theory of Probability, by W. Burnside. Journal of the Royal Statistical Society. Series A (General), 123(3):339, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342477.

Patterson:1960:BRB

[65] H. D. Patterson. Book review: Modern Sampling Methods, by Palmer O. Johnson; Munamarty S. Rao. Journal of the Royal Statistical Society. Series A (General), 123(3):339–340, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342478.

Hammersley:1960:BRBb

[66] J. M. Hammersley. Book review: A Source Book in Mathematics, by David Eugene Smith. Journal of the Royal Statistical Society. Series A (General), 123(3):340–341, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342479.

Ross:1960:BRB

[67] Alan S. C. Ross. Book review: Type-Token Mathematics: A Textbook of Mathematical Linguistics, by G. Herdan. Journal of the Royal Statistical Society. Series A (General), 123(3): 341–342, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342480.

Edwards:1960:BRB

[68] J. H. Edwards. Book review: Introduction to Quantitative Genetics, by D. S. Falconer. Journal of the Royal Statistical Society. Series A (General), 123 (3):342–343, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342481.

Vajda:1960:BRBc

[69] S. Vajda. Book review: Mathematical Methods and Theory in Games, Programming, and Economics, by Samuel Karlin. Journal of the Royal Statistical Society. Series A (General), 123 (3):343–344, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342482.

Robinson:1960:BRB

[70] Joan Robinson. Book review: Income and Wealth, by Raymond Goldsmith; Christopher Saunders. Journal of the Royal Statistical Society. Series A (General), 123(3):344–346, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342483.

Berman:1960:BRB

[71] L. S. Berman. Book review: A System of National Accounts and Supporting Tables. Journal of the Royal Statistical Society. Series A (General), 123 (3):346–347, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342484.

Peacock:1960:BRB

[72] Alan T. Peacock. Book review: Domestic Income and Expenditure: A Description of Sources and Methods, with Revised Calculations from 1954 to 1958. Journal of the Royal Statistical Society. Series A (General), 123(3): 347–348, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342485.

Carter:1960:BRB

[73] C. F. Carter. Book review: Methodological Aspects of Statistics on Research and Development Costs and Manpower. Journal of the Royal Statistical Society. Series A (General), 123(3):348, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342486.

Davies:1960:BRB

[74] O. L. Davies. Book review: Statistical Forecasting for Inventory Control, by Robert G. Brown. Journal of the Royal Statistical Society. Series A (General), 123(3):348–349, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342487.

Heady:1960:BRB

[75] J. A. Heady. Book review: Medical Surveys and Clinical Trials, by L. J. Witts. Journal of the Royal Statistical Society. Series A (General), 123(3): 349–350, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342488.

Rosenbaum:1960:BRB

[76] S. Rosenbaum. Book review: Tabulating Equipment and Army Medical Statistics, by Albert G. Love; Eugene L. Hamilton; Ida Levin Hellman. Journal of the Royal Statistical Society. Series A (General), 123(3):350, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342489.

Anonymous:1960:CNc

[77] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 123(3):351–354, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342490.

White:1960:ECS

[78] G. R. White. Ernest Charles Snow, C.B.E., D.Sc., 1886–1959. Journal of the Royal Statistical Society. Series A (General), 123(3):355–356, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342491.

Anonymous:1960:SEAc

[79] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 123(3):357–362, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342492.

Anonymous:1960:LALc

[80] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 123 (3):363–366, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342493.

Anonymous:1960:VIb

[81] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 123(4):538– 541, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343270.

Forsyth:1960:RBF

[82] F. G. Forsyth. The relationship between family size and family expenditure. Journal of the Royal Statistical Society. Series A (General), 123 (4):367–397, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343271.

Champernowne:1960:EIR

[83] D. G. Champernowne. An experimental investigation of the robustness of certain procedures for estimating means and regression coefficients. Journal of the Royal Statistical Society. Series A (General), 123(4):398–412, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343272.

Benjamin:1960:SPC

[84] B. Benjamin. Statistical problems connected with the 1961 population census. Journal of the Royal Statistical Society. Series A (General), 123(4):413– 426, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343273.

Nicholson:1960:OPC

[85] R. J. Nicholson and S. Gupta. Output and productivity changes in British manufacturing industry, 1948–1954: A study from census of production data.

Journal of the Royal Statistical Society. Series A (General), 123(4):427– 459, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343274.

Anonymous:1960:RMS

[86] Anonymous. Regulations of the medical section. Journal of the Royal Statistical Society. Series A (General), 123 (4):460–461, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343275.

Anonymous:1960:ARC

[87] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 123 (4):462–480, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343276.

Anonymous:1960:POH

[88] Anonymous. Proceedings of the one hundred and twenty-sixth Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 123 (4):481, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343277.

Cox:1960:BRB

[89] D. R. Cox. Book review: The Analysis of Variance, by H. Scheffe. Journal of the Royal Statistical Society. Series A (General), 123(4):482–483, ???? 1960. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343278.

Kendall:1960:BRB

[90] D. G. Kendall. Book review: Characteristic Functions, by E. Lukacs. Journal of the Royal Statistical Society. Series A (General), 123(4):484, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343279.

Whittle:1960:BRB

[91] P. Whittle. Book review: An Introduction to Infinitely Many Variates, by E. A. Robinson. Journal of the Royal Statistical Society. Series A (General), 123(4):485, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343280.

Foster:1960:BRB

[92] F. G. Foster. Book review: The Theory of Storage, by P. A. P. Moran. Journal of the Royal Statistical Society. Series A (General), 123(4):485–486, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343281.

Cane:1960:BRB

[93] Violet Cane. Book review: Individual Choice Behavior: A Theoretical Analysis, by R. Duncan Luce. Journal of the Royal Statistical Society. Series A (General), 123(4):486–488, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343282.

Tocher:1960:BRBb

[94] K. D. Tocher. Book review: Basic Theorems in Matrix Theory, by Marvin Marcus. Journal of the Royal Statistical Society. Series A (General), 123 (4):488, ???? 1960. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343283.

Tocher:1960:BRBc

[95] K. D. Tocher. Book review: Tables of the Bivariate Normal Distribution Function and Related Functions. Journal of the Royal Statistical Society. Series A (General), 123(4):488, ???? 1960. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343284.

Tocher:1960:BRBd

[96] K. D. Tocher. Book review: Confluent Hypergeometric Functions, by L. J. Slater. Journal of the Royal Statistical Society. Series A (General), 123(4): 489, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343285.

Bodmer:1960:BRB

[97] W. F. Bodmer. Book review: Biometrical Genetics. Proceedings of an International Symposium Held at Ottawa, August 1958, by Oscar Kempthorne. Journal of the Royal Statistical Society. Series A (General), 123(4):489– 491, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343286.

Spicer:1960:BRB

[98] C. C. Spicer. Book review: Catalytic Models in Epidemiology, by Hugo Muench. Journal of the Royal Statistical Society. Series A (General), 123 (4):491, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343287.

Benjamin:1960:BRBc

[99] B. Benjamin. Book review: The Study of Population: An Inventory and Appraisal, by Philip M. Hauser; Otis Dudley Duncan. Journal of the Royal Statistical Society. Series A (General), 123 (4):492, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343288.

Morton:1960:BRBb

[100] G. Morton. Book review: An Introduction to Linear Programming and the Theory of Games, by S. Vajda. Journal of the Royal Statistical Society. Series A (General), 123(4):493, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343289.

Johnston:1960:BRB

[101] J. Johnston. Book review: Impact Multipliers and Dynamic Properties of the Klein-Goldberger Model, by Arthur S. Goldberger. Journal of the Royal Statistical Society. Series A (General), 123 (4):493–494, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343290.

Gorman:1960:BRB

[102] W. M. Gorman. Book review: Econometrics: An Introduction to Maximum Likelihood Methods, by Stefan Valavanis. Journal of the Royal Statistical Society. Series A (General), 123(4):494– 496, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343291.

Morton:1960:BRBc

[103] G. Morton. Book review: Management Technology, by Roger R. Crane; C.

West Churchman. Journal of the Royal Statistical Society. Series A (General), 123(4):496–497, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343292.

Barna:1960:BRB

[104] T. Barna. Book review: A Comparison of National Output and Productivity of the United Kingdom and the United States, by D. Paige; G. Bombach. Journal of the Royal Statistical Society. Series A (General), 123(4): 497, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343293.

Williams:1960:BRB

[105] Alan Williams. Book review: Public Finance in Theory and Practice, by A. R. Prest. Journal of the Royal Statistical Society. Series A (General), 123 (4):498–499, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343294.

Nicholson:1960:BRB

[106] J. L. Nicholson. Book review: Bibliography on Income and Wealth, Volume VII, 1955–56, by Phyllis Deane. Journal of the Royal Statistical Society. Series A (General), 123(4):499, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343295.

Paine:1960:BRBa

[107] G. Paine. Book review: Changes in the Share of Wealth Held, by Top Wealth-Holders, 1922–1956. by Robert J. Lampman. Journal of the Royal Statistical Society. Series A (General), 123 (4):500, ???? 1960. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2343296.

Deane:1960:BRBb

[108] Marjorie Deane. Book review: Concentration in British Industry, by Richard Evely; I. M. D. Little. Journal of the Royal Statistical Society. Series A (General), 123(4):500–501, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343297.

Deane:1960:BRBc

[109] Marjorie Deane. Book review: Merger Movements in American Industry, 1895–1956, by Ralph L. Nelson. Journal of the Royal Statistical Society. Series A (General), 123(4):501–502, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343298.

Paine:1960:BRBb

[110] G. Paine. Book review: State Income Differentials 1919–1954, by Frank A. Hanna. Journal of the Royal Statistical Society. Series A (General), 123 (4):502, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343299.

Maizels:1960:BRB

[111] A. Maizels. Book review: Trade Balances during Business Cycles: U.S. and Britain since 1880, by Ilse Mintz. Journal of the Royal Statistical Society. Series A (General), 123(4):503– 504, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343300.

Mannheim:1960:BRB

[112] Hermann Mannheim. Book review: Time Spent Awaiting Trial. A Home Office Research Unit Report, by Evelyn Gibson. Journal of the Royal Statistical Society. Series A (General), 123 (4):504–506, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343301.

Anstey:1960:BRB

[113] Vera Anstey. Book review: The City of Calcutta: A Socio-Economic Survey, 1954–55 to 1957–58, by S. N. Sen. Journal of the Royal Statistical Society. Series A (General), 123(4):506– 508, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343302.

Anonymous:1960:CNd

[114] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 123(4):509–517, ???? 1960. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343303.

Sagoroff:1960:OA

[115] S. Sagoroff and L. Isserlis. Oskar Anderson, 1887–1960. Journal of the Royal Statistical Society. Series A (General), 123(4):518–519, ???? 1960. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343304.

Roy:1960:GD

[116] R. Roy. Georges Darmois, 1888–1960. Journal of the Royal Statistical Society. Series A (General), 123(4):520– 522, ???? 1960. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2343305.

Anonymous:1960:SEAd

[117] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 123(4):523–530, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343306.

Anonymous:1960:LALd

[118] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 123 (4):531–537, ???? 1960. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343307.

Anonymous:1960:BM

[119] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 123(4):i, ???? 1960. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343308.

Anonymous:1961:FMa

[120] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 124(1):??, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343148.

Anonymous:1961:PA

[121] Anonymous. Presidential address. Journal of the Royal Statistical Society. Series A (General), 124(1):1–16, ???? 1961. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343149.

Anonymous:1961:PM

[122] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society. Series A (General), 124(1):16– 19, ???? 1961. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2343150.

Daw:1961:CMF

[123] R. H. Daw. The comparison of male and female mortality rates. Journal of the Royal Statistical Society. Series A (General), 124(1):20–43, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343151.

Cox:1961:DEC

[124] D. R. Cox. Design of experiments: The control of error. Journal of the Royal Statistical Society. Series A (General), 124(1):44–48, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343152.

Maxwell:1961:RTF

[125] A. E. Maxwell. Recent trends in factor analysis. Journal of the Royal Statistical Society. Series A (General), 124 (1):49–59, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343153.

Donnison:1961:MHE

[126] D. V. Donnison. The movement of households in England. Journal of the Royal Statistical Society. Series A (General), 124(1):60–80, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/

stable/2343154. See corrigendum [180].

Kevork:1961:ETW

[127] C. Kevork. The elimination of tariffs in a wide free trade area in Europe: An analysis for selected commodities. Journal of the Royal Statistical Society. Series A (General), 124(1):81–92, ???? 1961. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343155.

Vajda:1961:BRBa

[128] S. Vajda. Book review: Mathematical Methods in the Theory of Queueing, by A. Y. Kintchine; D. M. Andrews; M. H. Quenouille. Journal of the Royal Statistical Society. Series A (General), 124(1):93–94, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343156.

Moran:1961:BRB

[129] P. A. Moran. Book review: Elements of the Theory of Markov Processes and their Applications, by A. T. Bharucha-Reid. Journal of the Royal Statistical Society. Series A (General), 124(1):94– 95, ???? 1961. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343157.

Whittle:1961:BRB

[130] P. Whittle. Book review: Introduction to Probability and Random Variables, by George P. Wadsworth; Joseph G. Bryan. Journal of the Royal Statistical Society. Series A (General), 124(1):95– 96, ???? 1961. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343158.

Beale:1961:BRB

[131] E. M. L. Beale. Book review: Information and Decision Processes, by Robert E. Machol. Journal of the Royal Statistical Society. Series A (General), 124 (1):96–97, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343159.

Priestley:1961:BRB

[132] M. B. Priestley. Book review: Probability: An Introduction, by Samuel Goldberg. Journal of the Royal Statistical Society. Series A (General), 124(1):97, ???? 1961. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343160.

Newell:1961:BRB

[133] D. J. Newell. Book review: Introduction to Probability and Statistics, by Henry L. Alder; Edward B. Roessler. Journal of the Royal Statistical Society. Series A (General), 124(1):97–98, ???? 1961. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343161.

Wetherill:1961:BRB

[134] G. B. Wetherill. Book review: An Introduction to Experimental Design, by William S. Ray. Journal of the Royal Statistical Society. Series A (General), 124(1):98–99, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343162.

Hill:1961:BRB

[135] I. D. Hill. Book review: Statistics Manual with Examples taken from Ordnance Development, by Edwin L. Crow; Frances A. Davis; Margaret W. Maxfield. Journal of the Royal Statistical Society. Series A (General), 124(1):99– 100, ???? 1961. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2343163.

Page:1961:BRBa

[136] E. S. Page. Book review: Annual Review in Automatic Programming, Volume I, by Richard Goodman. Journal of the Royal Statistical Society. Series A (General), 124(1):100–101, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343164.

Benjamin:1961:BRBa

[137] B. Benjamin. Book review: A History of the International Statistical Institute, 1885–1960, by J. W. Nixon. Journal of the Royal Statistical Society. Series A (General), 124(1):102, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343165.

Parkinson:1961:BRB

[138] C. Northcote Parkinson. Book review: Administrative Vitality: The Conflict with Bureaucracy, by Marshall E. Dimock. Journal of the Royal Statistical Society. Series A (General), 124(1): 102–104, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343166.

Tippett:1961:BRB

[139] L. H. C. Tippett. Book review: Decimal Coinage and the Metric System. Should Britain Change?, by The British Association for the Advancement of Science; The Association of British Chambers of Commerce. Journal of the Royal Statistical Society. Series A (General), 124(1):104–105, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343167.

Oldham:1961:BRB

[140] P. D. Oldham. Book review: Controlled Clinical Trials. Journal of the Royal Statistical Society. Series A (General), 124(1):105–106, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343168.

Benjamin:1961:BRBb

[141] B. Benjamin. Book review: Introduction to Health Statistics for the Use of Health Officers, Students, Public Health and Social Workers, etc, by Satya Swaroop. Journal of the Royal Statistical Society. Series A (General), 124(1):106–107, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343169.

Hewitt:1961:BRB

[142] D. Hewitt. Book review: Report on Confidential Enquiries into Maternal Deaths in England and Wales, 1955– 1957, by A. L. Walker; A. J. Wrigley; G. S. W. Organe; Roma N. Chamberlain; W. J. Martin. Journal of the Royal Statistical Society. Series A (General), 124(1):107, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343170.

Gore:1961:BRB

[143] A. T. Gore. Book review: Morbidity Statistics from General Practice, Volume II (Occupation), by W. P. D. Logan. Journal of the Royal Statistical Society. Series A (General), 124(1): 108, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343171.

Grebenik:1961:BRB

[144] E. Grebenik. Book review: Demographic Yearbook, 1959. Journal of the Royal Statistical Society. Series A (General), 124(1):108–109, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343172.

Morant:1961:BRB

[145] G. M. Morant. Book review: Race Elements in Bengal: A Quantitative Study, by D. N. Majumdar; C. Radhakrishna Rao; P. C. Mahalanobis. Journal of the Royal Statistical Society. Series A (General), 124(1):109– 110, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343173.

Hocking:1961:BRBa

[146] W. S. Hocking. Book review: National Insurance Acts, 1946–59. Report, by the Government Actuary on the Second Quinquennial Review (H.C. 220, 1959–60). Journal of the Royal Statistical Society. Series A (General), 124 (1):110–113, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343174.

Hocking:1961:BRBb

[147] W. S. Hocking. Book review: National Insurance (Industrial Injuries) Acts, 1946–59. Report, by the Government Actuary on the Second Quinquennial Review (H.C. 300, 1959–60). Journal of the Royal Statistical Society. Series A (General), 124(1):113, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343175.

Dicks-Mireaux:1961:BRB

[148] L. A. Dicks-Mireaux. Book review: Regional Cycles of Manufacturing Employment in the United States, 1914– 1953, by George H. Borts. Journal of the Royal Statistical Society. Series A (General), 124(1):114, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343176.

Lomax:1961:BRBa

[149] K. S. Lomax. Book review: Productivity and Technical Change, by W. E. G. Salter. Journal of the Royal Statistical Society. Series A (General), 124 (1):114–115, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343177.

Lomax:1961:BRBb

[150] K. S. Lomax. Book review: Industrial Production, 1959 Revision, by Board of Governors of the Federal Reserve System. Journal of the Royal Statistical Society. Series A (General), 124 (1):115, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343178.

Anonymous:1961:CNa

[151] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 124(1):116–120, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343179.

Anonymous:1961:SEA

[152] Anonymous. Statistical and economic articles in recent periodicals and list of additions to the library. Journal of the Royal Statistical Society.

Series A (General), 124(1):120, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343180.

Anonymous:1961:BMa

[153] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 124(1):i–iv, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343181.

Anonymous:1961:FMb

[154] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 124(2):??, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984127.

Tocher:1961:RMO

[155] K. D. Tocher. The role of models in operational research. Journal of the Royal Statistical Society. Series A (General), 124(2):121–142, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2984128.

Scott:1961:RMS

[156] Christopher Scott. Research on Mail surveys. Journal of the Royal Statistical Society. Series A (General), 124 (2):143–205, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2984129.

Andic:1961:IDI

[157] Suphan Andic and Alan T. Peacock. The international distribution of income, 1949 and 1957. Journal of the Royal Statistical Society. Series A (General), 124(2):206–218, ????

1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984130.

Simpson:1961:ASD

[158] H. R. Simpson. The analysis of survey data on an electronic computer. Journal of the Royal Statistical Society. Series A (General), 124(2):219–226, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984131.

Kahan:1961:PDN

[159] B. C. Kahan. A practical demonstration of a needle experiment designed to give a number of concurrent estimates of π. Journal of the Royal Statistical Society. Series A (General), 124 (2):227–239, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2984132.

Bofinger:1961:NPW

[160] Eve Bofinger and V. J. Bofinger. A note on the paper by W. E. Thomson on “Ernie — A Mathematical and Statistical Analysis”. Journal of the Royal Statistical Society. Series A (General), 124 (2):240–243, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2984133. See

[2].

Borts:1961:RCR

[161] George Borts. The recent controversy over resale price maintenance. Journal of the Royal Statistical Society. Series A (General), 124(2):244–249, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984134.

David:1961:BRB

[162] F. N. David. Book review: Contributions to Probability and Statistics: Essays in Honor of Harold Hotelling, by I. Olkin; S. G. Ghurye; W. Hoeffding; W. G. Madow; H. B. Mann. Journal of the Royal Statistical Society. Series A (General), 124(2):250–251, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984135.

Welch:1961:BRB

[163] B. L. Welch. Book review: Principles of Regression Analysis, by R. L. Plackett. Journal of the Royal Statistical Society. Series A (General), 124 (2):251–252, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2984136.

Buckland:1961:BRBa

[164] W. R. Buckland. Book review: An Introduction to Stochastic Processes with Special Reference to Methods and Applications, by M. S. Bartlett. Journal of the Royal Statistical Society. Series A (General), 124(2):252, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984137.

Jenkins:1961:BRB

[165] G. M. Jenkins. Book review: Time-Series Analysis, by E. J. Hannan. Journal of the Royal Statistical Society. Series A (General), 124(2):252–253, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984138.

Kendall:1961:BRBa

[166] M. G. Kendall. Book review: Transvariazione, by Corrado Gini. Journal of the Royal Statistical Society. Series A (General), 124(2):253, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2984139.

Gorman:1961:BRB

[167] W. M. Gorman. Book review: “An Issue in Honor of Ragnar Frisch in the Year of his Sixty-Fifth Birthday.”. Journal of the Royal Statistical Society. Series A (General), 124(2):254– 255, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2984140.

Robinson:1961:BRB

[168] Joan Robinson. Book review: Statistical Cost Analysis, by J. Johnson. Journal of the Royal Statistical Society. Series A (General), 124(2):255, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2984141.

Nicholson:1961:BRB

[169] J. L. Nicholson. Book review: Collecting Financial Data, by Consumer Panel Techniques: A Pilot Study. by Robert Ferber. Journal of the Royal Statistical Society. Series A (General), 124 (2):255–256, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2984142.

Saunders:1961:BRB

[170] C. T. Saunders. Book review: The Quality and Economic Significance of Anticipations Data. Journal of the Royal Statistical Society. Series A (General), 124(2):256–258, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984143.

Roy:1961:BRB

[171] A. D. Roy. Book review: Investment Decisions Under Uncertainty, by R. A. D. Egerton. Journal of the Royal Statistical Society. Series A (General), 124 (2):258–259, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2984144.

Vajda:1961:BRBb

[172] S. Vajda. Book review: Self-Organizing Systems: Proceedings of an Inter-Disciplinary Conference, 5 and 6 May, 1959; Co-Sponsored, by Information Systems Branch of Office of Naval Research and Armour Research Foundation of Illinois Institute of Technology. by Marshall C. Yovitz; Scott Cameron. Journal of the Royal Statistical Society. Series A (General), 124(2):259– 260, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2984145.

Vajda:1961:BRBc

[173] S. Vajda. Book review: Management Sciences, Models and Techniques, by C. West Churchman; M. Verhulst. Journal of the Royal Statistical Society. Series A (General), 124(2):260–261, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984146.

Page:1961:BRBb

[174] D. J. Page. Book review: The Statistical Basis of Quality Control Charts: A Manual for Business and Factory Managers, by S. K. Ekambaram. Journal of the Royal Statistical Society. Series A (General), 124(2):262, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984147.

Nove:1961:BRB

[175] A. Nove. Book review: Industrial Statistics, by A. L. Ezhov. Journal of the Royal Statistical Society. Series A (General), 124(2):262–263, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984148.

Deane:1961:BRB

[176] P. M. Deane. Book review: Historical Statistics of the United States, Colonial Times to 1957: A Statistical Abstract Supplement. Journal of the Royal Statistical Society. Series A (General), 124 (2):263–264, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2984149.

Kelsall:1961:BRBa

[177] R. K. Kelsall. Book review: Newcomers: The West Indians in London, by Ruth Glass. Journal of the Royal Statistical Society. Series A (General), 124 (2):264–265, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2984150.

Ward:1961:BRB

[178] Barbara E. Ward. Book review: Upper Nankin Street, Singapore, by Barrington Kaye. Journal of the Royal Statistical Society. Series A (General), 124 (2):265–267, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2984151.

Wootton:1961:BRB

[179] Barbara Wootton. Book review: Tradition and Change: A Study of Banbury, by Margaret Stacey. Journal of the Royal Statistical Society. Series A (General), 124(2):267–268, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984152.

Anonymous:1961:CMH

[180] Anonymous. Corrigendum: The movement of households in England. Journal of the Royal Statistical Society. Series A (General), 124(2):268, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984153. See [126].

Anonymous:1961:CNb

[181] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 124(2):269–274, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984154.

Irwin:1961:ECF

[182] J. O. Irwin and E. D. Van Rest. Edgar Charles Fieller, 1907–1960. Journal of the Royal Statistical Society. Series A (General), 124(2):275–277, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2984155.

Anonymous:1961:LBRa

[183] Anonymous. List of books received for review. Journal of the Royal Statistical Society. Series A (General), 124 (2):278–283, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2984156.

Anonymous:1961:BMb

[184] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 124(2):i–viii, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2984157.

Anonymous:1961:FMc

[185] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 124(3):??, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343239.

Lindley:1961:EME

[186] D. V. Lindley. An experiment in the marking of an examination. Journal of the Royal Statistical Society. Series A (General), 124(3):285–313, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343240.

Drion:1961:TPN

[187] E. F. Drion. Two papers on nutrition: The intercorrelations between the nutrients consumed by a group of families in the Netherlands. Journal of the Royal Statistical Society. Series A (General), 124(3):314–335, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343241.

Heady:1961:DBC

[188] J. A. Heady. Diets of bank clerks development of a method of classifying the diets of individuals for use in epidemiological studies. Journal of the Royal Statistical Society. Series A (General), 124(3):336–371, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343242.

Johnson:1961:SAS

[189] N. L. Johnson. Sequential analysis: A survey. Journal of the Royal Statistical Society. Series A (General), 124 (3):372–411, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343243.

Friedlander:1961:TEC

[190] D. Friedlander. A technique for estimating a contingency table, given the marginal totals and some supplementary data. Journal of the Royal Statistical Society. Series A (General), 124 (3):412–420, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343244.

Ilersic:1961:SV

[191] A. R. Ilersic. Statistical Victoriana. Journal of the Royal Statistical Society. Series A (General), 124(3):421– 427, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343245.

Anonymous:1961:WP

[192] Anonymous. Wholesale prices in 1960. Journal of the Royal Statistical Society. Series A (General), 124(3):428– 434, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343246.

Wishart:1961:BRB

[193] D. M. G. Wishart. Book review: Modern Probability Theory and its Applications, by Emanuel Parzen. Journal of the Royal Statistical Society. Series A (General), 124(3):435–436, ????

1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343247.

Harris:1961:BRB

[194] C. C. Harris. Book review: Small Particle Statistics — An Account of Statistical Methods for the Investigation of Finely Divided Materials, by G. Herdan; M. L. Smith; W. H. Hardwick; P. Connor. Journal of the Royal Statistical Society. Series A (General), 124 (3):436–439, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343248.

Ehrenberg:1961:BRB

[195] A. S. C. Ehrenberg. Book review: Decisions, Values and Groups, by Dorothy Willner. Journal of the Royal Statistical Society. Series A (General), 124 (3):439–440, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343249.

Buckland:1961:BRBb

[196] W. R. Buckland. Book review: Exponentially Distributed Random Numbers, by Charles E. Clark; Betty W. Holz. Journal of the Royal Statistical Society. Series A (General), 124(3): 440, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343250.

Jessop:1961:BRB

[197] W. N. Jessop. Book review: Operations Research and Systems Engineering, by Charles D. Flagle; William H. Huggins; Robert H. Roy. Journal of the Royal Statistical Society. Series A (General), 124(3):440–441, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343251.

Vajda:1961:BRBd

[198] S. Vajda. Book review: Proceedings of the Second International Conference on Operational Research, by International Federation of Operational Research Societies. Journal of the Royal Statistical Society. Series A (General), 124 (3):441–442, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343252.

Brookes:1961:BRB

[199] B. C. Brookes. Book review: A Second Course in Statistics, by Robert Loveday. Journal of the Royal Statistical Society. Series A (General), 124(3):442– 443, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343253.

Peston:1961:BRB

[200] M. H. Peston. Book review: The Use of Economic Statistics, by Conrad Blyth. Journal of the Royal Statistical Society. Series A (General), 124(3):443– 444, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343254.

Peacock:1961:BRB

[201] Alan T. Peacock. Book review: The Structure of Property Ownership in Great Britain, by E. Victor Morgan. Journal of the Royal Statistical Society. Series A (General), 124(3):444– 445, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343255.

Reynolds:1961:BRB

[202] D. J. Reynolds. Book review: Capital in Transportation, Communications and Public Utilities: Its Formation and Financing, by Melville J. Ulmer. Journal of the Royal Statistical Society. Series A (General), 124(3):445–446, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343256.

Hawtrey:1961:BRB

[203] Ralph Hawtrey. Book review: Topical Comment: Essays in Dynamic Economics Applied, by Roy Harrod. Journal of the Royal Statistical Society. Series A (General), 124(3):446–447, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343257.

Paish:1961:BRB

[204] F. W. Paish. Book review: Money Under Review, by W. Manning Dacey. Journal of the Royal Statistical Society. Series A (General), 124(3):447– 448, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343258.

Corlett:1961:BRB

[205] W. J. Corlett. Book review: The Social Framework: An Introduction to Economics, by J. R. Hicks. Journal of the Royal Statistical Society. Series A (General), 124(3):449, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343259.

Ainsworth:1961:BRB

[206] R. B. Ainsworth. Book review: Carroll Wright and Labor Reform: The Origin of Labor Statistics, by James Leiby. Journal of the Royal Statistical Society. Series A (General), 124(3): 449–450, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343260.

Hazlewood:1961:BRB

[207] Arthur Hazlewood. Book review: Employment in an Underdeveloped Area: A Sample Survey of Kingston, Jamaica, by W. F. Maunder. Journal of the Royal Statistical Society. Series A (General), 124(3):450–451, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343261.

Benjamin:1961:BRBc

[208] B. Benjamin. Book review: The Institute of Actuaries Students’ Society: The First Fifty Years, 1910–1960 — An Appreciation, by F. A. A. Menzler. Journal of the Royal Statistical Society. Series A (General), 124(3):451, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343262.

Lafitte:1961:BRB

[209] F. Lafitte. Book review: Housing Needs and Planning Policy: A Restatement of the Problems of Housing Need and “Overspill” in England and Wales, by J. B. Cullingworth. Journal of the Royal Statistical Society. Series A (General), 124(3):451–452, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343263.

Kelsall:1961:BRBb

[210] R. K. Kelsall. Book review: Charles Booth, Social Scientist, by T. S. Simey; M. B. Simey. Journal of the Royal Statistical Society. Series A (General), 124 (3):453–454, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343264.

Morris:1961:BRB

[211] Terence Morris. Book review: Delinquent Generations, by L. T. Wilkins. Journal of the Royal Statistical Society. Series A (General), 124(3):454, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343265.

Anonymous:1961:CNc

[212] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 124(3):455–458, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343266.

Renyi:1961:CJ

[213] A. Renyi. Charles Jordan, 1871–1959. Journal of the Royal Statistical Society. Series A (General), 124(3):459, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343267.

Anonymous:1961:LBRb

[214] Anonymous. List of books received for review. Journal of the Royal Statistical Society. Series A (General), 124 (3):460–462, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343268.

Anonymous:1961:BMc

[215] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 124(3):i–vi, ????

1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343269.

Anonymous:1961:VI

[216] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 124(4):600– 603, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342928.

Anonymous:1961:FMd

[217] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 124(4):??, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342929.

Sukhatme:1961:WHF

[218] P. V. Sukhatme. The world’s hunger and future needs in food supplies. Journal of the Royal Statistical Society. Series A (General), 124(4):463–525, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342930.

Nixon:1961:CPE

[219] J. W. Nixon. Classification of the population by economic activities. Journal of the Royal Statistical Society. Series A (General), 124(4):526–542, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342931.

Prais:1961:FCE

[220] S. J. Prais. Family composition and expenditure: A comment on Mr Forsyth’s paper. Journal of the Royal Statistical Society. Series A (General), 124(4):

543–545, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342932.

Anonymous:1961:BNR

[221] Anonymous. Babbage’s note respecting the origin of the Statistical Society. Journal of the Royal Statistical Society. Series A (General), 124(4):546, ???? 1961. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2342933.

Anonymous:1961:RGA

[222] Anonymous. Regulations of the general applications section and the industrial applications section. Journal of the Royal Statistical Society. Series A (General), 124(4):547–550, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342934.

Anonymous:1961:ARC

[223] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 124 (4):551–571, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342935.

Anonymous:1961:POH

[224] Anonymous. Proceedings of the one hundred and twenty-seventh Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 124 (4):572, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342936.

Moore:1961:BRB

[225] P. G. Moore. Book review: Statistical Theory and Methodology in Science and Engineering, by K. A. Brownlee. Journal of the Royal Statistical Society. Series A (General), 124(4):573–574, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342937.

Plackett:1961:BRB

[226] R. L. Plackett. Book review: An Introduction to Linear Statistical Models. Vol. 1, by Franklin A. Graybill. Journal of the Royal Statistical Society. Series A (General), 124(4):574–575, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342938.

Kendall:1961:BRBb

[227] M. G. Kendall. Book review: Le Medie, by Corrado Gini. Journal of the Royal Statistical Society. Series A (General), 124(4):575–576, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342939.

Moser:1961:BRB

[228] C. A. Moser. Book review: Sampling Methods for Censuses and Surveys, by Frank Yates. Journal of the Royal Statistical Society. Series A (General), 124 (4):576–577, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342940.

Nixon:1961:BRBa

[229] J. W. Nixon. Book review: Stichproben in der Amtlichen Statistik. Journal of the Royal Statistical Society. Series A (General), 124(4):577, ????

1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342941.

Nixon:1961:BRBb

[230] J. W. Nixon. Book review: Lehrbuch der Praktischen Statistik. Bevolkerungs-Wirtschafts-Sozialstatistik, by P. Quante. Journal of the Royal Statistical Society. Series A (General), 124(4):577– 578, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342942.

Good:1961:BRB

[231] I. J. Good. Book review: Entropy and the Unity of Knowledge, by P. T. Landsberg. Journal of the Royal Statistical Society. Series A (General), 124 (4):578, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342943.

Herdan:1961:BRB

[232] G. Herdan. Book review: Contributions to the Statistical Study of the Codex Sinaiticus, by Christian Tindall. Journal of the Royal Statistical Society. Series A (General), 124(4):578– 579, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342944.

Tocher:1961:BRBa

[233] K. D. Tocher. Book review: Tables of Random Normal Deviates, by J. Sengupta; N. Bhattacharya. Journal of the Royal Statistical Society. Series A (General), 124(4):579–580, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342945.

Tocher:1961:BRBb

[234] K. D. Tocher. Book review: Tables of Tolerance-Limit Factors for Normal Distributions, by Alfred Weissberg; Glenn H. Beatty. Journal of the Royal Statistical Society. Series A (General), 124(4):580, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342946.

Tocher:1961:BRBc

[235] K. D. Tocher. Book review: Tables of ln Γ(Z) for Complex Argument, by A. A. Abramov; Tables of Airy Functions and Special Confluent Hypergeometric Functions for Asymptotic Solutions of Differential Equations of the Second Order, by A. D. Smirnov; D. G. Fry. Journal of the Royal Statistical Society. Series A (General), 124 (4):580–581, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342947.

Tocher:1961:BRBd

[236] K. D. Tocher. Book review: Tables of Values of the Function w(z) = e^{−z^{2}}(1+ \frac{2i}{\sqrt{π}} ∫_{0}^{z} e^{t^{2}} dt) for Complex Argument, by V. N. Faddeyeva; N. M. Terent’ev. Journal of the Royal Statistical Society. Series A (General), 124(4):581, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342948.

Anstey:1961:BRB

[237] Vera Anstey. Book review: The Environs of Tagore: A Preliminary Assessment for a Pilot Project in Regional Survey Planning Around the Visva-Bharati, by Hashim Amir Ali; Then and Now (1933–1958): A Study of Socio-Economic Structure and Change in Some Villages near Visva-Bharati University, Bengal, by Hashim Amir Ali. Journal of the Royal Statistical Society. Series A (General), 124(4):581– 583, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342949.

Cockburn:1961:BRB

[238] Christine Cockburn. Book review: The Cost of Social Security, 1949–1957. Journal of the Royal Statistical Society. Series A (General), 124(4):583– 584, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342950.

Cottrell:1961:BRB

[239] H. C. Cottrell. Book review: Statistical Measures of Corporate Bond Financing since 1900, by W. Braddock Hickman; Elizabeth T. Simpson. Journal of the Royal Statistical Society. Series A (General), 124(4):584, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342951.

Tinbergen:1961:BRB

[240] J. Tinbergen. Book review: A Demonstration Planning Model for India, by Jan Sandee. Journal of the Royal Statistical Society. Series A (General), 124 (4):585, ???? 1961. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342952.

Baines:1961:BRB

[241] A. H. J. Baines. Book review: Studies on Consumer Behaviour, by Amalendu Ganguly; J. Roy; R. G. Laha; Ashok Rudra; Bina Roy; Deb Kumar Bose; S. K. Dhar; I. M. Chakravarti; A. K. Chakraverti. Journal of the Royal Statistical Society. Series A (General), 124 (4):585–586, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342953.

Barna:1961:BRB

[242] T. Barna. Book review: Capital in Manufacturing and Mining: Its Formation and Financing, by D. Creamer; S. P. Dobrovolsky; I. Borenstein. Journal of the Royal Statistical Society. Series A (General), 124(4):587, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342954.

Ray:1961:BRB

[243] G. F. Ray. Book review: Energy in the American Economy, 1850–1975: An Economic Study of its History and Prospects, by Sam H. Schurr; Bruce C. Netschert; Vera F. Eliasberg; Joseph Lerner; Hans H. Landsberg. Journal of the Royal Statistical Society. Series A (General), 124(4):587–589, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342955.

Anonymous:1961:CNd

[244] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 124(4):590–594, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342956.

Anonymous:1961:LBRc

[245] Anonymous. List of books received for review. Journal of the Royal Statistical Society. Series A (General), 124 (4):595–599, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342957.

Anonymous:1961:BMd

[246] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 124(4):i–iv, ???? 1961. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342958.

Anonymous:1962:FMa

[247] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 125(1):??, ???? 1962. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343208.

Liddell:1962:CM

[248] F. D. K. Liddell. Coalminers’ morbidity. Journal of the Royal Statistical Society. Series A (General), 125(1):1–30, ???? 1962. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2343209.

Hill:1962:SID

[249] I. D. Hill. Sampling inspection and defence specification DEF-131. Journal of the Royal Statistical Society. Series A (General), 125(1):31–87, ???? 1962. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343210.

Lewis:1962:AIM

[250] B. N. Lewis. On the analysis of interaction in multi-dimensional contingency tables. Journal of the Royal Statistical Society. Series A (General), 125 (1):88–117, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343211.

Butler:1962:USS

[251] E. B. Butler. University students, staff and recurrent grants, pre-War and post-War. Journal of the Royal Statistical Society. Series A (General), 125 (1):118–143, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343212.

Rudd:1962:SAC

[252] J. Rudd. The statistics of accidents collected by H.M. Factory Inspectorate, Ministry of Labour. Journal of the Royal Statistical Society. Series A (General), 125(1):144–150, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343213.

Lindley:1962:BRB

[253] D. V. Lindley. Book review: Queues, by D. R. Cox; Walter L. Smith. Journal of the Royal Statistical Society. Series A (General), 125(1):151–152, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343214.

Cox:1962:BRB

[254] D. R. Cox. Book review: An Elementary Introduction to the Theory of Probability, by B. V. Gnedenko; A. Ya Khinchin. Journal of the Royal Statistical Society. Series A (General), 125 (1):152–153, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343215.

Johnson:1962:BRB

[255] N. L. Johnson. Book review: Probability with Statistical Applications, by Frederick Mosteller; Robert E. K. Rourke; George B. Thomas,; Probability: A First Course. by Frederick Mosteller; Robert E. K. Rourke; George B. Thomas. Journal of the Royal Statistical Society. Series A (General), 125(1):153–154, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343216.

Downton:1962:BRBa

[256] F. Downton. Book review: Axiomatics of Classical Statistical Mechanics, by R. Kurth. Journal of the Royal Statistical Society. Series A (General), 125 (1):154–155, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343217.

Barton:1962:BRBa

[257] D. E. Barton. Book review: Decomposition of Superpositions of Distribution Functions, by Pal Medgyessy. Journal of the Royal Statistical Society. Series A (General), 125(1):155–156, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343218.

Maxwell:1962:BRB

[258] A. E. Maxwell. Book review: The Rhetoric of Science — A Methodological Discussion of the Two, by Two Table. by Roy G. Francis. Journal of the Royal Statistical Society. Series A (General), 125(1):156–157, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343219.

Oldham:1962:BRB

[259] P. D. Oldham. Book review: Principles of Medical Statistics, by A. Bradford Hill. Journal of the Royal Statistical Society. Series A (General), 125 (1):157–158, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343220.

Gray:1962:BRB

[260] P. G. Gray. Book review: Sample Design in Business Research, by W. Edwards Deming. Journal of the Royal Statistical Society. Series A (General), 125(1):158–159, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343221.

Simpson:1962:BRB

[261] H. R. Simpson. Book review: A Short Manual on Sampling, Vol. 1: Elements of Sample Survey Theory. Journal of the Royal Statistical Society. Series A (General), 125(1):159–160, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343222.

Foster:1962:BRB

[262] F. G. Foster. Book review: Introduction to Congestion Theory in Telephone Systems, by R. Syski; Thomas Eades; H. H. G. Morris. Journal of the Royal Statistical Society. Series A (General), 125(1):160, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343223.

Vajda:1962:BRBa

[263] S. Vajda. Book review: Electronic Computers, Fundamentals, Systems, and Applications, by Paul von Handel. Journal of the Royal Statistical Society. Series A (General), 125(1):161, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343224.

Page:1962:BRBa

[264] E. S. Page. Book review: Adaptive Control Processes: A Guided Tour, by Richard Bellman. Journal of the Royal Statistical Society. Series A (General), 125(1):161–162, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343225.

Vajda:1962:BRBb

[265] S. Vajda. Book review: Progress in Operations Research, by Russell L. Ackoff. Journal of the Royal Statistical Society. Series A (General), 125(1):162– 163, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343226.

Vajda:1962:BRBc

[266] S. Vajda. Book review: Problems in Basic Operations Research Methods for Management, by Randolph W. Cabell; Almarin Phillips. Journal of the Royal Statistical Society. Series A (General), 125(1):163, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343227.

Bound:1962:BRBa

[267] J. A. Bound. Book review: Statistics– with Applications in Management and Economics, by Earl K. Bowen. Journal of the Royal Statistical Society. Series A (General), 125(1):163–164, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343228.

Champernowne:1962:BRB

[268] D. G. Champernowne. Book review: Principles of Economics, by Alfred Marshall; C. W. Guillebaud. Journal of the Royal Statistical Society. Series A (General), 125(1):164–165, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343229.

Baykov:1962:BRB

[269] Alexander Baykov. Book review: Soviet Statistics of Physical Output of Industrial Commodities: Their Compilation and Quality, by Gregory Grossman. Journal of the Royal Statistical Society. Series A (General), 125(1): 165–166, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343230.

Anstey:1962:BRB

[270] Vera Anstey. Book review: The Corporate Sector in India: A Factual Presentation of Long and Short-Term Trends, by Raj K. Nigam; N. C. Chaudhuri. Journal of the Royal Statistical Society. Series A (General), 125(1):166– 167, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343231.

Grebenik:1962:BRBa

[271] E. Grebenik. Book review: Demographic and Economic Change in Developed Countries: Conference of the Universities–National Bureau Committee for Economic Research, New York. Journal of the Royal Statistical Society. Series A (General), 125(1):167– 168, ???? 1962. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343232.

Lafitte:1962:BRB

[272] Francois Lafitte. Book review: Social Security Statistics. Journal of the Royal Statistical Society. Series A (General), 125(1):168–169, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343233.

Halsey:1962:BRB

[273] A. H. Halsey. Book review: The Chosen Few: An Examination of Some Aspects of University Selection in Britain, by W. D. Furneaux. Journal of the Royal Statistical Society. Series A (General), 125(1):169–171, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343234.

Donnison:1962:BRBa

[274] D. V. Donnison. Book review: Housing and Economic Progress: A Study of the Housing Experiences of Boston’s Middle-Income Families, by Lloyd Rodwin. Journal of the Royal Statistical Society. Series A (General), 125(1):171–172, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343235.

Anonymous:1962:CNa

[275] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 125(1):173–177, ???? 1962. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343236.

Anonymous:1962:LBRa

[276] Anonymous. List of books received for review. Journal of the Royal Statistical Society. Series A (General), 125 (1):178–182, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343237.

Anonymous:1962:BMa

[277] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 125(1):i–vi, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343238.

Anonymous:1962:FMb

[278] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 125(2):??, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982324.

Cane:1962:LI

[279] Violet R. Cane. Learning and inference. Journal of the Royal Statistical Society. Series A (General), 125(2): 183–209, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982325.

Ilersic:1962:PHL

[280] A. R. Ilersic. The prospect for householders’ local rates. Journal of the Royal Statistical Society. Series A (General), 125(2):210–231, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982326.

Strauss:1962:SEM

[281] E. Strauss. The structure of English milk distribution. Journal of the Royal Statistical Society. Series A (General), 125(2):232–242, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982327.

Thom:1962:MUL

[282] A. Thom. The megalithic unit of length. Journal of the Royal Statistical Society. Series A (General), 125 (2):243–251, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982328.

Farrell:1962:EEP

[283] M. J. Farrell and M. Fieldhouse. Estimating efficient production functions under increasing returns to scale. Journal of the Royal Statistical Society. Series A (General), 125(2):252–267, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982329.

Haybittle:1962:EYS

[284] J. L. Haybittle. The estimation of T - year survival rate in patients treated for cancer. Journal of the Royal Statistical Society. Series A (General), 125 (2):268–283, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982330.

Plackett:1962:BRB

[285] R. L. Plackett. Book review: The Advanced Theory of Statistics: Vol. 2 — Inference and Relationship, by M. G. Kendall; A. Stuart. Journal of the Royal Statistical Society. Series A (General), 125(2):284–286, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982331.

Hitchcock:1962:BRB

[286] Shirley Hitchcock. Book review: Elementary Statistical Exercises, by F.

N. David; E. S. Pearson. Journal of the Royal Statistical Society. Series A (General), 125(2):286–287, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982332.

Bartholomew:1962:BRB

[287] D. J. Bartholomew. Book review: Principes de Statistique Mathématique, by A. Tortrat. Journal of the Royal Statistical Society. Series A (General), 125(2):287, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982333.

Finch:1962:BRB

[288] P. D. Finch. Book review: On the Use of Markov Processes in One-Server Waiting-Time Problems and Renewal Theory, by J. Th. Runnenburg. Journal of the Royal Statistical Society. Series A (General), 125(2):288, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982334.

Page:1962:BRBb

[289] E. S. Page. Book review: Numerical Analysis: With Emphasis on the Application of Numerical Techniques to Problems of Infinitesimal Calculus in a Single Variable, by Zdenek Kopal. Journal of the Royal Statistical Society. Series A (General), 125(2):288– 289, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982335.

Spicer:1962:BRB

[290] C. C. Spicer. Book review: Sequential Medical Trials, by P. Armitage. Journal of the Royal Statistical Society.

Series A (General), 125(2):289, ???? 1962. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982336.

Wetherill:1962:BRBa

[291] G. B. Wetherill. Book review: Analysing Qualitative Data, by A. E. Maxwell. Journal of the Royal Statistical Society. Series A (General), 125 (2):289–290, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982337.

Walker:1962:BRBa

[292] Monica A. Walker. Book review: Studies in Item Analysis and Prediction, by Herbert Solomon. Journal of the Royal Statistical Society. Series A (General), 125(2):290–291, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2982338.

Jenkins:1962:BRB

[293] G. M. Jenkins. Book review: Introduction to the Statistical Dynamics of Automatic Control Systems, by V. V. Solodovnikov; J. B. Thomas; L. A. Zadek. Journal of the Royal Statistical Society. Series A (General), 125 (2):291–292, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982339.

Mercer:1962:BRB

[294] A. Mercer. Book review: Les Phénomènes d’Attente–Théorie et Applications, by A. Kaufmann; R. Cruon. Journal of the Royal Statistical Society. Series A (General), 125(2):292– 293, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982340.

Vajda:1962:BRBd

[295] S. Vajda. Book review: Pattern in Organization Analysis: A Critical Examination, by Sherman Krupp. Journal of the Royal Statistical Society. Series A (General), 125(2):293–294, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982341.

Benjamin:1962:BRBa

[296] B. Benjamin. Book review: Population, by William Petersen. Journal of the Royal Statistical Society. Series A (General), 125(2):294, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982342.

Benjamin:1962:BRBb

[297] B. Benjamin. Book review: Demographic Trends, 1956–76, in Western Europe and in the United States. Journal of the Royal Statistical Society. Series A (General), 125(2):295, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982343.

Benjamin:1962:BRBc

[298] B. Benjamin. Book review: British Towns: A Statistical Study of their Social and Economic Differences, by C. A. Moser; Wolf Scott. Journal of the Royal Statistical Society. Series A (General), 125(2):295–298, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982344.

Searle:1962:BRB

[299] W. F. Searle. Book review: Family Living Studies: A Symposium. Journal of the Royal Statistical Society. Series A (General), 125(2):298–300, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982345.

Nixon:1962:BRB

[300] J. W. Nixon. Book review: Statistical Data: Population, Employment, Agriculture, Industry, Trade, Transport, External Trade, Finance. Journal of the Royal Statistical Society. Series A (General), 125(2):300–301, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982346.

Cunliffe:1962:BRB

[301] S. V. Cunliffe. Book review: Census of Production Reports. Journal of the Royal Statistical Society. Series A (General), 125(2):301, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982347.

Ridley:1962:BRBa

[302] T. M. Ridley. Book review: The Growth of Public Expenditure in the United Kingdom, by Alan T. Peacock; Jack Wiseman; J. Veverka. Journal of the Royal Statistical Society. Series A (General), 125(2):302–303, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982348.

Peacock:1962:BRB

[303] Alan T. Peacock. Book review: Public Finances: Needs, Sources and Utilization. Journal of the Royal Statistical Society. Series A (General), 125 (2):303–305, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982349.

Deane:1962:BRB

[304] P. M. Deane. Book review: Real Wages in Manufacturing, 1890–1914, by Albert Rees; Donald P. Jacobs. Journal of the Royal Statistical Society. Series A (General), 125(2):305, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982350.

Berman:1962:BRB

[305] L. S. Berman. Book review: Output, Input and Productivity Measurement: Studies in Income and Wealth. Vol. 25, by Conference on Research in Income and Wealth. Journal of the Royal Statistical Society. Series A (General), 125 (2):305–306, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982351.

Maizels:1962:BRB

[306] A. Maizels. Book review: English Overseas Trade Statistics, 1697–1808, by Elizabeth Boody Schumpeter. Journal of the Royal Statistical Society. Series A (General), 125(2):306–307, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982352.

Hemming:1962:BRBa

[307] M. F. W. Hemming. Book review: Production of Commodities by Means of Commodities, by Piero Sraffa. Journal of the Royal Statistical Society. Series A (General), 125(2):307–309, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982353.

Donnison:1962:BRBb

[308] D. V. Donnison. Book review: The Postwar Residential Mortgage Market, by Saul B. Klaman. Journal of the Royal Statistical Society. Series A (General), 125(2):309–310, ???? 1962. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982354.

Kaser:1962:BRB

[309] M. C. Kaser. Book review: Central Planning in Czechoslovakia. Organization for Growth in a Mature Economy, by J. M. Michal. Journal of the Royal Statistical Society. Series A (General), 125(2):310–311, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2982355.

McMahon:1962:BRB

[310] C. W. McMahon. Book review: Europe’s Needs and Resources: Trends and Prospects in Eighteen Countries, by J. Frederic Dewhurst; John O. Coppock; P. Lamartine Yates. Journal of the Royal Statistical Society. Series A (General), 125(2):311–313, ???? 1962. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982356.

Anonymous:1962:CNb

[311] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 125(2):314–317, ???? 1962. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982357.

Anonymous:1962:LBRb

[312] Anonymous. List of books received for review. Journal of the Royal Statistical Society. Series A (General), 125 (2):318–320, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982358.

Anonymous:1962:BMb

[313] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 125(2):i–viii, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982359.

Anonymous:1962:FMc

[314] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 125(3):??, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982405.

Barnard:1962:LIT

[315] G. A. Barnard, G. M. Jenkins, and C. B. Winsten. Likelihood inference and time series. Journal of the Royal Statistical Society. Series A (General), 125(3):321–372, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982406.

Calder:1962:MH

[316] Ritchie Calder. The mathematics of hunger. Journal of the Royal Statistical Society. Series A (General), 125 (3):373–377, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982407.

Benjamin:1962:WPT

[317] B. Benjamin. World population trends. Journal of the Royal Statistical Society. Series A (General), 125(3):378– 386, ???? 1962. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982408.

Passmore:1962:EFR

[318] R. Passmore. Estimation of food requirements. Journal of the Royal Statistical Society. Series A (General), 125 (3):387–398, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982409.

Pirie:1962:FSF

[319] N. W. Pirie. Future sources of food supply: Scientific problems. Journal of the Royal Statistical Society. Series A (General), 125(3):399–417, ???? 1962. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982410.

Clark:1962:FSF

[320] Colin Clark. Future sources of food supply: Economic problems. Journal of the Royal Statistical Society. Series A (General), 125(3):418–448, ???? 1962. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982411.

Thomas:1962:CR

[321] Brinley Thomas. Concluding remarks. Journal of the Royal Statistical Society. Series A (General), 125(3):449– 450, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982412.

Holland:1962:RTB

[322] J. D. Holland. The Reverend Thomas Bayes, F.R.S. (1702–61). Journal of the Royal Statistical Society. Series A (General), 125(3):451–461, ????

1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/

stable/2982413. See corrigenda [443].

Dawson:1962:EER

[323] R. F. F. Dawson. Estimated expenditure on road transport in Great Britain, 1960. Journal of the Royal Statistical Society. Series A (General), 125(3): 462–472, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982414.

Sutherland:1962:SHC

[324] Ian Sutherland. Statistics of human conflict. Journal of the Royal Statistical Society. Series A (General), 125 (3):473–483, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982415.

Irwin:1962:BRB

[325] J. O. Irwin. Book review: Essays on Probability and Statistics, by M. S. Bartlett. Journal of the Royal Statistical Society. Series A (General), 125 (3):484–487, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982416.

Good:1962:BRBa

[326] I. J. Good. Book review: Theory of Probability, by Harold Jeffreys. Journal of the Royal Statistical Society. Series A (General), 125(3):487–489, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982417.

Lloyd:1962:BRB

[327] E. H. Lloyd. Book review: Elements of Mathematical Statistics, by Howard W.

Alexander. Journal of the Royal Statistical Society. Series A (General), 125 (3):489–490, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982418.

Kendall:1962:BRB

[328] M. G. Kendall. Book review: Mathematisches Wörterbuch, by Josef Naas; H. L. Schmid. Journal of the Royal Statistical Society. Series A (General), 125 (3):490–491, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982419.

Booth:1962:BRB

[329] A. D. Booth. Book review: Charles Babbage and his Calculating Engines: Selected Writings, by Charles Babbage and Others. by Philip Morrison; Emily Morrison. Journal of the Royal Statistical Society. Series A (General), 125 (3):491–493, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982420.

Farlie:1962:BRB

[330] D. J. G. Farlie. Book review: Contributions to the Theory of Estimation from Grouped and Partially Grouped Samples, by Gunnar Kulldorf. Journal of the Royal Statistical Society. Series A (General), 125(3):493–494, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982421.

Whittle:1962:BRBa

[331] P. Whittle. Book review: Markov Chains with Stationary Transition Probabilities, by Kai Lai Chung. Journal of the Royal Statistical Society. Series A (General), 125(3):494, ????

1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982422.

Whittle:1962:BRBb

[332] P. Whittle. Book review: Theory of Markov Processes, by E. B. Dynkin; D. E. Brown; T. Kovary. Journal of the Royal Statistical Society. Series A (General), 125(3):495, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982423.

Land:1962:BRB

[333] A. H. Land. Book review: Mathematical Programming, by S. Vajda. Journal of the Royal Statistical Society. Series A (General), 125(3):495–496, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982424.

Owen:1962:BRB

[334] A. R. G. Owen. Book review: Introduction to the Mathematical Theory of Genetic Linkage, by N. T. J. Bailey. Journal of the Royal Statistical Society. Series A (General), 125(3):496–497, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982425.

Downton:1962:BRBb

[335] F. Downton. Book review: Elementary Principles in Statistical Mechanics Developed with Especial Reference to Thermodynamics, by J. Willard Gibbs. Journal of the Royal Statistical Society. Series A (General), 125(3):497, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982426.

Gore:1962:BRB

[336] A. T. Gore. Book review: Refraction and its Components during the Growth of the Eye from the Age of Three, by Arnold Sorsby; B. Benjamin; Michael Sheridan; Janet Stone; G. A. Leary. Journal of the Royal Statistical Society. Series A (General), 125(3):497– 498, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982427.

Tanner:1962:BRB

[337] J. M. Tanner. Book review: The Physical Anthropology of Southern Nigeria: A Biometric Study in Statistical Method, by P. A. Talbot; H. Mulhall. Journal of the Royal Statistical Society. Series A (General), 125(3):498– 499, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982428.

Hill:1962:BRB

[338] A. Bradford Hill. Book review: The Genesis of the British National Health Service, by John Jewkes; Sylvia Jewkes. Journal of the Royal Statistical Society. Series A (General), 125 (3):499, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982429.

Grebenik:1962:BRBb

[339] E. Grebenik. Book review: The Economic Circumstances of Old People, by Dorothy Cole; J. Utting; National Assistance and National Prosperity, by Tony Lynes. Journal of the Royal Statistical Society. Series A (General), 125 (3):499–500, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982430.

Martin:1962:BRB

[340] J. P. Martin. Book review: Murder: A Home Office Research Unit Report, by Evelyn Gibson; S. Klein. Journal of the Royal Statistical Society. Series A (General), 125(3):500–501, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982431.

Benjamin:1962:BRBd

[341] B. Benjamin. Book review: Introducing Population Statistics, by Norma McArthur. Journal of the Royal Statistical Society. Series A (General), 125 (3):501–502, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982432.

Baines:1962:BRB

[342] A. H. J. Baines. Book review: Change in the American National Diet, 1879–1959, by Merrill K. Bennett; Rosamund H. Peirce. Journal of the Royal Statistical Society. Series A (General), 125(3):502–503, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982433.

Barnard:1962:BRBa

[343] G. A. Barnard. Book review: The Case Against Decimalisation, by A. C. Aitken. Journal of the Royal Statistical Society. Series A (General), 125 (3):503–504, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982434.

Robinson:1962:BRB

[344] Joan Robinson. Book review: Planning for India: Selected Explorations in Methodology, by Ragnar Frisch. Journal of the Royal Statistical Society. Series A (General), 125(3):504–506, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982435.

Ridley:1962:BRBb

[345] T. M. Ridley. Book review: The Guardian Wage Indexes: A Series of Indexes of Wage Rates in British Industry since 1948, Weekly and Hourly, together with Indexes of Average Earnings, by Ely Devons; J. R. Crossley. Journal of the Royal Statistical Society. Series A (General), 125(3):506, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982436.

Holmans:1962:BRB

[346] A. E. Holmans. Book review: Capital in the American Economy: Its Formation and Financing, by Simon Kuznets; Elizabeth Jenks. Journal of the Royal Statistical Society. Series A (General), 125(3):506–508, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982437.

Feinstein:1962:BRB

[347] C. H. Feinstein. Book review: Business Cycle Indicators. Vol. I. Contributions to the Analysis of Current Business Conditions; Business Cycle Indicators. Vol. 2. Basic Data on Cyclical Indicators, by Geoffrey H. Moore; Signals of Recession and Recovery: An Experiment with Monthly Reporting, by Julius Shiskin. Journal of the Royal Statistical Society. Series A (General), 125(3):508–509, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982438.

Hemming:1962:BRBb

[348] M. F. W. Hemming. Book review: Equilibrium and Growth in the World Economy, by Ragnar Nurkse. Journal of the Royal Statistical Society. Series A (General), 125(3):509–514, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982439.

Anonymous:1962:CNc

[349] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 125(3):515–519, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982440.

Anonymous:1962:LBRc

[350] Anonymous. List of books received for review. Journal of the Royal Statistical Society. Series A (General), 125 (3):520–523, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982441.

Anonymous:1962:BMc

[351] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 125(3):i–vi, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982442.

Anonymous:1962:VI

[352] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 125(4):680– 684, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982609.

Anonymous:1962:FMd

[353] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 125(4):??, ???? 1962. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982610.

Finney:1962:SEE

[354] D. J. Finney. The statistical evaluation of educational allocation and selection. Journal of the Royal Statistical Society. Series A (General), 125 (4):525–564, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982611.

Stein:1962:RLP

[355] Charles Stein. A remark on the likelihood principle. Journal of the Royal Statistical Society. Series A (General), 125(4):565–568, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2982612.

See comments [356].

Barnard:1962:CSR

[356] G. A. Barnard. Comments on Stein’s “A Remark on the Likelihood Principle”. Journal of the Royal Statistical Society. Series A (General), 125 (4):569–573, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982613. See

[355].

Healy:1962:WYE

[357] M. J. R. Healy and E. L. Jones. Wheat yields in England, 1815–59. Journal of the Royal Statistical Society. Series A (General), 125(4):574–579, ????

1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982614.

Zarkovich:1962:SNH

[358] S. S. Zarkovich. A supplement to “Note on the History of Sampling Methods in Russia”. Journal of the Royal Statistical Society. Series A (General), 125 (4):580–582, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982615.

Jeffcoate:1962:CBK

[359] Grace O. Jeffcoate. The chance of being killed in a road accident for people Born during different years. Journal of the Royal Statistical Society. Series A (General), 125(4):583–587, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982616.

Rao:1962:NSV

[360] A. S. Prakasa Rao. A note on small value transactions in India’s foreign trade. Journal of the Royal Statistical Society. Series A (General), 125(4): 588–591, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982617.

Andrews:1962:RCR

[361] P. W. S. Andrews and Frank A. Friday. The recent controversy over resale price maintenance: A rejoinder. Journal of the Royal Statistical Society. Series A (General), 125(4):592–595, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982618.

Anonymous:1962:WP

[362] Anonymous. Wholesale prices in 1961. Journal of the Royal Statistical Society. Series A (General), 125(4):596– 602, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982619.

Anonymous:1962:RRS

[363] Anonymous. Regulations of the research section. Journal of the Royal Statistical Society. Series A (General), 125(4):603–604, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2982620.

Anonymous:1962:ARC

[364] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 125 (4):605–625, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982621.

Anonymous:1962:PSG

[365] Anonymous. Proceedings of a special general meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 125 (4):626, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982622.

Anonymous:1962:POH

[366] Anonymous. Proceedings of the one hundred and twenty-eighth Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 125 (4):627, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982623.

Anonymous:1962:LR

[367] Anonymous. [list of reviews]. Journal of the Royal Statistical Society. Series A (General), 125(4):628, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982624.

Barnard:1962:BRBb

[368] G. A. Barnard, H. D. Miller, C. Domb, and Cedric A. B. Smith. Book review: Proceedings of the Fourth Berkeley Symposium on Mathematical Statistics and Probability, Held at the Statistical Laboratory, University of California, June 20 through June 30, 1960, by Jerzy Neyman. Journal of the Royal Statistical Society. Series A (General), 125(4):629–633, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982625.

Beaumont:1962:BRB

[369] G. P. Beaumont. Book review: Elements of Mathematical Statistics, by J. F. Ractliffe. Journal of the Royal Statistical Society. Series A (General), 125 (4):633–634, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982626.

Newell:1962:BRB

[370] D. J. Newell. Book review: Introduction to Probability and Statistics, by Henry L. Alder; Edward B. Roessler. Journal of the Royal Statistical Society. Series A (General), 125(4):634– 635, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982627.

Wetherill:1962:BRBb

[371] G. B. Wetherill. Book review: Statistical Decision Theory, by L. Weiss. Journal of the Royal Statistical Society. Series A (General), 125(4):635, ???? 1962. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982628.

Walker:1962:BRBb

[372] Monica A. Walker. Book review: Statistics in Psychological Research, by William S. Ray. Journal of the Royal Statistical Society. Series A (General), 125(4):635–636, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2982629.

Maurice:1962:BRB

[373] Rita J. Maurice. Book review: Use and Abuse of Statistics, by W. J. Reichmann. Journal of the Royal Statistical Society. Series A (General), 125 (4):636–637, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982630.

Stuart:1962:BRB

[374] Alan Stuart. Book review: Bibliography of Nonparametric Statistics, by I. Richard Savage. Journal of the Royal Statistical Society. Series A (General), 125(4):637, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982631.

Ruben:1962:BRB

[375] Harold Ruben. Book review: A Course in the Geometry of n Dimensions, by M. G. Kendall. Journal of the Royal Statistical Society. Series A (General), 125(4):638–639, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982632.

Gilchrist:1962:BRB

[376] W. G. Gilchrist. Book review: Combinatorial Chance, by F. N. David; D. E. Barton. Journal of the Royal Statistical Society. Series A (General), 125 (4):639–640, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982633.

Miller:1962:BRB

[377] H. D. Miller. Book review: The Algebra of Probable Inference, by Richard T. Cox. Journal of the Royal Statistical Society. Series A (General), 125 (4):640–641, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982634.

Barton:1962:BRBb

[378] D. E. Barton. Book review: Tables of Neyman-Shortest Unbiased Confidence Intervals (a) for the Binomial Parameter, (b) for the Poisson Parameter, by Colin R. Blyth; David W. Hutchinson. Journal of the Royal Statistical Society. Series A (General), 125(4):641– 642, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982635.

Barnett:1962:BRB

[379] V. D. Barnett. Book review: Numerical Methods for Scientists and Engineers, by R. W. Hamming. Journal of the Royal Statistical Society. Series A (General), 125(4):642–643, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982636.

Good:1962:BRBb

[380] I. J. Good. Book review: Coding Theorems of Information Theory, by J. Wolfowitz. Journal of the Royal Statistical Society. Series A (General), 125 (4):643–645, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982637.

Barnard:1962:BRBc

[381] G. A. Barnard. Book review: Transmission of Information: A Statistical Theory of Communications, by Robert M. Fano. Journal of the Royal Statistical Society. Series A (General), 125 (4):645, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982638.

Vajda:1962:BRBe

[382] S. Vajda. Book review: Programmation Linéaire, by M. Simonnard. Journal of the Royal Statistical Society. Series A (General), 125(4):645–646, ???? 1962. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982639.

Vajda:1962:BRBf

[383] S. Vajda. Book review: Operations Research/Management Science, by Robert S. Titchen; Arnold J. Rosenthal. Journal of the Royal Statistical Society. Series A (General), 125(4): 646–647, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982640.

Vajda:1962:BRBg

[384] S. Vajda. Book review: A Guide to Operational Research, by W. E. Duckworth. Journal of the Royal Statistical Society. Series A (General), 125(4):

647, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982641.

Vajda:1962:BRBh

[385] S. Vajda. Book review: Some Techniques of Operational Research, by Operational Research Group of the National Coal Board; B. T. Houlden. Journal of the Royal Statistical Society. Series A (General), 125(4):647– 648, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982642.

Mann:1962:BRB

[386] D. F. C. Mann. Book review: Quality Control in the Office, by P. N. Wallis. Journal of the Royal Statistical Society. Series A (General), 125(4):648– 649, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982643.

Menzler:1962:BRB

[387] F. A. A. Menzler. Book review: Equitable Assurances, by Maurice Edward Ogborn. Journal of the Royal Statistical Society. Series A (General), 125 (4):649–652, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982644.

Barr:1962:BRB

[388] A. Barr. Book review: Annual Epidemiological and Vital Statistics, 1958. Journal of the Royal Statistical Society. Series A (General), 125(4):652, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982645.

Rowntree:1962:BRB

[389] J. A. Rowntree. Book review: Manuals on Methods of Estimating Population. Manual 1. Methods of Estimating Total Population for Current Dates. Journal of the Royal Statistical Society. Series A (General), 125(4):653, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982646.

Carrier:1962:BRBa

[390] N. H. Carrier. Book review: The Registrar General’s Statistical Review of England and Wales for the Year 1960, Part II, Tables, Population. Journal of the Royal Statistical Society. Series A (General), 125(4):653–654, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982647.

Carrier:1962:BRBb

[391] N. H. Carrier. Book review: The Registrar General’s Quarterly Return for England and Wales. Third Quarter, 1960. Journal of the Royal Statistical Society. Series A (General), 125 (4):654–655, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982648.

Smeed:1962:BRB

[392] R. J. Smeed. Book review: Road Risks of the Elderly: The Dangers to which they are Exposed, and how they may Reduce the Risk of Serious Injury, by G. Prys Williams; Harold S. Goodwin. Journal of the Royal Statistical Society. Series A (General), 125(4):655– 656, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982649.

Ainsworth:1962:BRB

[393] R. B. Ainsworth. Book review: Man-Power Statistics, 1950–1960. Journal of the Royal Statistical Society. Series A (General), 125(4):656, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982650.

Saunders:1962:BRB

[394] C. T. Saunders. Book review: Facts and Figures about the Church of England, by Statistical Unit of the Central Board of Finance of the Church of England. Journal of the Royal Statistical Society. Series A (General), 125 (4):656–658, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982651.

Finney:1962:BRB

[395] D. J. Finney. Book review: Measurement of Achievement in Mathematics. A Statistical Study on Effectiveness of Board and University Examinations in India, by A. K. Gayen; P. B. Nanda; R. H. Mathur; P. Duari; S. D. Dubey; N. Bhattacharyya. Journal of the Royal Statistical Society. Series A (General), 125(4):658–659, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982652.

Revell:1962:BRB

[396] J. R. S. Revell. Book review: The Share of Top Wealth-Holders in National Wealth, 1922–56, by Robert J. Lampman. Journal of the Royal Statistical Society. Series A (General), 125 (4):659–660, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982653.

Bound:1962:BRBb

[397] J. A. Bound. Book review: The Industries of London since 1861, by P. G. Hall. Journal of the Royal Statistical Society. Series A (General), 125 (4):660, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982654.

Pankhurst:1962:BRB

[398] Kenneth Pankhurst. Book review: The Welsh Economy: Studies in Expansion, by Brinley Thomas. Journal of the Royal Statistical Society. Series A (General), 125(4):660–662, ???? 1962. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982655.

Parkinson:1962:BRB

[399] J. R. Parkinson. Book review: Scotland’s Scientific Heritage, by A. G. Clement; Robert H. S. Robertson. Journal of the Royal Statistical Society. Series A (General), 125(4):662– 663, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982656.

Anonymous:1962:CNd

[400] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 125(4):664–667, ???? 1962. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982657.

Anonymous:1962:SRA

[401] Anonymous. Sir Ronald Aylmer Fisher, 1890–1962. Journal of the Royal Statistical Society. Series A (General), 125 (4):668, ???? 1962. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2982658.

Forte:1962:LE

[402] Francesco Forte. Luigi Einaudi, 1874– 1961. Journal of the Royal Statistical Society. Series A (General), 125(4): 668–669, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982659.

Menzler:1962:SWP

[403] F. A. A. Menzler. Sir William Palin Elderton, 1877–1962. Journal of the Royal Statistical Society. Series A (General), 125(4):669–672, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982660.

Ogborn:1962:JFS

[404] M. E. Ogborn. Johan Frederik Steffensen, 1873–1961. Journal of the Royal Statistical Society. Series A (General), 125(4):672–673, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982661.

Anonymous:1962:LBRd

[405] Anonymous. List of books received for review. Journal of the Royal Statistical Society. Series A (General), 125 (4):674–679, ???? 1962. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982662.

Anonymous:1962:BMd

[406] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 125(4):i–vi, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982663.

Anonymous:1963:VIa

[407] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 126(1):??, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982443.

Anonymous:1963:FMa

[408] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 126(1):??, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982444.

Irwin:1963:PMM

[409] J. O. Irwin. The place of mathematics in medical and biological statistics. Journal of the Royal Statistical Society. Series A (General), 126(1):1–45, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982445.

Foster:1963:ESB

[410] C. D. Foster and M. E. Beesley. Estimating the social benefit of constructing an underground railway in London. Journal of the Royal Statistical Society. Series A (General), 126(1):46–93, ???? 1963. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982446.

Goodman:1963:MCC

[411] Leo A. Goodman. On methods for comparing contingency tables. Journal of the Royal Statistical Society. Series A (General), 126(1):94–108, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982447.

Edwards:1963:MAT

[412] A. W. F. Edwards. The measure of association in a 2 × 2 table. Journal of the Royal Statistical Society. Series A (General), 126(1):109–114, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982448.

Glasser:1963:RNS

[413] Gerald J. Glasser. Random numbers, sample selection and occupancy problems. Journal of the Royal Statistical Society. Series A (General), 126(1): 115–119, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982449.

Draper:1963:DAA

[414] Norman R. Draper. Does age affect master chess? Journal of the Royal Statistical Society. Series A (General), 126(1):120–127, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982450.

Whittle:1963:BRB

[415] P. Whittle. Book review: Mathematical Statistics, by S. S. Wilks. Journal of the Royal Statistical Society. Series A (General), 126(1):128–129, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982451.

Silvey:1963:BRB

[416] S. D. Silvey. Book review: Mathematical Statistics, by John E. Freund. Journal of the Royal Statistical Society. Series A (General), 126(1):130, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982452.

Kendall:1963:BRB

[417] D. G. Kendall. Book review: Random Variables and Probability Distributions, by Harald Cramér. Journal of the Royal Statistical Society. Series A (General), 126(1):130–131, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982453.

Loynes:1963:BRB

[418] R. M. Loynes. Book review: Random Processes, by Murray Rosenblatt. Journal of the Royal Statistical Society. Series A (General), 126(1):131, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982454.

Bartholomew:1963:BRB

[419] D. J. Bartholomew. Book review: An Introduction to Probability and Mathematical Statistics, by Howard G. Tucker. Journal of the Royal Statistical Society. Series A (General), 126 (1):131–132, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982455.

James:1963:BRB

[420] G. S. James. Book review: Elements of Probability and Statistics, by F. L. Wolf. Journal of the Royal Statistical Society. Series A (General), 126(1): 132–133, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982456.

Vajda:1963:BRBa

[421] S. Vajda. Book review: Studies in Applied Probability and Management Science, by Kenneth J. Arrow; Samuel Karlin; Herbert Scarf. Journal of the Royal Statistical Society. Series A (General), 126(1):133–134, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982457.

Stuart:1963:BRB

[422] Alan Stuart. Book review: Handbook of Nonparametric Statistics: Investigations of Randomness, Moments, Percentiles, and Distributions, by John E. Walsh. Journal of the Royal Statistical Society. Series A (General), 126 (1):134, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982458.

Welch:1963:BRB

[423] B. L. Welch. Book review: The Foundations of Statistical Inference, by L. G. Savage. Journal of the Royal Statistical Society. Series A (General), 126 (1):134–135, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982459.

Cochran:1963:BRB

[424] W. G. Cochran. Book review: An Introduction to Sampling Theory (with Applications to Agriculture), by M. R. Sampford. Journal of the Royal Statistical Society. Series A (General), 126 (1):135–136, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982460.

Downton:1963:BRB

[425] F. Downton. Book review: Elements of Queueing Theory: With Applications, by T. L. Saaty. Journal of the Royal Statistical Society. Series A (General), 126(1):136–137, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2982461.

Thatcher:1963:BRBa

[426] A. R. Thatcher. Book review: Computers in Education, by J. A. P. Hall. Journal of the Royal Statistical Society. Series A (General), 126(1):137, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982462.

Nixon:1963:BRB

[427] E. Nixon. Book review: An Introduction to Computational Methods, by K. A. Redish. Journal of the Royal Statistical Society. Series A (General), 126 (1):138, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982463.

Hamaker:1963:BRB

[428] H. C. Hamaker. Book review: Defence Specification DEF-131: Sampling Procedures and Tables for Inspection by Attributes; Defence Guide to Sampling Inspection: A Guide to the Use of Defence Specification DEF-131. Journal of the Royal Statistical Society. Series A (General), 126(1):138–140, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982464.

Forsyth:1963:BRB

[429] F. G. Forsyth. Book review: Family Composition and Consumption, by M. H. David. Journal of the Royal Statistical Society. Series A (General), 126 (1):140–141, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982465.

Heady:1963:BRBa

[430] J. A. Heady. Book review: Defects, Uses, and Abuses of National Food Supply and Consumption Data, by Helen C. Farnsworth. Journal of the Royal Statistical Society. Series A (General), 126(1):141, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982466.

Grebenik:1963:BRBa

[431] E. Grebenik. Book review: Income Distribution and Social Change: A Study in Criticism, by Richard M. Titmuss. Journal of the Royal Statistical Society. Series A (General), 126(1):142, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982467.

Carrier:1963:BRBa

[432] N. H. Carrier. Book review: The Registrar General’s Statistical Review of England and Wales for the Year 1959, Part III, Commentary. Journal of the Royal Statistical Society. Series A (General), 126(1):143–144, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982468.

Tanner:1963:BRB

[433] J. C. Tanner. Book review: Nigerian Motor Vehicle Traffic. An Economic Forecast, by V. W. Hogg; C. M. Roelandts. Journal of the Royal Statistical Society. Series A (General), 126 (1):144, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982469.

Knapp:1963:BRB

[434] J. A. Knapp. Book review: National Income and Expenditure, by Richard Stone; Giovanna Stone. Journal of the Royal Statistical Society. Series A (General), 126(1):144–145, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982470.

Hawtrey:1963:BRB

[435] Ralph Hawtrey. Book review: Studies in an Inflationary Economy: The United Kingdom 1948–1961, by F. W. Paish. Journal of the Royal Statistical Society. Series A (General), 126(1): 145–146, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982471.

Clark:1963:BRBa

[436] Colin Clark. Book review: The Growth of Industrial Production in the Soviet Union, by G. Warren Nutter. Journal of the Royal Statistical Society. Series A (General), 126(1):146–147, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982472.

McMahon:1963:BRB

[437] C. W. McMahon. Book review: Productivity Trends in the United States, by John W. Kendrick; Maude R. Pech. Journal of the Royal Statistical Society. Series A (General), 126(1):148– 149, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982473.

Freeman:1963:BRB

[438] C. Freeman. Book review: The Rate and Direction of Inventive Activity:

Economic and Social Factors. Journal of the Royal Statistical Society. Series A (General), 126(1):149–150, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982474.

Feinstein:1963:BRB

[439] C. H. Feinstein. Book review: Postwar Cycles in Manufacturers’ Inventories, by Thomas M. Stanback. Journal of the Royal Statistical Society. Series A (General), 126(1):150–152, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982475.

Anonymous:1963:CNa

[440] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 126(1):153–158, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982476.

Irwin:1963:SRA

[441] J. O. Irwin, G. A. Barnard, Kenneth Mather, F. Yates, and M. J. R. Healy. Sir Ronald Aylmer Fisher, 1890–1962. Journal of the Royal Statistical Society. Series A (General), 126(1):159– 178, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982477.

Anonymous:1963:LAa

[442] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 126(1):179–185, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982478.

Holland:1963:CRT

[443] J. D. Holland. Corrigenda: The Reverend Thomas Bayes, F.R.S. (1702–61). Journal of the Royal Statistical Society. Series A (General), 126(1):185, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982479. See [322].

Anonymous:1963:BMa

[444] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 126(1):i–vi, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982480.

Anonymous:1963:FMb

[445] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 126(2):??, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982360.

Barnett:1963:SRB

[446] V. D. Barnett and T. Lewis. A study of the relation between G.C.E. and degree results. Journal of the Royal Statistical Society. Series A (General), 126 (2):187–226, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982361.

Saunders:1963:ICP

[447] C. T. Saunders and R. E. Crum. International comparisons of productivity growth in the 1950’s. Journal of the Royal Statistical Society. Series A (General), 126(2):227–236, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982362.

Johnston:1963:PMC

[448] J. Johnston. The productivity of management consultants. Journal of the Royal Statistical Society. Series A (General), 126(2):237–249, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982363.

Dickinson:1963:DWR

[449] H. D. Dickinson. Do we really want productivity? Journal of the Royal Statistical Society. Series A (General), 126 (2):250–254, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982364.

Barnard:1963:NMQ

[450] G. A. Barnard. New methods of quality control. Journal of the Royal Statistical Society. Series A (General), 126 (2):255–258, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982365.

Daisley:1963:QAW

[451] P. A. Daisley. A quality approach to waste reduction. Journal of the Royal Statistical Society. Series A (General), 126(2):259–262, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982366.

Tippett:1963:PMA

[452] L. H. C. Tippett. Productivity measurement as an aid to productivity improvement. Journal of the Royal Statistical Society. Series A (General), 126 (2):262–266, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982367.

Sears:1963:PIT

[453] G. W. Sears. Productivity improvement through operational research. Journal of the Royal Statistical Society. Series A (General), 126(2):267– 269, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982368.

Healy:1963:SIK

[454] M. J. R. Healy. A subject index to the Kendall-doig bibliography of statistical literature. Journal of the Royal Statistical Society. Series A (General), 126 (2):270–275, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982369.

Tanner:1963:CMO

[455] J. C. Tanner. Car and motorcycle ownership in the counties of Great Britain in 1960. Journal of the Royal Statistical Society. Series A (General), 126 (2):276–284, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982370.

Bain:1963:GDN

[456] A. D. Bain. The growth of demand for new commodities. Journal of the Royal Statistical Society. Series A (General), 126(2):285–299, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982371.

Swift:1963:DWN

[457] B. M. Swift. The duration of wage negotiations, 1950–58. Journal of the Royal Statistical Society. Series A (General), 126(2):300–314, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982372.

Wishart:1963:BRB

[458] D. M. G. Wishart. Book review: Renewal Theory, by D. R. Cox. Journal of the Royal Statistical Society. Series A (General), 126(2):315–316, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982373.

Armitage:1963:BRB

[459] P. Armitage. Book review: Statistical Methods in Clinical and Preventive Medicine, by Austin Bradford Hill. Journal of the Royal Statistical Society. Series A (General), 126(2):316– 317, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982374.

Plackett:1963:BRB

[460] R. L. Plackett. Book review: Contributions to Order Statistics, by Ahmed E. Sarhan; Bernard G. Greenberg. Journal of the Royal Statistical Society. Series A (General), 126(2):317–318, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982375.

Healy:1963:BRBa

[461] M. J. R. Healy. Book review: Mathematics of Design and Analysis of Experiments, by M. C. Chakrabarti. Journal of the Royal Statistical Society. Series A (General), 126(2):318, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982376.

Farlie:1963:BRBa

[462] D. Farlie. Book review: Introduction to Statistical Inference, by E. S. Keeping. Journal of the Royal Statistical Society. Series A (General), 126(2):318– 319, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982377.

Thatcher:1963:BRBb

[463] A. R. Thatcher. Book review: Games, Gods and Gambling: The Origins and History of Probability and Statistical Ideas from the Earliest Times to the Newtonian Era, by F. N. David. Journal of the Royal Statistical Society. Series A (General), 126(2):319–320, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982378.

Beaumont:1963:BRBa

[464] G. P. Beaumont. Book review: Statistics. An Intermediate Text-book, by N. L. Johnson; H. Tetley. Journal of the Royal Statistical Society. Series A (General), 126(2):320–321, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982379.

Gales:1963:BRB

[465] K. E. Gales. Book review: Statistics, An Intuitive Approach, by George H. Weinberg; John A. Schumaker. Journal of the Royal Statistical Society. Series A (General), 126(2):321, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982380.

Heady:1963:BRBb

[466] J. A. Heady. Book review: How to Read Statistics, by I. R. Vesselo. Journal of the Royal Statistical Society. Series A (General), 126(2):322, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982381.

Barton:1963:BRBa

[467] D. E. Barton. Book review: The Theory of Graphs and its Applications, by Claude Berge; Alison Doig. Journal of the Royal Statistical Society. Series A (General), 126(2):322–323, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982382.

Barton:1963:BRBb

[468] D. E. Barton. Book review: Bibliography of Statistical Literature, 1950– 1958, by M. G. Kendall; Alison Doig. Journal of the Royal Statistical Society. Series A (General), 126(2):323– 324, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982383.

Barnett:1963:BRB

[469] V. D. Barnett. Book review: Handbook of Statistical Tables, by D. B. Owen. Journal of the Royal Statistical Society. Series A (General), 126(2):325– 326, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982384.

Champernowne:1963:BRBa

[470] D. G. Champernowne. Book review: Basic Mathematics, by R. G. D. Allen. Journal of the Royal Statistical Society. Series A (General), 126(2):326, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982385.

Vajda:1963:BRBb

[471] S. Vajda. Book review: Statistical Management of Inventory Systems, by Harvey M. Wagner. Journal of the Royal Statistical Society. Series A (General), 126(2):326–327, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982386.

Lindley:1963:BRB

[472] D. V. Lindley. Book review: Stochastic Service Systems, by John Riordan. Journal of the Royal Statistical Society. Series A (General), 126(2):327– 328, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982387.

Segal:1963:BRB

[473] G. Segal and E. Nixon. Book review: Electronic Digital Computers, by G. D. Smirnov. Journal of the Royal Statistical Society. Series A (General), 126 (2):328, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982388.

Healy:1963:BRBb

[474] M. J. R. Healy. Book review: An Introduction to Symbolic Programming, by P. Wegner. Journal of the Royal Statistical Society. Series A (General), 126 (2):328, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982389.

Revell:1963:BRBa

[475] J. R. S. Revell. Book review: The Flow-of-Funds Approach to Social Accounting. Appraisal, Analysis and Applications. Journal of the Royal Statistical Society. Series A (General), 126 (2):328–329, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982390.

Dosser:1963:BRBa

[476] Douglas Dosser. Book review: Systems of National Accounts in Africa, by P. Ady; M. Courcier. Journal of the Royal Statistical Society. Series A (General), 126(2):330, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982391.

Dosser:1963:BRBb

[477] Douglas Dosser. Book review: Nigerian National Accounts, 1950–57, by P. M. C. Okigbo. Journal of the Royal Statistical Society. Series A (General), 126 (2):330–331, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982392.

Henderson:1963:BRB

[478] P. D. Henderson. Book review: Capital Funds in Underdeveloped Countries. The Role of Financial Institutions, by Edward Nevin. Journal of the Royal Statistical Society. Series A (General), 126(2):331–332, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982393.

Champernowne:1963:BRBb

[479] D. G. Champernowne. Book review: Economic Philosophy, by Joan Robinson. Journal of the Royal Statistical Society. Series A (General), 126(2):332, ???? 1963. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982394.

Corlett:1963:BRB

[480] W. J. Corlett. Book review: Statistical Measurements for Economics and Administration, by M. Budin. Journal of the Royal Statistical Society.

Series A (General), 126(2):333, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982395.

Buckland:1963:BRBa

[481] W. R. Buckland. Book review: Business and Economic Forecasting: An Econometric Approach, by M. H. Spencer; C. G. Clark; P. W. Hoguet. Journal of the Royal Statistical Society. Series A (General), 126(2):333– 334, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982396.

Buckland:1963:BRBb

[482] W. R. Buckland. Book review: Business and Economic Statistics, by W. A. Spurr; L. S. Kellogg; J. H. Smith. Journal of the Royal Statistical Society. Series A (General), 126(2):334–335, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982397.

Florence:1963:BRB

[483] P. Sargant Florence. Book review: Diversification and Integration in American Industry, by Michael Gort. Journal of the Royal Statistical Society. Series A (General), 126(2):335–336, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982398.

Baykov:1963:BRB

[484] Alexander Baykov. Book review: Small-Scale Industry in the Soviet Union, by Adam Kaufman. Journal of the Royal Statistical Society. Series A (General), 126(2):336–338, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982399.

Glover:1963:BRB

[485] F. J. Glover. Book review: Trends in Natural Resource Commodities: Statistics of Prices, Output, Consumption, Foreign Trade, and Employment in the United States, 1870–1957, by Neal Potter; Francis T. Christy. Journal of the Royal Statistical Society. Series A (General), 126(2):338–340, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982400.

Anonymous:1963:CNb

[486] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 126(2):341–345, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982401.

Ridley:1963:HSB

[487] T. M. Ridley. Harold Scott booker, 1903–1963. Journal of the Royal Statistical Society. Series A (General), 126 (2):346–347, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982402.

Anonymous:1963:LAb

[488] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 126(2):348–352, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982403.

Anonymous:1963:BMb

[489] Anonymous. Back matter. Journal of the Royal Statistical Society.

Series A (General), 126(2):i–vi, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982404.

Anonymous:1963:FMc

[490] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 126(3):??, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982221.

Pearce:1963:UCN

[491] S. C. Pearce. The use and classification of non-orthogonal designs. Journal of the Royal Statistical Society. Series A (General), 126(3):353–377, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982222.

Wagenfuhr:1963:SPE

[492] Rolf Wagenführ. The statistical problems of European integration. Journal of the Royal Statistical Society. Series A (General), 126(3):378–399, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982223.

Gani:1963:FPE

[493] J. Gani. Formulae for projecting enrolments and degrees awarded in universities. Journal of the Royal Statistical Society. Series A (General), 126 (3):400–409, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982224.

Matuszewski:1963:ATI

[494] T. I. Matuszewski, Paul R. Pitts, and John A. Sawyer. Alternative treatments of imports in input-output models: A Canadian study. Journal of the Royal Statistical Society. Series A (General), 126(3):410–432, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982225.

Gupta:1963:IOT

[495] S. Gupta. Input and output trends in British manufacturing industry, 1948– 54. Journal of the Royal Statistical Society. Series A (General), 126(3):433– 445, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982226.

Dorrance:1963:EFT

[496] Graeme S. Dorrance. The entries in financial transactions and balance sheet accounts. Journal of the Royal Statistical Society. Series A (General), 126 (3):446–465, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982227.

Buckland:1963:BRBc

[497] W. R. Buckland. Book review: Studi in Onore di Corrado Gini. Journal of the Royal Statistical Society. Series A (General), 126(3):466–467, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982228.

Buckland:1963:BRBd

[498] W. R. Buckland. Book review: A Guide to Tables in Mathematical Statistics, by J. Arthur Greenwood; H. O. Hartley. Journal of the Royal Statistical Society. Series A (General), 126 (3):467–468, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982229.

Stone:1963:BRBa

[499] M. Stone. Book review: Méthodes Statistiques dans les Sciences Humaines, by P. Pèpe; M. Tisserand-Perrier. Journal of the Royal Statistical Society. Series A (General), 126 (3):468–469, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982230.

Stone:1963:BRBb

[500] M. Stone. Book review: Mathematical Methods in Small Group Processes, by Joan Criswell; Herbert Solomon; Patrick Suppes. Journal of the Royal Statistical Society. Series A (General), 126(3):469, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982231.

Farlie:1963:BRBb

[501] D. J. Farlie. Book review: Statistical Analysis in the Geological Sciences, by R. L. Miller; J. S. Kahn. Journal of the Royal Statistical Society. Series A (General), 126(3):469–470, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982232.

Beale:1963:BRB

[502] E. M. L. Beale. Book review: Readings in Mathematical Programming, by S. Vajda. Journal of the Royal Statistical Society. Series A (General), 126 (3):470, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982233.

Priestley:1963:BRB

[503] M. B. Priestley. Book review: Random Wavelets and Cybernetic Systems, by Enders A. Robinson. Journal of the Royal Statistical Society. Series A (General), 126(3):471, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982234.

King:1963:BRB

[504] P. J. H. King. Book review: Proceedings of a Harvard Symposium on Digital Computers and their Applications, by A. C. Oettinger. Journal of the Royal Statistical Society. Series A (General), 126(3):471–472, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982235.

Palca:1963:BRB

[505] H. Palca. Book review: Variance and Production Function Analyses of Farm Accounts. With Special Reference to Inter-Farm and Random Variations, by Knud Rasmussen. Journal of the Royal Statistical Society. Series A (General), 126(3):472, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982236.

Mills:1963:BRB

[506] G. Mills. Book review: Approaches to Dynamic Investment Planning, by Stephen A. Marglin. Journal of the Royal Statistical Society. Series A (General), 126(3):473, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982237.

Ashton:1963:BRB

[507] T. S. Ashton. Book review: Politics and Economics. Papers in Political Economy, by Lord Robbins. Journal of the Royal Statistical Society. Series A (General), 126(3):473–474, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982238.

Revell:1963:BRBb

[508] J. R. S. Revell. Book review: The National Wealth of the United States in the Postwar Period, by Raymond W. Goldsmith. Journal of the Royal Statistical Society. Series A (General), 126 (3):474–475, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982239.

Searle:1963:BRB

[509] W. F. Searle. Book review: Consumption of Exotic Consumer Goods as an Indicator of Economic Achievement in Ten Countries of Tropical Africa, by William O. Jones; Christian Merát. Journal of the Royal Statistical Society. Series A (General), 126(3):475, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982240.

West:1963:BRB

[510] Christine West. Book review: The Ownership of Major Consumer Durables; A Statistical Survey of Motor-Cars, Refrigerators, Washing Machines and Television Sets in the Oxford Savings Survey of 1953. Journal of the Royal Statistical Society. Series A (General), 126(3):475–476, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982241.

Turner:1963:BRB

[511] H. A. Turner. Book review: Unemployment and Structural Change. Journal of the Royal Statistical Society. Series A (General), 126(3):476–477, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982242.

Williams:1963:BRB

[512] Gertrude Williams. Book review: The Reluctant Job Changer: Studies in Work Attachments and Aspirations, by Gladys L. Palmer; Herbert S. Parnes; Richard C. Wilcock; Mary W. Herman; Carol P. Branard. Journal of the Royal Statistical Society. Series A (General), 126(3):477–479, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982243.

Leser:1963:BRB

[513] C. E. V. Leser. Book review: A Study of United Kingdom Imports, by M. FG. Scott. Journal of the Royal Statistical Society. Series A (General), 126(3): 479–480, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982244.

Lloyd:1963:BRB

[514] E. H. Lloyd. Book review: Design of Water-Resource Systems; New Techniques for Relating Economic Objectives, Engineering Analysis, and Government Planning, by Arthur Maass; Maynard M. Hufschmidt; Robert Dorfman; Harold A. Thomas; Stephen A. Marglin; Gordon Masken Fair. Journal of the Royal Statistical Society. Series A (General), 126(3):480–481, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982245.

Carrier:1963:BRBb

[515] N. H. Carrier. Book review: The Registrar General’s Quarterly Return for England and Wales. Fourth Quarter, 1962. Journal of the Royal Statistical Society. Series A (General), 126(3): 481–482, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982246.

Heady:1963:BRBc

[516] J. A. Heady. Book review: Morbidity Statistics from General Practice, by The Research Committee of the Council of the College of General Practitioners. Journal of the Royal Statistical Society. Series A (General), 126 (3):482–483, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982247.

Anonymous:1963:CNc

[517] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 126(3):484–486, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982248.

Anonymous:1963:VS

[518] Anonymous. Viscount Samuel, 1870– 1963. Journal of the Royal Statistical Society. Series A (General), 126(3): 487, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982249.

Anonymous:1963:LAc

[519] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 126(3):488–491, ????

1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982250.

Anonymous:1963:BMc

[520] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 126(3):i–vi, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982251.

Anonymous:1963:VIb

[521] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 126(4):626– 629, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982575.

Anonymous:1963:FMd

[522] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 126(4):??, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982576.

Moore:1963:SAA

[523] P. G. Moore. A statistical approach to the allocation of technical effort in some industrial situations. Journal of the Royal Statistical Society. Series A (General), 126(4):493–536, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982577.

Sutherland:1963:JGT

[524] Ian Sutherland. John Graunt: A tercentenary tribute. Journal of the Royal Statistical Society. Series A (General), 126(4):537–556, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2982578.

Kokan:1963:OAM

[525] A. R. Kokan. Optimum allocation in multivariate surveys. Journal of the Royal Statistical Society. Series A (General), 126(4):557–565, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/

stable/2982579. See errata [776].

Anonymous:1963:WP

[526] Anonymous. Wholesale prices in 1962. Journal of the Royal Statistical Society. Series A (General), 126(4):566– 571, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982580.

Anonymous:1963:ARC

[527] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 126 (4):572–591, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982581.

Anonymous:1963:POH

[528] Anonymous. Proceedings of the one hundred and twenty-ninth Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 126 (4):592, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982582.

Anonymous:1963:PSG

[529] Anonymous. Proceedings of a special general meeting. Journal of the Royal Statistical Society. Series A (General), 126(4):593–594, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982583.

Beaumont:1963:BRBb

[530] G. P. Beaumont. Book review: Introduction to Statistical Inference, by H. F. Freeman. Journal of the Royal Statistical Society. Series A (General), 126 (4):595–596, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982584.

Ilersic:1963:BRB

[531] A. R. Ilersic. Book review: Statistics: Methods and Applications, by John I. Griffin. Journal of the Royal Statistical Society. Series A (General), 126 (4):596–597, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982585.

Champernowne:1963:BRBc

[532] D. G. Champernowne. Book review: A Priori Information and Time Series Analysis: Essays in Economic Theory and Measurement, by Franklin M. Fisher. Journal of the Royal Statistical Society. Series A (General), 126 (4):597–598, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982586.

Tocher:1963:BRB

[533] K. D. Tocher. Book review: Introduction to the Theory of Queues, by L. Takacs. Journal of the Royal Statistical Society. Series A (General), 126 (4):598, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982587.

Bather:1963:BRB

[534] J. A. Bather. Book review: Applied Dynamic Programming, by Richard E. Bellman; Stuart E. Dreyfus. Journal of the Royal Statistical Society. Series A (General), 126(4):598–599, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982588.

Champernowne:1963:BRBd

[535] D. G. Champernowne. Book review: Decision, Order and Time in Human Affairs, by G. L. S. Shackle. Journal of the Royal Statistical Society. Series A (General), 126(4):599–600, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982589.

Nelder:1963:BRB

[536] J. A. Nelder. Book review: The Essence of Biometry, by John Stanley. Journal of the Royal Statistical Society. Series A (General), 126(4):600– 601, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982590.

Farlie:1963:BRBc

[537] D. J. Farlie. Book review: The Calculus of Linguistic Observations, by G. Herdan. Journal of the Royal Statistical Society. Series A (General), 126 (4):601, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982591.

Shankleman:1963:BRB

[538] Eric Shankleman. Book review: Quantitative Techniques in Marketing Analysis, by Ronald E. Frank; Alfred A.

Kuehn; William F. Massy. Journal of the Royal Statistical Society. Series A (General), 126(4):602–603, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982592.

Clark:1963:BRBb

[539] Colin Clark. Book review: A Programme for Growth, 1. A Computable Model of Economic Growth, by Richard Stone; Alan Brown. Journal of the Royal Statistical Society. Series A (General), 126(4):603–604, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982593.

Barna:1963:BRB

[540] T. Barna. Book review: A Programme for Growth, 2. A Social Accounting Matrix for 1960, by Richard Stone; Alan Brown. Journal of the Royal Statistical Society. Series A (General), 126 (4):604–605, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982594.

Seton:1963:BRB

[541] F. Seton. Book review: Input-Output Tables (their Compilation and Use), by O. Lukács; Gy. Cukor; P. Havas; Z. Román. Journal of the Royal Statistical Society. Series A (General), 126 (4):605–607, ???? 1963. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982595.

Saunders:1963:BRB

[542] C. T. Saunders. Book review: The British Economy in the Nineteenth-Fifties, by G. D. N. Worswick; P. H. Ady. Journal of the Royal Statistical Society. Series A (General), 126(4):

607–608, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982596.

Champernowne:1963:BRBe

[543] D. G. Champernowne. Book review: Inflation: Proceedings of a Conference held, by the International Economic Association by D. C. Hague. Journal of the Royal Statistical Society. Series A (General), 126(4):608–609, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982597.

Prest:1963:BRB

[544] A. R. Prest. Book review: African Studies in Income and Wealth, by L. H. Samuels. Journal of the Royal Statistical Society. Series A (General), 126 (4):609, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982598.

Rybczynski:1963:BRB

[545] T. M. Rybczynski. Book review: Problems of the United States as World Trader and Banker, by Hal B. Lary. Journal of the Royal Statistical Society. Series A (General), 126(4):610– 611, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982599.

Grebenik:1963:BRBb

[546] E. Grebenik. Book review: Problems in Stable Population Theory, by Alvaro Lopez. Journal of the Royal Statistical Society. Series A (General), 126(4): 611, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982600.

Benjamin:1963:BRB

[547] B. Benjamin. Book review: Measurement of Mortality, by Harry Gershenson. Journal of the Royal Statistical Society. Series A (General), 126(4):612, ???? 1963. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982601.

Seear:1963:BRB

[548] B. N. Seear. Book review: Woman, Work and Age: To Study the Employment of Working Women Throughout Their Middle Lives, by F. Le Gros Clark. Journal of the Royal Statistical Society. Series A (General), 126(4): 612–614, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982602.

Payne:1963:BRB

[549] P. M. Payne. Book review: St. Bartholomew’s Hospital Reports, by A. W. Franklin; M. P. Curwen. Journal of the Royal Statistical Society. Series A (General), 126(4):614, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982603.

Anonymous:1963:CNd

[550] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 126(4):615–617, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982604.

Chester:1963:LB

[551] D. N. Chester. Lord Beveridge, 1879– 1963. Journal of the Royal Statistical Society. Series A (General), 126(4):

618–620, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982605.

Kendall:1963:WFS

[552] W. L. Kendall. William Frederick Searle, 1909–1963. Journal of the Royal Statistical Society. Series A (General), 126(4):620–621, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2982606.

Anonymous:1963:LAd

[553] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 126(4):622–625, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982607.

Anonymous:1963:BMd

[554] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 126(4):i–vi, ???? 1963. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982608.

Anonymous:1964:VIa

[555] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 127(1):??, ???? 1964. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982288.

Anonymous:1964:FMa

[556] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 127(1):??, ???? 1964. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982289.

Newell:1964:SAD

[557] D. J. Newell. Statistical aspects of the demand for maternity beds. Journal of the Royal Statistical Society. Series A (General), 127(1):1–41, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982290.

Oliver:1964:IRM

[558] F. R. Oliver. Inter-regional migration and unemployment, 1951–61. Journal of the Royal Statistical Society. Series A (General), 127(1):42–75, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982291.

Allen:1964:SCE

[559] R. G. D. Allen. Sampling for current economic statistics. Journal of the Royal Statistical Society. Series A (General), 127(1):76–88, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982292.

Dean:1964:FIB

[560] Geoffry Dean. Fixed investment in Britain and Norway an experiment in international comparison. Journal of the Royal Statistical Society. Series A (General), 127(1):89–107, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982293.

Idaikkadar:1964:NUM

[561] N. M. Idaikkadar. A note on undernourishment in Malaya. Journal of the Royal Statistical Society. Series A (General), 127(1):108–110, ????

1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982294.

Thomas:1964:DBE

[562] Brinley Thomas. The dimensions of British economic growth, 1688–1959. Journal of the Royal Statistical Society. Series A (General), 127(1):111– 123, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982295.

Welch:1964:BRB

[563] B. L. Welch. Book review: Introduction to the Theory of Statistics, by A. M. Mood; F. A. Graybill. Journal of the Royal Statistical Society. Series A (General), 127(1):124–125, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982296.

James:1964:BRB

[564] G. S. James. Book review: Probability and Statistical Methods, by R. R. Umarji. Journal of the Royal Statistical Society. Series A (General), 127 (1):125–126, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982297.

Hitchcook:1964:BRB

[565] S. E. Hitchcook. Book review: Statistical Analysis, Ideas and Methods, by E. Vernon Lewis. Journal of the Royal Statistical Society. Series A (General), 127(1):126–127, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982298.

Kingman:1964:BRB

[566] J. F. C. Kingman. Book review: Probability Theory, by Michel Loève. Journal of the Royal Statistical Society. Series A (General), 127(1):127–128, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982299.

Cochran:1964:BRB

[567] W. G. Cochran. Book review: Basic Ideas of Scientific Sampling, by Alan Stuart. Journal of the Royal Statistical Society. Series A (General), 127 (1):128–129, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982300.

Kalton:1964:BRB

[568] G. Kalton. Book review: Chance, Luck and Statistics, by Horace C. Levinson. Journal of the Royal Statistical Society. Series A (General), 127(1):129– 130, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982301.

Lomnicki:1964:BRB

[569] Z. A. Lomnicki. Book review: Statistical Theory of Reliability, by Mathematics Research Center. Journal of the Royal Statistical Society. Series A (General), 127(1):130–131, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982302.

Walters:1964:BRB

[570] A. A. Walters. Book review: Mathematical Models in the Social Sciences, by John G. Kemeny; J. Laurie Snell. Journal of the Royal Statistical Society. Series A (General), 127(1):131– 132, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982303.

Miller:1964:BRB

[571] A. J. Miller. Book review: Mathematical Theories of Traffic Flow, by Frank A. Haight. Journal of the Royal Statistical Society. Series A (General), 127 (1):132, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982304.

Barton:1964:BRB

[572] D. E. Barton. Book review: An Index of Mathematical Tables, by A. Fletcher; J. C. P. Miller; L. Rosenhead; L. J. Comrie. Journal of the Royal Statistical Society. Series A (General), 127 (1):133–134, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982305.

Vajda:1964:BRBa

[573] S. Vajda. Book review: Tables Describing Small-Sample Properties of the Mean, Median, Standard Deviation and Other Statistics in Sampling from Various Distributions, by Churchill Eisenhart. Journal of the Royal Statistical Society. Series A (General), 127(1): 134, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982306.

Rowntree:1964:BRBa

[574] J. A. Rowntree. Book review: U.S. Bureau of the Census. The Current Population Survey, Report on Methodology. Journal of the Royal Statistical Society. Series A (General), 127(1):134– 135, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982307.

Rowntree:1964:BRBb

[575] J. A. Rowntree. Book review: U.S. Bureau of the Census. The Current Population Survey Reinterview Program, Some Notes and Discussion. Journal of the Royal Statistical Society. Series A (General), 127(1):135–136, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982308.

Martin:1964:BRBa

[576] Leslie V. Martin. Book review: The Chester Beatty Research Institute Serial Abridged Life Tables, England and Wales 1841–1960, by R. A. M. Case; Christine Coghill; Joyce L. Harley; Joan T. Pearson. Journal of the Royal Statistical Society. Series A (General), 127(1):136–137, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982309.

Grebenik:1964:BRBa

[577] E. Grebenik. Book review: So Much Alive. The Life and Work of Wladimir S. Woytinski, by Emma S. Woytinski. Journal of the Royal Statistical Society. Series A (General), 127(1):137, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982310.

Hills:1964:BRBa

[578] M. Hills. Book review: Elements of Medical Statistics, by J. V. Smart. Journal of the Royal Statistical Society. Series A (General), 127(1):137– 138, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982311.

Grebenik:1964:BRBb

[579] E. Grebenik. Book review: The Use of Vital and Health Statistics for Genetic and Radiation Studies, by United Nations and the World Health Organization. Journal of the Royal Statistical Society. Series A (General), 127 (1):138–139, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982312.

Martin:1964:BRBb

[580] J. P. Martin. Book review: Persistent Criminals, by W. H. Hammond; Edna Chayen. Journal of the Royal Statistical Society. Series A (General), 127 (1):139, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982313.

Kemsley:1964:BRBa

[581] W. F. F. Kemsley. Book review: Household Survey Carried out in the Urban Areas of Greece During 1957– 58. Journal of the Royal Statistical Society. Series A (General), 127(1):140, ???? 1964. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2982314.

Corlett:1964:BRB

[582] W. J. Corlett. Book review: Research Methods in Economics and Business, by Robert Ferber; P. J. Verdoorn. Journal of the Royal Statistical Society. Series A (General), 127(1):140–141, ???? 1964. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982315.

Geary:1964:BRB

[583] R. C. Geary. Book review: Structural Interdependence and Economic Development, by Tibor Barna; William I. Abraham; Zoltan Kenessey. Journal of the Royal Statistical Society. Series A (General), 127(1):141–143, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982316.

Singer:1964:BRB

[584] H. W. Singer. Book review: Industrial Growth and World Trade: World Trends in Production, Consumption and Trade in Manufactures, by A. Maizels. Journal of the Royal Statistical Society. Series A (General), 127 (1):144–146, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982317.

Deane:1964:BRB

[585] Phyllis Deane. Book review: Economic Backwardness in Historical Perspective; A Book of Essays, by Alexander Gerschenkron. Journal of the Royal Statistical Society. Series A (General), 127(1):146–147, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982318.

Saunders:1964:BRBa

[586] C. T. Saunders. Book review: Britain 1984: A Forecast Prepared for Unilever, by Ronald Brech. Journal of the Royal Statistical Society. Series A (General), 127(1):147–149, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982319.

Holmans:1964:BRB

[587] A. E. Holmans. Book review: Capital and Rates of Return in Manufacturing Industries, by George J. Stigler.

Journal of the Royal Statistical Society. Series A (General), 127(1):149, ???? 1964. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982320.

Anonymous:1964:CNa

[588] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 127(1):150–159, ???? 1964. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982321.

Anonymous:1964:LAa

[589] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 127(1):160–164, ???? 1964. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982322.

Anonymous:1964:BMa

[590] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 127(1):i–vi, ???? 1964. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982323.

Anonymous:1964:FMb

[591] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 127(2):??, ???? 1964. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344001.

Benjamin:1964:PIM

[592] B. Benjamin and R. Ash. Prescribing information and management of the n.h.s. pharmaceutical services. Journal of the Royal Statistical Society. Series A (General), 127(2):165–198, ????

1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344002.

Yates:1964:HSW

[593] F. Yates and M. J. R. Healy. How should we reform the teaching of statistics? Journal of the Royal Statistical Society. Series A (General), 127(2): 199–210, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344003.

Bishop:1964:TGS

[594] H. E. Bishop. The training of government statisticians. Journal of the Royal Statistical Society. Series A (General), 127(2):211–215, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344004.

Deane:1964:TSE

[595] Marjorie Deane. The training of statisticians for economics and business. Journal of the Royal Statistical Society. Series A (General), 127(2):216– 218, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344005.

Spicer:1964:TMS

[596] C. C. Spicer. The training of medical statisticians. Journal of the Royal Statistical Society. Series A (General), 127 (2):219–233, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344006.

Horowitz:1964:NHP

[597] Ira Horowitz. A note on the hart-Paris measure of changes of business concentration. Journal of the Royal Statistical Society. Series A (General), 127 (2):234–237, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344007.

Maxwell:1964:CML

[598] A. E. Maxwell. Calculating maximumlikelihood factor loadings. Journal of the Royal Statistical Society. Series A (General), 127(2):238–241, ???? 1964. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344008.

Pratten:1964:CFA

[599] C. F. Pratten. A comparison of forecast and actual steel production in 1962. Journal of the Royal Statistical Society. Series A (General), 127(2):242– 250, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344009.

Amey:1964:DMB

[600] L. R. Amey. Diversified manufacturing business. Journal of the Royal Statistical Society. Series A (General), 127 (2):251–290, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344010.

Kendall:1964:BRB

[601] M. G. Kendall. Book review: Proceedings of the Symposium on Time Series Analysis Held at Brown University, June 11–14, 1962, by Murray Rosenblatt. Journal of the Royal Statistical Society. Series A (General), 127 (2):291–292, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344011.

Healy:1964:BRBa

[602] M. J. R. Healy. Book review: Smoothing, Forecasting and Prediction of Discrete Time Series, by R. G. Brown. Journal of the Royal Statistical Society. Series A (General), 127(2):292– 293, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344012.

Vajda:1964:BRBb

[603] S. Vajda. Book review: Linear Programming and Extensions, by George B. Dantzig. Journal of the Royal Statistical Society. Series A (General), 127 (2):294–295, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344013.

Curnow:1964:BRB

[604] R. N. Curnow. Book review: The Method of Paired Comparisons, by H. A. David. Journal of the Royal Statistical Society. Series A (General), 127 (2):295, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344014.

Bartlett:1964:BRBa

[605] M. S. Bartlett. Book review: Factor Analysis as a Statistical Method, by D. N. Lawley; A. E. Maxwell. Journal of the Royal Statistical Society. Series A (General), 127(2):295–296, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344015.

Jenkins:1964:BRB

[606] G. M. Jenkins. Book review: Introduction to Scientific Inference, by Robert Hooke. Journal of the Royal Statistical Society. Series A (General), 127 (2):296–297, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344016.

Gilchrist:1964:BRB

[607] W. G. Gilchrist. Book review: Probability and Statistics for Everyman: How to Understand and Use the Laws of Chance, by Irving Adler. Journal of the Royal Statistical Society. Series A (General), 127(2):297–298, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344017.

Ilersic:1964:BRB

[608] A. R. Ilersic. Book review: Descriptive Statistics, by Freda Conway. Journal of the Royal Statistical Society. Series A (General), 127(2):298–299, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344018.

Armitage:1964:BRBa

[609] P. Armitage. Book review: Bibliography of Basic Texts and Monographs on Statistical Methods, 1945–1960, by William R. Buckland; Ronald A. Fox. Journal of the Royal Statistical Society. Series A (General), 127(2):299, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344019.

Vajda:1964:BRBc

[610] S. Vajda. Book review: Methods and Models of Operations Research, by Arnold Kaufmann. Journal of the Royal Statistical Society. Series A (General), 127(2):299–300, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344020.

Healy:1964:BRBb

[611] M. J. R. Healy. Book review: An Introduction to ALGOL Programming, by R. Wooldridge; J. F. Ractliffe. Journal of the Royal Statistical Society. Series A (General), 127(2):300, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344021.

Loynes:1964:BRB

[612] R. M. Loynes. Book review: Probabilities on Algebraic Structures, by Ulf Grenander. Journal of the Royal Statistical Society. Series A (General), 127 (2):300–302, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344022.

Wetherill:1964:BRB

[613] G. B. Wetherill. Book review: Tables of Random Permutations, by Lincoln E. Moses; Robert E. Oakford. Journal of the Royal Statistical Society. Series A (General), 127(2):302, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344023.

Armitage:1964:BRBb

[614] P. Armitage. Book review: Tables for Testing Significance in a 2 × 2 Contingency Table, by D. J. Finney; R. Latscha; B. M. Bennett; P. Hsu; E. S. Pearson. Journal of the Royal Statistical Society. Series A (General), 127 (2):303, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344024.

Benjamin:1964:BRB

[615] B. Benjamin. Book review: Expert Committee on Health Statistics, Eighth Report. Journal of the Royal Statistical Society. Series A (General), 127 (2):303–304, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344025.

Freeman:1964:BRBa

[616] C. Freeman. Book review: Census 1961, Great Britain, Scientific and Technological Qualifications. Journal of the Royal Statistical Society. Series A (General), 127(2):304–305, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344026.

Saunders:1964:BRBb

[617] C. T. Saunders. Book review: Productivity in British Industry, by K. G. B. Bakewell; Graham Hutton. Journal of the Royal Statistical Society. Series A (General), 127(2):305–306, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344027.

Streeten:1964:BRBa

[618] Paul Streeten. Book review: International Trade Theory in a Developing World, by Roy Harrod; Douglas Hague. Journal of the Royal Statistical Society. Series A (General), 127 (2):306–307, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344028.

Streeten:1964:BRBb

[619] Paul Streeten. Book review: Pricing and Employment in the Trade Cycle: A Study of British Manufacturing Industry, 1950–61, by R. R. Neild. Journal of the Royal Statistical Society. Series A (General), 127(2):307–308, ????

1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344029.

Freeman:1964:BRBb

[620] C. Freeman. Book review: Research and Development: Its Growth and Composition, by N. E. Terleckyj. Journal of the Royal Statistical Society. Series A (General), 127(2):308–309, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344030.

Anonymous:1964:CLO

[621] Anonymous. The calculus of linguistic observations. Journal of the Royal Statistical Society. Series A (General), 127(2):309–310, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344031.

Anonymous:1964:CNb

[622] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 127(2):311–319, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344032.

George:1964:JRH

[623] R. F. George. John Reginald Hugh Shaul, 1897–1963. Journal of the Royal Statistical Society. Series A (General), 127(2):320–322, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344033.

Anonymous:1964:LAb

[624] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 127(2):323–326, ????

1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/4130656.

Anonymous:1964:BMb

[625] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 127(2):i–vi, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344034.

Anonymous:1964:FMc

[626] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 127(3):??, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343588.

Dean:1964:SFC

[627] Geoffrey Dean and J. O. Irwin. The stock of fixed capital in the United Kingdom in 1961. Journal of the Royal Statistical Society. Series A (General), 127(3):327–358, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343589.

Kendall:1964:SAL

[628] M. G. Kendall and J. D. Murchland. Statistical aspects of the legality of gambling. Journal of the Royal Statistical Society. Series A (General), 127 (3):359–391, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343590.

McArthur:1964:SFI

[629] Margaret McArthur. Some factors involved in estimating calorie requirements, with special reference to persons engaged in agricultural labour in Asian countries. Journal of the Royal Statistical Society. Series A (General), 127 (3):392–408, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343591.

Somers:1964:SMA

[630] Robert H. Somers. Simple measures of association for the triple dichotomy. Journal of the Royal Statistical Society. Series A (General), 127(3):409– 415, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343592.

Cashen:1964:SPS

[631] R. O. Cashen. A study of the prediction of success in selecting artificer apprentices for the Royal Navy. Journal of the Royal Statistical Society. Series A (General), 127(3):416–431, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343593.

Anonymous:1964:WP

[632] Anonymous. Wholesale prices in 1963. Journal of the Royal Statistical Society. Series A (General), 127(3):432– 437, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343594.

Irwin:1964:BRB

[633] J. O. Irwin. Book review: The Causation of Bus Drivers’ Accidents. Journal of the Royal Statistical Society. Series A (General), 127(3):438–451, ????

1964. URL http://www.jstor.org/
stable/2343595.

Silvey:1964:BRB

[634] S. D. Silvey. Book review: Statistical Theory, by B. W. Lindgren. Journal of the Royal Statistical Society. Series A (General), 127(3):452–453, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343596.

Maxwell:1964:BRBa

[635] A. E. Maxwell. Book review: Statistical Estimation in Factor Analysis: A New Technique and Its Foundation, by K. G. Jöreskog. Journal of the Royal Statistical Society. Series A (General), 127(3):453–454, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343597.

Good:1964:BRB

[636] I. J. Good. Book review: Information Theory and Coding, by Norman Abramson. Journal of the Royal Statistical Society. Series A (General), 127 (3):454–455, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343598.

Hills:1964:BRBb

[637] M. Hills. Book review: Tables of the Negative Binomial Probability Distribution, by Eric Williamson; Michael H. Bretherton. Journal of the Royal Statistical Society. Series A (General), 127 (3):455, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343599.

Beaumont:1964:BRB

[638] G. P. Beaumont. Book review: Basic Concepts of Probability and Statistics, by E. L. Lehmann; J. L. Hodges. Journal of the Royal Statistical Society. Series A (General), 127(3):455–456, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343600.

Revell:1964:BRB

[639] J. R. S. Revell. Book review: Studies in the National Balance Sheet of the United States, by Raymond W. Goldsmith; Robert E. Lipsey; Morris Mendelson. Journal of the Royal Statistical Society. Series A (General), 127 (3):456–457, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343601.

Vajda:1964:BRBd

[640] S. Vajda. Book review: Méthodes et Modèles de la Recherche Opérationnelle (Les Mathématiques de I’Entreprise), Tome 2, by A. Kaufmann. Journal of the Royal Statistical Society. Series A (General), 127(3):457–458, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343602.

Vajda:1964:BRBe

[641] S. Vajda. Book review: The Art of Simulation, by K. D. Tocher. Journal of the Royal Statistical Society. Series A (General), 127(3):458, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343603.

Butler:1964:BRB

[642] D. E. Butler. Book review: The English Voter, by A. J. Allen. Journal of the Royal Statistical Society. Series A (General), 127(3):458–459, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343604.

Seear:1964:BRB

[643] B. N. Seear. Book review: Coalminers’ Attendance at Work: The Report of an Investigation by a National Coal Board Research Team, by R. B. Buzzard; F. D. K. Liddell. Journal of the Royal Statistical Society. Series A (General), 127 (3):459–460, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343605.

Kaser:1964:BRB

[644] M. C. Kaser. Book review: Multi-Sector Economic Accounts, by D. I. Oparin; P. F. Knightsfield; K. J. Lancaster. Journal of the Royal Statistical Society. Series A (General), 127 (3):460–461, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343606.

Bartlett:1964:BRBb

[645] M. S. Bartlett. Book review: The Theory of Branching Processes, by Theodore E. Harris. Journal of the Royal Statistical Society. Series A (General), 127(3):461–462, ???? 1964. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343607.

Taylor:1964:BRB

[646] Wallis Taylor. Book review: Birmingham City, Central Statistical Office: Abstract of Statistics, No. 7, 1962, by Alan B. Neale. Journal of the Royal Statistical Society. Series A (General), 127(3):462–463, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343608.

Routh:1964:BRBa

[647] Guy Routh. Book review: White-Collar Redundancy: A Case Study, by Dorothy Wedderburn. Journal of the Royal Statistical Society. Series A (General), 127(3):463, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343609.

Anonymous:1964:CNc

[648] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 127(3):464–465, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343610.

Anonymous:1964:NC

[649] Anonymous. Norman Crump. Journal of the Royal Statistical Society. Series A (General), 127(3):466, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343611.

Anonymous:1964:LAc

[650] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 127(3):467–471, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343612.

Anonymous:1964:BMc

[651] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 127(3):i–vi, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343613.

Anonymous:1964:VIb

[652] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 127(4):605– 608, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343931.

Anonymous:1964:FMd

[653] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 127(4):??, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343932.

Moser:1964:PSH

[654] C. A. Moser, P. R. G. Layard, and J. O. Irwin. Planning the scale of higher education in Britain: Some statistical problems. Journal of the Royal Statistical Society. Series A (General), 127 (4):473–526, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343933.

Thom:1964:LUL

[655] A. Thom. The larger units of length of megalithic man. Journal of the Royal Statistical Society. Series A (General), 127(4):527–533, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343934.

Winkler:1964:RBC

[656] Wilhelm Winkler. Relations between crude and life tables death rates. Journal of the Royal Statistical Society. Series A (General), 127(4):534–543, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343935.

Anonymous:1964:ARC

[657] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 127 (4):544–565, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343936.

Anonymous:1964:POH

[658] Anonymous. Proceedings of the one hundred and thirtieth Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 127(4):566, ???? 1964. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343937.

Kerridge:1964:BRB

[659] D. Kerridge. Book review: Introduction to Statistical Method, by Sylvain Ehrenfeld; Sebastian Littauer. Journal of the Royal Statistical Society. Series A (General), 127(4):567–568, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343938.

Maxwell:1964:BRBb

[660] James Maxwell. Book review: Statistical Research Methods in Education and Psychology, by Norman K. Henderson. Journal of the Royal Statistical Society. Series A (General), 127(4):568, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343939.

Levison:1964:BRB

[661] M. Levison. Book review: Computer Programming and Autocodes, by D. G. Burnett-Hall; L. A. G. Dresel; P. A.

Samet. Journal of the Royal Statistical Society. Series A (General), 127 (4):569, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343940.

Armitage:1964:BRBc

[662] P. Armitage. Book review: Stochastic Models in Medicine and Biology, by J. Gurland. Journal of the Royal Statistical Society. Series A (General), 127 (4):569–570, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343941.

Bartholomew:1964:BRB

[663] D. J. Bartholomew. Book review: Linear Programming: Methods and Applications, by Saul I. Gass. Journal of the Royal Statistical Society. Series A (General), 127(4):570, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343942.

Page:1964:BRB

[664] E. S. Page. Book review: Monte Carlo Methods, by J. M. Hammersley; D. C. Handscomb. Journal of the Royal Statistical Society. Series A (General), 127 (4):570–571, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343943.

Davies:1964:BRBa

[665] Hilda M. Davies. Book review: Basic Statistics for Business and Economics, by Dick A. Leabo; C. Frank Smith. Journal of the Royal Statistical Society. Series A (General), 127(4):571– 572, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343944.

Fisher:1964:BRB

[666] Gordon Fisher. Book review: Méthodes Statistiques de l’Économetrie, by Edmond Malinvaud. Journal of the Royal Statistical Society. Series A (General), 127(4):572–573, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343945.

Leser:1964:BRBa

[667] C. E. V. Leser. Book review: A Contribution to Demand Analysis, by I. F. Pearce; Théorie des Choix et Fonctions de Demande. by André Nataf. Journal of the Royal Statistical Society. Series A (General), 127(4):573–574, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343946.

Clark:1964:BRBa

[668] Colin Clark. Book review: A Guide to Australian Economic Statistics, by George R. Palmer. Journal of the Royal Statistical Society. Series A (General), 127(4):574–575, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343947.

Clark:1964:BRBb

[669] Colin Clark. Book review: Economic Development with Special Reference to East Asia, by Kenneth Berrill. Journal of the Royal Statistical Society. Series A (General), 127(4):575–576, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343948.

Clark:1964:BRBc

[670] Colin Clark. Book review: A Programme for Growth, 4. Capital Output and Employment 1948–60. Journal of the Royal Statistical Society. Series A (General), 127(4):577–578, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343949.

Davies:1964:BRBb

[671] R. W. Davies. Book review: Planning and Statistics in Socialist Countries. Journal of the Royal Statistical Society. Series A (General), 127(4):578, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343950.

Hawtrey:1964:BRB

[672] Ralph Hawtrey. Book review: Sterling in the Sixties, by Christopher McMahon. Journal of the Royal Statistical Society. Series A (General), 127(4): 578–579, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343951.

Dickinson:1964:BRB

[673] H. D. Dickinson. Book review: The Economic Theory of ‘Managerial’ Capitalism, by Robin Marris. Journal of the Royal Statistical Society. Series A (General), 127(4):579–581, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343952.

Routh:1964:BRBb

[674] Guy Routh. Book review: The Economics of Capital Utilisation: A Report on Multiple-Shift Work, by Robin Marris; Ian Maclean; Simon Bernau. Journal of the Royal Statistical Society. Series A (General), 127(4):581, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343953.

Rybczynski:1964:BRB

[675] T. M. Rybczynski. Book review: The Finance and Analysis of Capital Projects, by A. J. Merrett; Allen Sykes. Journal of the Royal Statistical Society. Series A (General), 127(4):581– 582, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343954.

Anstey:1964:BRB

[676] Vera Anstey. Book review: Recent Trends in Union Finance: A Critique of the Budgetary Policy of the Government of India since 1935, by S. K. Mukherjee. Journal of the Royal Statistical Society. Series A (General), 127 (4):582–583, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343955.

Leser:1964:BRBb

[677] C. E. V. Leser. Book review: Models of Income Determination, by The Conference on Research in Income and Wealth. Journal of the Royal Statistical Society. Series A (General), 127 (4):583–584, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343956.

Barna:1964:BRB

[678] T. Barna. Book review: The Economics of Take-off into Sustained Growth. Proceedings of a Conference held by the International Economic Association, by W. W. Rostow. Journal of the Royal Statistical Society. Series A (General), 127(4):584–585, ????

1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343957.

Jahoda:1964:BRB

[679] Marie Jahoda. Book review: Adolescent Personality and Behavior: M.M.P.I. Patterns of Normal, Delinquent, Dropout, and other Outcomes, by Starke R. Hathaway; Elio D. Monachesi. Journal of the Royal Statistical Society. Series A (General), 127 (4):585–586, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343958.

Abrams:1964:BRBa

[680] Mark Abrams. Book review: Party and Society: Voting Behaviour in the Anglo-American Democracies, by Robert T. Alford. Journal of the Royal Statistical Society. Series A (General), 127(4):586–588, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343959.

Abrams:1964:BRBb

[681] M. A. Abrams. Book review: A Review of Elections, 1961–1962. Journal of the Royal Statistical Society. Series A (General), 127(4):588, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343960.

Kemsley:1964:BRBb

[682] W. F. F. Kemsley. Book review: IPA National Readship Surveys 1963. Tables Relating to January-December, 1963. Journal of the Royal Statistical Society. Series A (General), 127(4): 588–590, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343961.

Amey:1964:BRB

[683] L. R. Amey. Book review: Advanced Cost Accountancy, by J. E. Smith; J. C. W. Day. Journal of the Royal Statistical Society. Series A (General), 127 (4):590, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343962.

Anonymous:1964:CNd

[684] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 127(4):591–595, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343963.

Kiser:1964:PKW

[685] Clyde V. Kiser. Pascal Kidder Whelpton, 1893–1964. Journal of the Royal Statistical Society. Series A (General), 127(4):596–597, ???? 1964. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343964.

Pearson:1964:SSW

[686] E. S. Pearson. Samuel Stanley wilks, 1906–1964. Journal of the Royal Statistical Society. Series A (General), 127 (4):597–599, ???? 1964. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343965.

Anonymous:1964:LAd

[687] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 127(4):600–604, ???? 1964. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343966.

Anonymous:1964:BMd

[688] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 127(4):i–vi, ???? 1964. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343967.

Anonymous:1965:VIa

[689] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 128(1):??, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343433.

Anonymous:1965:FMa

[690] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 128(1):??, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343434.

Chambers:1965:SII

[691] S. Paul Chambers. Statistics and intellectual integrity. Journal of the Royal Statistical Society. Series A (General), 128(1):1–16, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343435.

Berrington:1965:GE

[692] H. B. Berrington. The general election of 1964. Journal of the Royal Statistical Society. Series A (General), 128(1):17– 66, ???? 1965. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2343436.

Beesley:1965:VLS

[693] M. E. Beesley and C. D. Foster. The Victoria Line: Social benefit and finances. Journal of the Royal Statistical Society. Series A (General), 128(1):67– 88, ???? 1965. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343437. See note [861]

and response [862].

Gini:1965:CIS

[694] Corrado Gini. On the characteristics of Italian statistics. Journal of the Royal Statistical Society. Series A (General), 128(1):89–109, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343438.

Dalenius:1965:MNS

[695] Tore Dalenius. The mode — a neglected statistical parameter. Journal of the Royal Statistical Society. Series A (General), 128(1):110–117, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343439.

Kemsley:1965:IVE

[696] W. F. F. Kemsley. Interviewer variability in expenditure surveys. Journal of the Royal Statistical Society. Series A (General), 128(1):118–139, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343440.

Jenkins:1965:BRB

[697] G. M. Jenkins. Book review: Prediction and Regulation, by P. Whittle. Journal of the Royal Statistical Society. Series A (General), 128(1):140–142, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343441.

Vajda:1965:BRBa

[698] S. Vajda. Book review: Einführung in die Wahrscheinlichkeitsrechnung und Mathematische Statistik, by Dietrich Morgenstern. Journal of the Royal Statistical Society. Series A (General), 128 (1):142, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343442.

Loynes:1965:BRB

[699] R. M. Loynes. Book review: Applications of Characteristic Functions, by E. Lukacs; R. G. Laha. Journal of the Royal Statistical Society. Series A (General), 128(1):142–143, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343443.

Tweedie:1965:BRB

[700] M. C. K. Tweedie. Book review: Principles of Random Walk, by F. Spitzer. Journal of the Royal Statistical Society. Series A (General), 128(1):143– 145, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343444.

Stone:1965:BRBa

[701] M. Stone. Book review: Miscellaneous Studies in Probability and Statistics: Distribution Theory, Small-Sample Problems and Occasional Tables, by The Statistical Engineering Laboratory, National Bureau of Standards. Journal of the Royal Statistical Society. Series A (General), 128(1): 145, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343445.

Armitage:1965:BRB

[702] P. Armitage. Book review: Statistical Method in Biological Assay, by D. J. Finney. Journal of the Royal Statistical Society. Series A (General), 128 (1):145–146, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343446.

Wetherill:1965:BRBa

[703] G. B. Wetherill. Book review: Tables of Natural and Common Logarithms to 110 Decimals, by W. E. Mansell; A. J. Thompson. Journal of the Royal Statistical Society. Series A (General), 128 (1):146, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343447.

Wetherill:1965:BRBb

[704] G. B. Wetherill. Book review: Statistical Tables for Biological, Agricultural and Medical Research, by R. A. Fisher; F. Yates. Journal of the Royal Statistical Society. Series A (General), 128 (1):146, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343448.

Conway:1965:BRBa

[705] Freda Conway. Book review: What do Statistics Show?, by F. R. Oliver. Journal of the Royal Statistical Society. Series A (General), 128(1):146–147, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343449.

Kalton:1965:BRB

[706] G. Kalton. Book review: Descriptive Statistics: A Programed Textbook. Volume I, by Lassar G. Gotkin; Leo S.

Goldstein. Journal of the Royal Statistical Society. Series A (General), 128 (1):147–148, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343450.

Beale:1965:BRBa

[707] E. M. L. Beale. Book review: Elements of Linear Programming with Economic Applications, by R. C. Geary; M. D. McCarthy. Journal of the Royal Statistical Society. Series A (General), 128 (1):148, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343451.

Oliver:1965:BRB

[708] F. R. Oliver. Book review: Consumer Credit Costs 1949–1959, by Paul F. Smith. Journal of the Royal Statistical Society. Series A (General), 128(1): 149, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343452.

Leser:1965:BRBa

[709] C. E. V. Leser. Book review: Priority Patterns and the Demand for Household Durable Goods, by F. G. Pyatt. Journal of the Royal Statistical Society. Series A (General), 128(1):149– 150, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343453.

Peston:1965:BRB

[710] M. H. Peston. Book review: The Role of Direct and Indirect Taxes in the Federal Revenue System: A Conference Report of the National Bureau of Economic Research and the Brookings Institution. Journal of the Royal Statistical Society. Series A (General), 128 (1):150–151, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343454.

Mays:1965:BRB

[711] John B. Mays. Book review: Offenders as Employees, by J. P. Martin. Journal of the Royal Statistical Society. Series A (General), 128(1):151–152, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343455.

Heasman:1965:BRB

[712] M. A. Heasman. Book review: Problems and Progress in Medical Care: Essays in Current Research, by Gordon McLachlan. Journal of the Royal Statistical Society. Series A (General), 128 (1):152–153, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343456.

Scammell:1965:BRBa

[713] W. M. Scammell. Book review: The International Monetary Fund: Retrospect and Prospect, by Shigeo Horie; Roy Harrod. Journal of the Royal Statistical Society. Series A (General), 128 (1):153–155, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343457.

McCarthy:1965:BRB

[714] M. J. McCarthy. Book review: Plans for the Reform of the International Monetary System, by Fritz Machlup. Journal of the Royal Statistical Society. Series A (General), 128(1):155– 156, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343458.

Brechling:1965:BRB

[715] Frank Brechling. Book review: A Macroeconomic Theory of Workable Competition, by G. E. Ferguson. Journal of the Royal Statistical Society. Series A (General), 128(1):156–157, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343459.

Anonymous:1965:CNa

[716] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 128(1):158–161, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343460.

Morrell:1965:RC

[717] A. J. H. Morrell. L. R. Connor. Journal of the Royal Statistical Society. Series A (General), 128(1):162, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343461.

Anonymous:1965:LAa

[718] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 128(1):163–167, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343462.

Anonymous:1965:BMa

[719] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 128(1):i–vi, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343463.

Anonymous:1965:FMb

[720] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 128(2):??, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344177.

Morton:1965:AGP

[721] A. Q. Morton. The authorship of Greek prose. Journal of the Royal Statistical Society. Series A (General), 128(2): 169–233, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344178. See er-

rata [813].

Cochran:1965:POS

[722] W. G. Cochran and S. Paul Chambers. The planning of observational studies of human populations. Journal of the Royal Statistical Society. Series A (General), 128(2):234–266, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344179.

Mackay:1965:VRL

[723] D. I. Mackay and N. K. Buxton. A view of regional labour statistics. Journal of the Royal Statistical Society. Series A (General), 128(2):267–284, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344180.

Oliver:1965:YYA

[724] F. R. Oliver. A year-by-year analysis of inter-regional migration, 1951– 1961. Journal of the Royal Statistical Society. Series A (General), 128(2): 285–287, ???? 1965. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344181.

Harrod:1965:MBE

[725] Roy Harrod. The management of the British economy. Journal of the Royal Statistical Society. Series A (General), 128(2):288–293, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344182.

Vajda:1965:BRBb

[726] S. Vajda. Book review: Algèbre Linéaire et Programmation Linéaire, by J. Acher. Journal of the Royal Statistical Society. Series A (General), 128 (2):294, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344183.

Brass:1965:BRB

[727] W. Brass. Book review: Basic Statistics: A Primer for the Biomedical Sciences, by Olive Jean Dunn. Journal of the Royal Statistical Society. Series A (General), 128(2):295, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344184.

Thomas:1965:BRBa

[728] Brinley Thomas. Book review: Investment in Australian Economic Development, 1861–1900, by N. G. Butlin. Journal of the Royal Statistical Society. Series A (General), 128(2):295– 297, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344185.

Walker:1965:BRB

[729] David Walker. Book review: Economic Development, by John Kenneth Galbraith. Journal of the Royal Statistical Society. Series A (General), 128 (2):297–298, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344186.

Thomas:1965:BRBb

[730] Brinley Thomas. Book review: Estimates of Residential Building, United States, 1840–1939, by Manuel Gottlieb. Journal of the Royal Statistical Society. Series A (General), 128(2):298– 299, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344187.

Walters:1965:BRB

[731] A. A. Walters. Book review: Experiments with Input-Output Models; An Application to the Economy of the United Kingdom, 1948–55, by A. Ghosh. Journal of the Royal Statistical Society. Series A (General), 128 (2):299–300, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344188.

Devons:1965:BRB

[732] E. Devons. Book review: British Export Performance: A Comparative Study, by S. J. Wells. Journal of the Royal Statistical Society. Series A (General), 128(2):300–301, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344189.

Martin:1965:BRBa

[733] J. P. Martin. Book review: Pentonville: A Sociological Study of an English Prison, by Terence Morris; Pauline Morris; Barbara Barer. Journal of the Royal Statistical Society. Series A (General), 128(2):301–302, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344190.

Kemsley:1965:BRB

[734] W. F. F. Kemsley. Book review: Kent Development Plan Quinquennial Review 1963. Journal of the Royal Statistical Society. Series A (General), 128 (2):302–303, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344191.

Lewis:1965:BRBa

[735] J. Parry Lewis. Book review: The Use of Mathematics in Economics, by V. S. Nemchinov; A. Nove. Journal of the Royal Statistical Society. Series A (General), 128(2):303–304, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344192.

David:1965:BRB

[736] F. N. David. Book review: Calcul des Probabilités en Vue des Applications, by M. Girault. Journal of the Royal Statistical Society. Series A (General), 128(2):305, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344193.

Hitchcock:1965:BRBa

[737] Shirley Hitchcock. Book review: Introduction to Experimental Statistics, by C. C. Li. Journal of the Royal Statistical Society. Series A (General), 128 (2):305–306, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344194.

Downton:1965:BRB

[738] F. Downton. Book review: Statistical Assessment of the Life Characteristic; a Bibliographic Guide, by W. R. Buckland. Journal of the Royal Statistical Society. Series A (General), 128 (2):306, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344195.

Vajda:1965:BRBc

[739] S. Vajda. Book review: Actes de la 3ème Conférence Internationale de Recherche Opérationnelle, by G. Kreweras; G. Morlat. Journal of the Royal Statistical Society. Series A (General), 128(2):306–307, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344196.

Beale:1965:BRBb

[740] E. M. L. Beale. Book review: Nonlinear and Dynamic Programming, by G. Hadley. Journal of the Royal Statistical Society. Series A (General), 128 (2):307–308, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344197.

Bartlett:1965:BRB

[741] M. S. Bartlett. Book review: An Introduction to Traffic Flow Theory, by D. L. Gerlough; D. G. Capelle. Journal of the Royal Statistical Society. Series A (General), 128(2):308–309, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344198.

Page:1965:BRB

[742] E. S. Page. Book review: Annual Review in Automatic Programming, by Richard Goodman. Journal of the Royal Statistical Society. Series A (General), 128(2):309, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344199.

Leser:1965:BRBb

[743] C. E. V. Leser. Book review: Econometric Theory, by Arthur S. Goldberger. Journal of the Royal Statistical Society. Series A (General), 128 (2):309–310, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344200.

Leser:1965:BRBc

[744] C. E. V. Leser. Book review: Econometric Model Building-Essays on the Causal Chain Approach, by Herman O. A. Wold. Journal of the Royal Statistical Society. Series A (General), 128 (2):310–311, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344201.

Anonymous:1965:CNb

[745] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 128(2):312–317, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344202.

Anonymous:1965:LAb

[746] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 128(2):318–320, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344203.

Anonymous:1965:BMb

[747] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 128(2):i–iv, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344204.

Anonymous:1965:FMc

[748] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 128(3):??, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344205.

Berman:1965:FFU

[749] Lawrence S. Berman. Flow of funds in the United Kingdom. Journal of the Royal Statistical Society. Series A (General), 128(3):321–360, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344206.

Barnett:1965:SPC

[750] V. D. Barnett and H. F. Ross. Statistical properties of computer acceptance tests. Journal of the Royal Statistical Society. Series A (General), 128 (3):361–394, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344207.

Armstrong:1965:SPS

[751] Alan Armstrong and Aubrey Silberton. Size of plant, size of enterprise and concentration in British manufacturing industry 1935–58. Journal of the Royal Statistical Society. Series A (General), 128(3):395–420, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344208.

Swee-Hock:1965:MTM

[752] Saw Swee-Hock. Malaya: Tables of male working life, 1957. Journal of the Royal Statistical Society. Series A (General), 128(3):421–438, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344209.

Nisbet:1965:BRB

[753] John Nisbet. Book review: Half Our Future. Journal of the Royal Statistical Society. Series A (General), 128 (3):439–441, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344210.

Stone:1965:BRBb

[754] M. Stone. Book review: Studies in Subjective Probability, by Henry E. Kyburg; Howard E. Smokler. Journal of the Royal Statistical Society. Series A (General), 128(3):442–443, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344211.

Vajda:1965:BRBd

[755] S. Vajda. Book review: Théorie des Chaînes de Markov Finies et ses Applications, by Patrick Gordon. Journal of the Royal Statistical Society. Series A (General), 128(3):443, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344212.

Cox:1965:BRB

[756] D. R. Cox. Book review: Multivariate Statistical Analysis for Biologists, by H. L. Seal. Journal of the Royal Statistical Society. Series A (General), 128 (3):443–444, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344213.

Hawkes:1965:BRB

[757] A. G. Hawkes. Book review: Time-Dependent Results in Storage Theory, by N. U. Prabhu. Journal of the Royal Statistical Society. Series A (General), 128(3):444–445, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344214.

Vajda:1965:BRBe

[758] S. Vajda. Book review: Les Problèmes d’Ordonnancement: Applications et Methodes, by D. Carré. Journal of the Royal Statistical Society. Series A (General), 128(3):445, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344215.

Farlie:1965:BRB

[759] D. J. G. Farlie. Book review: Cumulative Sum Techniques, by R. H. Woodward; P. L. Goldsmith. Journal of the Royal Statistical Society. Series A (General), 128(3):445–446, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344216.

Pearce:1965:BRBa

[760] I. F. Pearce. Book review: Consumer Discretionary Behaviour: A Comparative Study in Alternative Methods of Empirical Research, by Lewis Schipper. Journal of the Royal Statistical Society. Series A (General), 128(3):446, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344217.

Conway:1965:BRBb

[761] Freda Conway. Book review: Statistics, by A. R. Ilersic. Journal of the Royal Statistical Society. Series A (General), 128(3):446, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344218.

Mills:1965:BRB

[762] G. Mills. Book review: Quadratic Programming, by J. C. G. Boot. Journal of the Royal Statistical Society. Series A (General), 128(3):447, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344219.

Granger:1965:BRB

[763] C. W. J. Granger. Book review: The Random Character of Stock Market Prices, by Paul H. Cootner. Journal of the Royal Statistical Society. Series A (General), 128(3):447–449, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344220.

Vajda:1965:BRBf

[764] S. Vajda. Book review: Recherche Opérationnelle et Problèmes du Tiers-Monde. Journal of the Royal Statistical Society. Series A (General), 128 (3):449–450, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344221.

Pearce:1965:BRBb

[765] I. F. Pearce. Book review: Anticipations and Purchases, by F. Thomas Juster. Journal of the Royal Statistical Society. Series A (General), 128 (3):450, ???? 1965. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2344222.

Ray:1965:BRB

[766] W. D. Ray. Book review: Short-Term Forecasting, by G. A. Coutie; O. L. Davies. Journal of the Royal Statistical Society. Series A (General), 128 (3):451, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344223.

Lupton:1965:BRBa

[767] Tom Lupton. Book review: Economics and Sociology of Industry, by P. Sargant Florence. Journal of the Royal Statistical Society. Series A (General), 128(3):452–453, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344224.

Streeten:1965:BRB

[768] Paul Streeten. Book review: Postwar Economic Growth: Four Lectures, by Simon Kuznets. Journal of the Royal Statistical Society. Series A (General), 128(3):453–454, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344225.

Lewis:1965:BRBb

[769] J. Parry Lewis. Book review: A Programme for Growth, 5. The Model in its Environment. Journal of the Royal Statistical Society. Series A (General), 128(3):454–455, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344226.

Sarkar:1965:BRB

[770] N. K. Sarkar. Book review: Influence of Foreign Trade on the Level and Growth of the National Income of Ceylon, 1926–1954, by Warnasena Rasaputram. Journal of the Royal Statistical Society. Series A (General), 128 (3):455, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344227.

Sukhatme:1965:BRB

[771] P. V. Sukhatme. Book review: Trends in Wheat Consumption. Journal of the Royal Statistical Society. Series A (General), 128(3):455–457, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344228.

Eastham:1965:BRB

[772] J. K. Eastham. Book review: The National Income of Iraq, 1953–1960, by K. Haseeb. Journal of the Royal Statistical Society. Series A (General), 128 (3):457–458, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344229.

Anonymous:1965:CNc

[773] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 128(3):459–461, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344230.

Hill:1965:EW

[774] Austin Bradford Hill. The Earl of Woolton, 1883–1964. Journal of the Royal Statistical Society. Series A (General), 128(3):462–463, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344231.

Anonymous:1965:LAc

[775] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 128(3):464–467, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344232.

Anonymous:1965:EOA

[776] Anonymous. Errata: Optimum allocation in multivariate surveys. Journal of the Royal Statistical Society. Series A (General), 128(3):468, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344233. See [525].

Anonymous:1965:BMc

[777] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 128(3):i–vii, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344234.

Anonymous:1965:VIb

[778] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 128(4):624– 627, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343464.

Anonymous:1965:FMd

[779] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 128(4):??, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343465.

Smith:1965:PPS

[780] Cedric A. B. Smith. Personal probability and statistical analysis. Journal of the Royal Statistical Society. Series A (General), 128(4):469–499, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343466.

Schneider:1965:CPT

[781] J. R. L. Schneider. The census pre-test of 1964. Journal of the Royal Statistical Society. Series A (General), 128 (4):500–533, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343467.

Burman:1965:MSA

[782] J. P. Burman. Moving seasonal adjustment of economic time series. Journal of the Royal Statistical Society. Series A (General), 128(4):534–558, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/

stable/2343468. See errata [852].

Anonymous:1965:WP

[783] Anonymous. Wholesale prices in 1964. Journal of the Royal Statistical Society. Series A (General), 128(4):559– 564, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343469.

Anonymous:1965:ARC

[784] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 128 (4):565–585, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343470.

Anonymous:1965:POH

[785] Anonymous. Proceedings of the one hundred and thirty-first Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 128 (4):586–587, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343471.

Hicks:1965:BRB

[786] Ursula K. Hicks. Book review: Report of the Committee of Inquiry Into the Impact of Rates on Households, by R. G. D. Allen. Journal of the Royal Statistical Society. Series A (General), 128 (4):588–592, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343472.

Barton:1965:BRBa

[787] D. E. Barton. Book review: Handbook of Mathematical Functions with Formulas, Graphs and Mathematical Tables, by M. Abramovitz; I. A. Stegun. Journal of the Royal Statistical Society. Series A (General), 128(4):593– 594, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343473.

Smith:1965:BRB

[788] Cedric A. B. Smith. Book review: The Elements of Stochastic Processes with Applications to the Natural Sciences, by Norman T. J. Bailey. Journal of the Royal Statistical Society. Series A (General), 128(4):594, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343474.

Barton:1965:BRBb

[789] D. E. Barton. Book review: Applied Combinatorial Mathematics, by Edwin F. Beckenbach. Journal of the Royal Statistical Society. Series A (General), 128(4):595, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343475.

Barton:1965:BRBc

[790] D. E. Barton. Book review: Inequalities on Distribution Functions, by H. J. Godwin. Journal of the Royal Statistical Society. Series A (General), 128 (4):595–596, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343476.

Barton:1965:BRBd

[791] D. E. Barton. Book review: Geometrical Probability, by M. G. Kendall; P. A. P. Moran. Journal of the Royal Statistical Society. Series A (General), 128 (4):596–597, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343477.

Kingman:1965:BRB

[792] J. F. C. Kingman. Book review: Convex Transformations of Random Variables, by W. R. Van Zwet. Journal of the Royal Statistical Society. Series A (General), 128(4):597–598, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343478.

Stuart:1965:BRB

[793] A. Stuart. Book review: Handbook of Nonparametric Statistics, II: Results for Two and Several Sample Problems, Symmetry and Extremes, by John E.

Walsh. Journal of the Royal Statistical Society. Series A (General), 128 (4):598, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343479.

Wetherill:1965:BRBc

[794] G. B. Wetherill. Book review: Tracts for Computers, by M. S. Bartlett. Journal of the Royal Statistical Society. Series A (General), 128(4):598, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343480.

Wetherill:1965:BRBd

[795] G. B. Wetherill. Book review: Tables of the Function f(z) = ∫_{0}^{z} dx in the Complex Domain, by K. A. Karpov; D. E. Brown; Tables of the Function w(z) = e^{−z^{2}} ∫_{0}^{z} e^{x^{2}} dx in the Complex Domain, by K. A. Karpov; D. E. Brown; Tables of the Normal Probability Integral, the Normal Density and its Normalised Derivatives, by N. V. Smirnov; D. E. Brown. Journal of the Royal Statistical Society. Series A (General), 128 (4):599, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343481.

Maxwell:1965:BRB

[796] A. E. Maxwell. Book review: An Introduction to Psychological Statistics, by Philip H. DuBois. Journal of the Royal Statistical Society. Series A (General), 128(4):599, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343482.

Conway:1965:BRBc

[797] Freda Conway. Book review: Statistics in Theory and Practice with Special Reference to Published Statistics, by L. R. Connor; A. J. H. Morrell. Journal of the Royal Statistical Society. Series A (General), 128(4):599–600, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343483.

Hitchcock:1965:BRBb

[798] Shirley Hitchcock. Book review: Statistics and Experimental Design in Engineering and the Physical Sciences, by Norman L. Johnson; Fred C. Leone. Journal of the Royal Statistical Society. Series A (General), 128(4):600– 601, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343484.

Barna:1965:BRB

[799] T. Barna. Book review: International Association for Research in Income and Wealth Income Redistribution and the Statistical Foundations of Economic Policy Income and Wealth: Series X. Income Redistribution and the Statistical Foundation of Economic Policy, by Colin Clark; G. Stuvel. Journal of the Royal Statistical Society. Series A (General), 128(4):601, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343485.

Conway:1965:BRBd

[800] Freda Conway. Book review: Index Numbers: Theory and Applications, by Walter R. Crowe. Journal of the Royal Statistical Society. Series A (General), 128(4):601–602, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343486.

Mitchell:1965:BRB

[801] B. R. Mitchell. Book review: Railroads and American Economic Growth: Essays in Econometric History, by Robert W. Fogel. Journal of the Royal Statistical Society. Series A (General), 128(4):602–603, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343487.

Scammell:1965:BRBb

[802] W. M. Scammell. Book review: The “Band” Proposal: The Limits of Permissible Exchange Rate Variations, by George N. Halm. Journal of the Royal Statistical Society. Series A (General), 128(4):603–605, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343488.

Feinstein:1965:BRB

[803] C. H. Feinstein. Book review: Studies in Profit, Business Saving and Investment in the United Kingdom, 1920– 1962, by P. E. Hart; James Bates. Journal of the Royal Statistical Society. Series A (General), 128(4):605–606, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343489.

Spraos:1965:BRB

[804] John Spraos. Book review: The Theory of Trade and Protection, by W. P. Travis. Journal of the Royal Statistical Society. Series A (General), 128 (4):607–608, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343490.

Saunders:1965:BRB

[805] Christopher Saunders. Book review: Glossary of Terms in Official Statistics, by J. W. Nixon. Journal of the Royal Statistical Society. Series A (General), 128(4):608, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343491.

Lupton:1965:BRBb

[806] T. Lupton. Book review: Repercussions of Redundancy. A. Local Survey, by Hilda R. Kahn. Journal of the Royal Statistical Society. Series A (General), 128(4):609, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343492.

Martin:1965:BRBb

[807] J. P. Martin. Book review: Social Deviance: Social Policy, Action and Research, by Leslie T. Wilkins. Journal of the Royal Statistical Society. Series A (General), 128(4):610, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343493.

Anonymous:1965:CNd

[808] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 128(4):611–613, ???? 1965. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343494.

Davies:1965:VSN

[809] R. W. Davies and G. R. Barker. Vasilii Sergeevich Nemchinov, 1894– 1964. Journal of the Royal Statistical Society. Series A (General), 128(4): 614–615, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343495.

Grebenik:1965:ECR

[810] E. Grebenik. Edmund Cecil Rhodes, 1892–1964. Journal of the Royal Statistical Society. Series A (General), 128 (4):615–616, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343496.

Irwin:1965:EBW

[811] J. O. Irwin. Edwin Bidwell Wilson, 1879–1964. Journal of the Royal Statistical Society. Series A (General), 128 (4):616–618, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343497.

Anonymous:1965:LAd

[812] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 128(4):619–622, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343498.

Anonymous:1965:EAG

[813] Anonymous. Errata: The authorship of Greek prose. Journal of the Royal Statistical Society. Series A (General), 128(4):623, ???? 1965. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343499. See

[721].

Anonymous:1965:BMd

[814] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 128(4):i–vi, ???? 1965. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343500.

Anonymous:1966:VIa

[815] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 129(1):??, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343892.

Anonymous:1966:FMa

[816] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 129(1):??, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343893.

Beckerman:1966:EPD

[817] W. Beckerman. Economic planning and data requirements. Journal of the Royal Statistical Society. Series A (General), 129(1):1–5, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343894.

Wigley:1966:DNP

[818] K. J. Wigley. The development of the national plan. Journal of the Royal Statistical Society. Series A (General), 129(1):6–25, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343895.

Emmett:1966:DIE

[819] B. P. Emmett. The design of investigations into the effects of radio and television programmes and other mass communications. Journal of the Royal Statistical Society. Series A (General), 129(1):26–60, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343896.

Feldstein:1966:BVM

[820] Martin S. Feldstein. A binary variable multiple regression method of analysing factors affecting peri-Natal mortality and other outcomes of pregnancy. Journal of the Royal Statistical Society. Series A (General), 129(1):61– 73, ???? 1966. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343897.

Marris:1966:MFA

[821] R. Marris and A. Singh. A measure of a Firm’s average share price. Journal of the Royal Statistical Society. Series A (General), 129(1):74–97, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343898.

Armitage:1966:MOD

[822] P. Armitage, L. M. Blendis, and H. C. Smyllie. The measurement of observer disagreement in the recording of signs. Journal of the Royal Statistical Society. Series A (General), 129(1):98–109, ???? 1966. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343899.

Copas:1966:MCR

[823] J. B. Copas. Monte Carlo results for estimation in a stable Markov time series. Journal of the Royal Statistical Society. Series A (General), 129(1):110– 116, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343900.

Domb:1966:BRB

[824] C. Domb. Book review: Equilibrium Statistical Mechanics, by Frank C. Andrews. Journal of the Royal Statistical Society. Series A (General), 129(1): 117–118, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343901.

Miller:1966:BRB

[825] H. D. Miller. Book review: Statistical Theory and Metholodogy in Science and Engineering, by K. A. Brownlee. Journal of the Royal Statistical Society. Series A (General), 129(1):118, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343902.

Brillinger:1966:BRBa

[826] D. R. Brillinger. Book review: Spectral Analysis of Economic Time Series, by C. W. J. Granger; M. Hatanaka. Journal of the Royal Statistical Society. Series A (General), 129(1):118–119, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343903.

Kingman:1966:BRBa

[827] J. F. C. Kingman. Book review: Green’s Function Methods in Probability Theory, by Julian Keilson. Journal of the Royal Statistical Society. Series A (General), 129(1):119–120, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343904.

Whittle:1966:BRB

[828] P. Whittle. Book review: Introduction to Probability and Statistics from a Bayesian Viewpoint. Part 1: Probability, by D. V. Lindley; Introduction to Probability and Statistics from a Bayesian Viewpoint.Part 2: Inference, by D. V. Lindley. Journal of the Royal Statistical Society. Series A (General), 129(1):120–122, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343905.

Healy:1966:BRB

[829] M. J. R. Healy. Book review: L’Analyse hiérarchique, by B. Matalon. Journal of the Royal Statistical Society. Series A (General), 129(1):122, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343906.

Smith:1966:BRBa

[830] T. M. F. Smith. Book review: Descriptive Statistics: A Programmed Textbook, Vol. 2, by Lassar G. Gotkin; Leo S. Goldstein. Journal of the Royal Statistical Society. Series A (General), 129 (1):123, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343907.

Leser:1966:BRBa

[831] C. E. V. Leser. Book review: Essays on Econometrics and Planning, by C. R. Rao; D. B. Lahiri; K. R. Nair; P. Pant; S. S. Shrikhande. Journal of the Royal Statistical Society. Series A (General), 129(1):123–124, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343908.

Leser:1966:BRBb

[832] C. E. V. Leser. Book review: Statistical Inference and Measurement of Structural Changes in an Economy, by K. S. Rao. Journal of the Royal Statistical Society. Series A (General), 129 (1):124–125, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343909.

Mills:1966:BRBa

[833] G. Mills. Book review: Models for Decision, by C. M. Berners-Lèe. Journal of the Royal Statistical Society. Series A (General), 129(1):125–126, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343910.

Mills:1966:BRBb

[834] G. Mills. Book review: Exercises in Business Decisions, by J. N. Fairhead; D. S. Pugh; W. J. Williams; Business Exercises: Some Developments, by A. Life; D. Pugh. Journal of the Royal Statistical Society. Series A (General), 129(1):126, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343911.

Amey:1966:BRB

[835] L. R. Amey. Book review: Management Goals and Accounting for Control, by Y. Ijiri. Journal of the Royal Statistical Society. Series A (General), 129 (1):126–127, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343912.

Whitaker:1966:BRB

[836] J. K. Whitaker. Book review: A Programme for Growth (6). Exploring 1970: Some Numerical Results. Journal of the Royal Statistical Society. Series A (General), 129(1):127–128, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343913.

Saunders:1966:BRB

[837] C. T. Saunders. Book review: On the Accuracy of Economic Observations, by Oskar Morgenstern. Journal of the Royal Statistical Society. Series A (General), 129(1):128–130, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343914.

Pyatt:1966:BRB

[838] Graham Pyatt. Book review: Capital– Output–Employment Ratios in Industrial Programming, by Witold Lissowski; J. Syskind; B. Przestepska; P. F. Knightsfield. Journal of the Royal Statistical Society. Series A (General), 129(1):130–131, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343915.

Morgan:1966:BRB

[839] E. Victor Morgan. Book review: Measuring the Nation’s Wealth. Journal of the Royal Statistical Society. Series A (General), 129(1):131–132, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343916.

Brothwell:1966:BRB

[840] J. F. Brothwell. Book review: Redistribution of Income in the United Kingdom in 1959, 1957 and 1953, by J. L. Nicholson. Journal of the Royal Statistical Society. Series A (General), 129 (1):132–133, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343917.

Clark:1966:BRBa

[841] Colin Clark. Book review: Feeding India’s Growing Millions, by P. V.

Sukhatme. Journal of the Royal Statistical Society. Series A (General), 129 (1):133–134, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343918.

Scammell:1966:BRB

[842] W. M. Scammell. Book review: Foreign Dollar Loans in Europe, by Paul Einzig. Journal of the Royal Statistical Society. Series A (General), 129 (1):134–135, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343919.

Hillmann:1966:BRBa

[843] H. C. Hillmann. Book review: Union Budgets: A Factual Study of the Finances of the Government of India [1950–51 to 1964–65]. Journal of the Royal Statistical Society. Series A (General), 129(1):135–137, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343920.

Laming:1966:BRB

[844] D. R. J. Laming. Book review: Psychological Tests and Personnel Decisions, by Lee J. Cronbach; Goldine C. Gleser. Journal of the Royal Statistical Society. Series A (General), 129(1):137, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343921.

Posner:1966:BRB

[845] Rebecca Posner. Book review: Quantitative Linguistics, by G. Herdan. Journal of the Royal Statistical Society. Series A (General), 129(1):138–140, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343922.

Grebenik:1966:BRB

[846] E. Grebenik. Book review: Mortality in Sweden, by Tage Larsson. Journal of the Royal Statistical Society. Series A (General), 129(1):140–141, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343923.

Mussett:1966:BRB

[847] M. V. Mussett. Book review: Collaborative Tests: Report of a Symposium held at the School of Pharmacy, Brighton College of Technology, May 30th 1964. Journal of the Royal Statistical Society. Series A (General), 129 (1):141, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343924.

Smith:1966:BRBb

[848] Cedric A. B. Smith. Book review: Experiments in Plant Hybridisation, by Gregor Mendel; Ronald A. Fisher. Journal of the Royal Statistical Society. Series A (General), 129(1):141– 142, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343925.

Anonymous:1966:CNa

[849] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 129(1):143–147, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343926.

Boldrini:1966:CG

[850] Marcello Boldrini. Corrado Gini. Journal of the Royal Statistical Society. Series A (General), 129(1):148–150, ????

1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343927.

Anonymous:1966:LAa

[851] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 129(1):151–154, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343928.

Burman:1966:EMS

[852] J. P. Burman. Errata: Moving seasonal adjustment of economic time series. Journal of the Royal Statistical Society. Series A (General), 129(1):155, ???? 1966. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2343929. See [782].

Anonymous:1966:BMa

[853] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 129(1):i–iv, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343930.

Anonymous:1966:FMb

[854] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 129(2):??, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343614.

Winsten:1966:URU

[855] C. B. Winsten and F. Savigear. The use of rooms and the use of land for housing. Journal of the Royal Statistical Society. Series A (General), 129 (2):157–209, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343615.

Glass:1966:FPG

[856] D. V. Glass. Fertility and population growth. Journal of the Royal Statistical Society. Series A (General), 129 (2):210–215, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343616.

Benjamin:1966:MTW

[857] B. Benjamin. Mortality trends in the world. Journal of the Royal Statistical Society. Series A (General), 129(2): 216–221, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343617.

Sukhatme:1966:WFS

[858] P. V. Sukhatme. The world’s food supplies. Journal of the Royal Statistical Society. Series A (General), 129(2): 222–248, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343618.

Plackett:1966:CTS

[859] R. L. Plackett. Current trends in statistical inference. Journal of the Royal Statistical Society. Series A (General), 129(2):249–267, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343619.

Good:1966:PW

[860] I. J. Good. The probability of war. Journal of the Royal Statistical Society. Series A (General), 129(2):268– 269, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343620.

Lloyd:1966:NBF

[861] F. J. Lloyd. Note on the Beesley–Foster paper — Victoria Line: Social benefit and finances. Journal of the Royal Statistical Society. Series A (General), 129 (2):270–272, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343621. See

[693].

Beesley:1966:NBF

[862] M. E. Beesley and C. D. Foster. Note on the Beesley–Foster paper — Victoria Line: Social benefits and finances: A rejoinder. Journal of the Royal Statistical Society. Series A (General), 129 (2):273, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343622. See

[693].

Burman:1966:MSA

[863] J. P. Burman. Moving seasonal adjustment of economic time series: Additional note. Journal of the Royal Statistical Society. Series A (General), 129 (2):274, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343623.

Beckerman:1966:BRB

[864] W. Beckerman and A. A. Walters. Book review: The British Economy in 1975. Journal of the Royal Statistical Society. Series A (General), 129(2): 275–280, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343624.

Anonymous:1966:WP

[865] Anonymous. Wholesale prices in 1965. Journal of the Royal Statistical Society. Series A (General), 129(2):281– 286, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343625.

Cormack:1966:BRB

[866] R. M. Cormack. Book review: Concepts of Statistical Inference, by W. C. Guenther. Journal of the Royal Statistical Society. Series A (General), 129 (2):287–288, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343626.

Forte:1966:BRB

[867] Francesco Forte. Book review: Metodologica Statistica, by Giovanni Lasorsa; Statistica Economica. by Giovanni Lasorsa. Journal of the Royal Statistical Society. Series A (General), 129(2):288–289, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343627.

Good:1966:BRBa

[868] I. J. Good. Book review: Mathematical Theory of Probability and Statistics, by Richard von Mises; Hilda Geiringer. Journal of the Royal Statistical Society. Series A (General), 129(2):289– 291, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343628.

Good:1966:BRBb

[869] I. J. Good. Book review: Nombres et Systèmes aléatoires, by Octav Onicescu. Journal of the Royal Statistical Society. Series A (General), 129(2): 291, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343629.

Barton:1966:BRBa

[870] D. E. Barton. Book review: Contributions to Statistics, by C. R. Rao; D. B. Lahiri; K. R. Nair; P. Pant; S. S. Shrikhande. Journal of the Royal Statistical Society. Series A (General), 129 (2):291–293, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343630.

Tweedie:1966:BRB

[871] M. C. K. Tweedie. Book review: Lady Luck. The Theory of Probability, by Warren Weaver. Journal of the Royal Statistical Society. Series A (General), 129(2):293–295, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343631.

Kingman:1966:BRBb

[872] J. F. C. Kingman. Book review: Bibliography on Time Series and Stochastic Processes: An International Team Project, by Herman O. A. Wold. Journal of the Royal Statistical Society. Series A (General), 129(2):295, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343632.

Kingman:1966:BRBc

[873] J. F. C. Kingman. Book review: Mathematical Theory of Connecting Networks and Telephone Traffic, by V. E. Beneš. Journal of the Royal Statistical Society. Series A (General), 129(2): 295–296, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343633.

Wilkinson:1966:BRBa

[874] R. Wilkinson. Book review: Quality Control and Industrial Statistics, by Acheson J. Duncan. Journal of the Royal Statistical Society. Series A (General), 129(2):296–297, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343634.

Wilkinson:1966:BRBb

[875] R. Wilkinson. Book review: Essentials of Quality Control, by Alan Huitson; Joan Keen. Journal of the Royal Statistical Society. Series A (General), 129(2):297, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343635.

Barton:1966:BRBb

[876] D. E. Barton. Book review: Bibliography of Statistical Literature 1940–1949, by M. G. Kendall; A. G. Doig. Journal of the Royal Statistical Society. Series A (General), 129(2):297–298, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343636.

Lloyd:1966:BRBa

[877] E. H. Lloyd. Book review: Statistics for Scientists and Engineers, by R. Lowell Wine. Journal of the Royal Statistical Society. Series A (General), 129 (2):298–299, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343637.

Deane:1966:BRBa

[878] Phyllis M. Deane. Book review: Evidences of Long Swings in Aggregate Construction since the Civil War, by Moses Abramovitz. Journal of the Royal Statistical Society. Series A (General), 129(2):299, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343638.

Welham:1966:BRB

[879] P. J. Welham. Book review: Determinants and Effects of Changes in the Stock of Money, 1875–1960, by Phillip Cagan. Journal of the Royal Statistical Society. Series A (General), 129 (2):300, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343639.

Nevin:1966:BRB

[880] Edward Nevin. Book review: External Trade and the Economic Structure of Ceylon, 1900–1955, by Elaine Gunewardena. Journal of the Royal Statistical Society. Series A (General), 129(2):300–301, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343640.

Beesley:1966:BRB

[881] M. E. Beesley. Book review: Rail and Road in East Africa: Transport Co-Ordination in Undeveloped Countries, by Arthur Hazlewood. Journal of the Royal Statistical Society. Series A (General), 129(2):301–303, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343641.

Leser:1966:BRBc

[882] C. E. V. Leser. Book review: A Short-Term Model for Pakistan Economy: An Econometric Analysis, by Nurul Islam. Journal of the Royal Statistical Society. Series A (General), 129(2):303– 304, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343642.

Clark:1966:BRBb

[883] Colin Clark. Book review: The Impact of Price Movements on Areas under Selected Crops in India, 1900–39, by Dharm Narain. Journal of the Royal Statistical Society. Series A (General), 129(2):304, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343643.

Cooper:1966:BRB

[884] R. A. Cooper. Book review: Proceedings of the United Nations Conference on Trade and Development, Geneva, 23 March–16 June 1964. Vol. VI: Trade Expansion and Regional Groupings, Part I. Journal of the Royal Statistical Society. Series A (General), 129 (2):305, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343644.

Wilkinson:1966:BRBc

[885] R. Wilkinson. Book review: London’s Housing Needs, by Ruth Glass; John Westergaard. Journal of the Royal Statistical Society. Series A (General), 129 (2):306, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343645.

Kalton:1966:BRB

[886] G. Kalton. Book review: The Aged in the Welfare State, by Peter Townsend; Dorothy Wedderburn. Journal of the Royal Statistical Society. Series A (General), 129(2):306– 307, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343646.

Deane:1966:BRBb

[887] Phyllis M. Deane. Book review: Historical Statistics of Canada, by M. C. Urquhart; K. A. M. Buckley. Journal of the Royal Statistical Society. Series A (General), 129(2):307–308, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343647.

Brass:1966:BRB

[888] W. Brass. Book review: Trends in the Study of Morbidity and Mortality. Journal of the Royal Statistical Society. Series A (General), 129(2):309, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343648.

Anonymous:1966:CNb

[889] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 129(2):310–312, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343649.

Anonymous:1966:LAb

[890] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 129(2):313–315, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343650.

Anonymous:1966:BMb

[891] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 129(2):i–vii, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343651.

Anonymous:1966:FMc

[892] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 129(3):i–iii, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343501.

Chatfield:1966:PSM

[893] C. Chatfield, A. S. C. Ehrenberg, and G. J. Goodhardt. Progress on a simplified model of stationary purchasing behaviour. Journal of the Royal Statistical Society. Series A (General), 129 (3):317–367, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343502.

Zelen:1966:AEM

[894] M. Zelen. Application of exponential models to problems in cancer research. Journal of the Royal Statistical Society. Series A (General), 129(3):368– 398, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343503.

Moore:1966:SOR

[895] P. G. Moore. A survey of operational research. Journal of the Royal Statistical Society. Series A (General), 129 (3):399–447, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343504.

Peacock:1966:SAE

[896] Alan Peacock and Robert Lavers. The social accounting of education. Journal of the Royal Statistical Society. Series A (General), 129(3):448–466, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343505.

Lindley:1966:RUS

[897] D. V. Lindley. A report on university staff and students in statistics. Journal of the Royal Statistical Society. Series A (General), 129(3):467–470, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343506.

Zelen:1966:BRB

[898] Marvin Zelen. Book review: Mathematical Theory of Reliability, by Richard E. Barlow; Frank Proschan; Larry C. Hunter. Journal of the Royal Statistical Society. Series A (General), 129(3):471–472, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343507.

Kingman:1966:BRBd

[899] J. F. C. Kingman. Book review: Ergodic Theory and Information, by Patrick Billingsley. Journal of the Royal Statistical Society. Series A (General), 129(3):472, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343508.

Bartholomew:1966:BRBa

[900] D. J. Bartholomew. Book review: The Theory of Stochastic Processes, by D. R. Cox; H. D. Miller. Journal of the Royal Statistical Society. Series A (General), 129(3):473, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343509.

Brillinger:1966:BRBb

[901] D. R. Brillinger. Book review: Nonparametric Trend Analysis, by G. A.

Ferguson. Journal of the Royal Statistical Society. Series A (General), 129 (3):474, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343510.

Bather:1966:BRB

[902] J. A. Bather. Book review: Differential Games: A Mathematical Theory with Applications to Warfare and Pursuit, Control and Optimization, by Rufus Isaacs. Journal of the Royal Statistical Society. Series A (General), 129 (3):474–475, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343511.

Kingman:1966:BRBe

[903] J. F. C. Kingman. Book review: Mathematical Foundations of the Calculus of Probability, by Jacques Neveu. Journal of the Royal Statistical Society. Series A (General), 129(3):475–476, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343512.

Loynes:1966:BRB

[904] R. M. Loynes. Book review: Studies in the Theory of Random Processes, by A. V. Skorokhod. Journal of the Royal Statistical Society. Series A (General), 129(3):476, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343513.

Cox:1966:BRB

[905] D. R. Cox. Book review: Theory and Practice of the Evaluation of Measurements, by L. Jánossy. Journal of the Royal Statistical Society. Series A (General), 129(3):476–477, ????

1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343514.

Priestley:1966:BRB

[906] M. B. Priestley. Book review: Detection Theory, by Ivan Selin. Journal of the Royal Statistical Society. Series A (General), 129(3):477–478, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343515.

Beaumont:1966:BRB

[907] G. P. Beaumont. Book review: Principles of Statistics, by M. G. Bulmer. Journal of the Royal Statistical Society. Series A (General), 129(3):478– 479, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343516.

Niculescu:1966:BRB

[908] B. M. Niculescu. Book review: Calcul Economic, by Manea Mǎnescu. Journal of the Royal Statistical Society. Series A (General), 129(3):479–480, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343517.

Hillman:1966:BRB

[909] H. C. Hillman. Book review: Distribution of Family Incomes in Puerto Rico: A Case Study of the Impact of Economic Development on Income Distribution, by Faut M. Andic. Journal of the Royal Statistical Society. Series A (General), 129(3):480–481, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343518.

Anonymous:1966:CNc

[910] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 129(3):482–484, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343519.

Anonymous:1966:LAc

[911] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 129(3):485–488, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343520.

Anonymous:1966:BMc

[912] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 129(3):v–xi, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343521.

Anonymous:1966:VIb

[913] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 129(4):621– 624, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982252.

Anonymous:1966:FMd

[914] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 129(4):??, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982253.

Tippett:1966:SMI

[915] L. H. C. Tippett. Statistics in manufacturing industry. Journal of the Royal Statistical Society. Series A (General), 129(4):489–503, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982254.

Rutovitz:1966:PR

[916] Denis Rutovitz. Pattern recognition. Journal of the Royal Statistical Society. Series A (General), 129(4):504– 530, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982255.

Smith:1966:RRT

[917] T. M. F. Smith. Ratios of ratios and their applications. Journal of the Royal Statistical Society. Series A (General), 129(4):531–533, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982256.

Fleming:1966:LTM

[918] M. C. Fleming. The long-term measurement of construction costs in the United Kingdom. Journal of the Royal Statistical Society. Series A (General), 129(4):534–556, ???? 1966. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982257. See correction

[1024].

Anonymous:1966:ARC

[919] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 129 (4):557–575, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982258.

Anonymous:1966:POH

[920] Anonymous. Proceedings of the one hundred and thirty-second Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 129 (4):576, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982259.

Wilkinson:1966:BRBd

[921] R. K. Wilkinson. Book review: Report of the Committee on Housing in Greater London, by Committee on Housing in Greater London. Journal of the Royal Statistical Society. Series A (General), 129(4):577–586, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982260.

Bartholomew:1966:BRBb

[922] D. J. Bartholomew. Book review: The Statistical Analysis of Series of Events, by D. R. Cox; P. A. W. Lewis. Journal of the Royal Statistical Society. Series A (General), 129(4):587–588, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982261.

Constable:1966:BRB

[923] R. L. Constable. Book review: Basic Statistical Methods, by N. M. Downie; R. W. Heath. Journal of the Royal Statistical Society. Series A (General), 129 (4):588–589, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982262.

Kerridge:1966:BRB

[924] D. Kerridge. Book review: Logic of Statistical Inference, by Ian Hacking. Journal of the Royal Statistical Society. Series A (General), 129(4):589, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982263.

Hitchcock:1966:BRB

[925] Shirley Hitchcock. Book review: Introduction to Mathematical Statistics, by Robert V. Hogg; Allen T. Craig. Journal of the Royal Statistical Society. Series A (General), 129(4):589–590, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982264.

Wise:1966:BRB

[926] M. E. Wise. Book review: Mathematische Statistik, by B. L. van der Waerden. Journal of the Royal Statistical Society. Series A (General), 129 (4):590–591, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982265.

Oliver:1966:BRB

[927] F. R. Oliver. Book review: Basic Concepts of Measurement, by Brian Ellis. Journal of the Royal Statistical Society. Series A (General), 129(4):591, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982266.

Lloyd:1966:BRBb

[928] E. H. Lloyd. Book review: Long-Term Storage: An Experimental Study, by H. E. Hurst; R. P. Black; Y. M. Simaika. Journal of the Royal Statistical Society. Series A (General), 129(4):591– 593, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982267.

Gilchrist:1966:BRB

[929] W. G. Gilchrist. Book review: Introductory Engineering Statistics, by Irwin Guttman; S. S. Wilks. Journal of the Royal Statistical Society. Series A (General), 129(4):593–594, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982268.

Garside:1966:BRB

[930] M. J. Garside. Book review: Problems for Computer Solution, by F. Gruenberger; G. Jaffray. Journal of the Royal Statistical Society. Series A (General), 129(4):594–595, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982269.

Newell:1966:BRB

[931] D. J. Newell. Book review: Mathematics and Computer Science in Biology and Medicine, by Medical Research Council. Journal of the Royal Statistical Society. Series A (General), 129 (4):595–596, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982270.

Freeman:1966:BRB

[932] C. Freeman. Book review: Administering Research and Development: The Behaviour of Scientists and Engineers in Organizations, by Charles D. Orth. Journal of the Royal Statistical Society. Series A (General), 129(4):596– 597, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982271.

Vajda:1966:BRB

[933] S. Vajda. Book review: Operations Research: Process and Strategy, by David S. Stoller. Journal of the Royal Statistical Society. Series A (General), 129 (4):597–598, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982272.

Hillmann:1966:BRBb

[934] H. C. Hillmann. Book review: Das Wachstum der Deutschen Wirtschaft seit der Mitte des 19. Jahrhunderts, by Walther G. Hoffmann. Journal of the Royal Statistical Society. Series A (General), 129(4):598–599, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982273.

Shepherd:1966:BRB

[935] J. R. Shepherd. Book review: Costs, Prices and Profits: Their Cyclical Relations, by Thor Hultgren. Journal of the Royal Statistical Society. Series A (General), 129(4):599–600, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982274.

Ball:1966:BRB

[936] R. J. Ball. Book review: Random Processes and the Growth of Firms: A Study of the Pareto Law, by Josef Steindl. Journal of the Royal Statistical Society. Series A (General), 129 (4):600–601, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982275.

Maurice:1966:BRB

[937] Rita Maurice. Book review: Systems of Social Accounts, by G. Stuvel. Journal of the Royal Statistical Society. Series A (General), 129(4):601–602, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982276.

Thomas:1966:BRB

[938] Brinley Thomas. Book review: Building Cycles and Britain’s Growth, by J. Parry Lewis. Journal of the Royal Statistical Society. Series A (General), 129 (4):602–603, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982277.

Stone:1966:BRB

[939] P. A. Stone. Book review: The Construction Industry of Great Britain, by J. R. Colclough. Journal of the Royal Statistical Society. Series A (General), 129(4):603–604, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982278.

Stamp:1966:BRB

[940] L. Dudley Stamp. Book review: Population Growth and Planning Policy: An Analysis of Social and Economic Factors Affecting Housing and Employment Location in the West Midlands, by D. E. C. Eversley; Valerie J. Jackson; G. M. Lomas. Journal of the Royal Statistical Society. Series A (General), 129(4):604–605, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982279.

Wheatcroft:1966:BRB

[941] Stephen Wheatcroft. Book review: Transportation Economics: A Conference of the Universities — National Bureau Committee for Economic Research, by National Bureau Committee for Economic Research. Journal of the Royal Statistical Society. Series A (General), 129(4):605, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982280.

Berrington:1966:BRB

[942] Hugh Berrington. Book review: British Parliamentary Election Results 1950– 1964, by B. R. Mitchell; Klaus Boehm. Journal of the Royal Statistical Society. Series A (General), 129(4):606, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982281.

Prothero:1966:BRB

[943] R. Mansell Prothero. Book review: Malaria Eradication and Population Growth, with Special Reference to Ceylon and British Guiana, by Peter Newman. Journal of the Royal Statistical Society. Series A (General), 129(4): 606–607, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982282.

Anonymous:1966:CNd

[944] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 129(4):608–611, ???? 1966. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982283.

Irwin:1966:LIM

[945] J. O. Irwin. Leon Isserlis, M.A., D.Sc. (1881–1966). Journal of the Royal Statistical Society. Series A (General), 129 (4):612–616, ???? 1966. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982284.

Dobb:1966:OL

[946] M. H. Dobb. Oskar lange. Journal of the Royal Statistical Society. Series A (General), 129(4):616–617, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982285.

Anonymous:1966:LAd

[947] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 129(4):618–620, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982286.

Anonymous:1966:BMd

[948] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 129(4):i–vii, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982287.

Anonymous:1967:VIa

[949] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 130(1):??, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344035.

Anonymous:1967:FMa

[950] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 130(1):??, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344036.

Ehrenberg:1967:MTA

[951] A. S. C. Ehrenberg and W. A. Twyman. On measuring television audiences. Journal of the Royal Statistical Society. Series A (General), 130(1):1– 60, ???? 1967. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344037.

Keyfitz:1967:RPM

[952] Nathan Keyfitz. Reconciliation of population models: Matrix, integral equation and partial fraction. Journal of the Royal Statistical Society. Series A (General), 130(1):61–83, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344038.

Iyengar:1967:SEE

[953] N. Sreenivasa Iyengar. Some estimates of Engel elasticities based on national sample survey data. Journal of the Royal Statistical Society. Series A (General), 130(1):84–101, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344039.

Good:1967:GST

[954] I. J. Good and T. N. Gover. The generalized serial test and the binary expansion of \sqrt{2}. Journal of the Royal Statistical Society. Series A (General), 130 (1):102–107, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344040. See

corrigendum [1165].

Anonymous:1967:LR

[955] Anonymous. [list of reviews]. Journal of the Royal Statistical Society. Series A (General), 130(1):108, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344041.

Kingman:1967:BRB

[956] J. F. C. Kingman. Book review: An Introduction to Probability Theory and its Applications, by William Feller. Journal of the Royal Statistical Society.

Series A (General), 130(1):109, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344042.

Pearce:1967:BRBa

[957] S. C. Pearce. Book review: The Analysis of Variance, by Alan Huitson. Journal of the Royal Statistical Society. Series A (General), 130(1):109–110, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344043.

Daley:1967:BRB

[958] D. J. Daley. Book review: Probability Theory, by Klaus Krickeberg. Journal of the Royal Statistical Society. Series A (General), 130(1):110–111, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344044.

Miller:1967:BRBa

[959] H. D. Miller. Book review: Certain Studies on the Structure and Statistics of the Logarithmic Series Distribution and Related Tables, by G. P. Patil; A. R. Kamat; J. K. Wani. Journal of the Royal Statistical Society. Series A (General), 130(1):111, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344045.

Walker:1967:BRB

[960] A. M. Walker. Book review: Linear Statistical Inference and its Applications, by C. R. Rao. Journal of the Royal Statistical Society. Series A (General), 130(1):111–112, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344046.

Pearce:1967:BRBb

[961] S. C. Pearce. Book review: The Linear Hypothesis: A General Theory, by G. A. F. Seber. Journal of the Royal Statistical Society. Series A (General), 130(1):113, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344047.

Miller:1967:BRBb

[962] H. D. Miller. Book review: Research Papers in Statistics: Festschrift for J. Neyman, by F. N. David; Evelyn Fix. Journal of the Royal Statistical Society. Series A (General), 130(1):113– 114, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344048.

Bather:1967:BRBa

[963] J. A. Bather. Book review: A First Course in Stochastic Processes, by Samuel Karlin. Journal of the Royal Statistical Society. Series A (General), 130(1):114, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344049.

Page:1967:BRB

[964] E. S. Page. Book review: The Monte Carlo Method: The Method of Statistical Trials, by Yu. A. Shreider; G. J. Tee. Journal of the Royal Statistical Society. Series A (General), 130(1):114– 115, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344050.

Wilkinson:1967:BRBa

[965] R. Wilkinson. Book review: Statistics: An Introduction to Quantitative Economic Research, by Daniel B. Suits.

Journal of the Royal Statistical Society. Series A (General), 130(1):115– 117, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344051.

Brass:1967:BRBa

[966] W. Brass. Book review: Statistical Methods in Malaria Eradication, by Satya Swaroop; A. B. Gilroy; K. Uemura. Journal of the Royal Statistical Society. Series A (General), 130(1): 117, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344052.

Healy:1967:BRBa

[967] M. J. R. Healy. Book review: Théorie et Pratique des Sondages, by J. Desabie. Journal of the Royal Statistical Society. Series A (General), 130(1):118, ???? 1967. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344053.

Ahamad:1967:BRB

[968] B. Ahamad. Book review: Sampling Methods and Censuses, by S. S. Zarkovich. Journal of the Royal Statistical Society. Series A (General), 130 (1):118, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344054.

Smith:1967:BRBa

[969] T. M. F. Smith. Book review: Elementary Statistics, by Paul G. Hoel. Journal of the Royal Statistical Society. Series A (General), 130(1):118–119, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344055.

Ilersic:1967:BRB

[970] A. R. Ilersic. Book review: A First Course in Statistics, by Robert Loveday; Statistics for Business. by Leonard J. Holman. Journal of the Royal Statistical Society. Series A (General), 130 (1):119–121, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344056.

Healy:1967:BRBb

[971] M. J. R. Healy. Book review: Matrix Algebra for the Biological Sciences (Including Applications in Statistics), by S. R. Searle. Journal of the Royal Statistical Society. Series A (General), 130 (1):121, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344057.

Barton:1967:BRB

[972] D. E. Barton. Book review: Programming, Games and Transportation Networks, by Claude Berge; A. Ghouila-Houri; M. Merrington; C. Ramanujacharyulu. Journal of the Royal Statistical Society. Series A (General), 130 (1):121–122, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344058.

Simpson:1967:BRB

[973] H. R. Simpson. Book review: OM-NITAB: A Computer Program for Statistical and Numerical Analysis, by Joseph Hilsenrath; Guy G. Ziegler; Carla G. Messina; Philip J. Walsh; Robert J. Herbold. Journal of the Royal Statistical Society. Series A (General), 130(1):122–123, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344059.

Chambers:1967:BRBa

[974] John M. Chambers. Book review: Random Number Generators, by Birger Jansson. Journal of the Royal Statistical Society. Series A (General), 130 (1):123–124, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344060.

Hill:1967:BRB

[975] I. D. Hill. Book review: Computers in the Hospital Service. Journal of the Royal Statistical Society. Series A (General), 130(1):124, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344061.

Mishan:1967:BRB

[976] E. J. Mishan. Book review: Analytical Economics: Issues and Problems, by Nicholas Georgescu-Roegen. Journal of the Royal Statistical Society. Series A (General), 130(1):125–126, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344062.

Welham:1967:BRB

[977] P. J. Welham. Book review: Studies in Short-Term National Accounts and Long-Term Economic Growth, by Simon Goldberg; Phyllis Deane. Journal of the Royal Statistical Society. Series A (General), 130(1):126–127, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344063.

Savage:1967:BRB

[978] C. I. Savage. Book review: Economic Enquiry in Australia, by Crauford D.

W. Goodwin. Journal of the Royal Statistical Society. Series A (General), 130 (1):127–128, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344064.

Lury:1967:BRB

[979] D. A. Lury. Book review: An Economic Survey of the Sisal Industry of Tanganyika, by C. W. Guillebaud. Journal of the Royal Statistical Society. Series A (General), 130(1):128–129, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344065.

Coppock:1967:BRB

[980] D. J. Coppock. Book review: Economic Growth in Britain, by P. D. Henderson. Journal of the Royal Statistical Society. Series A (General), 130 (1):129–131, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344066.

Leser:1967:BRBa

[981] C. E. V. Leser. Book review: Consumer Demand in the United States 1929– 1970: Analyses and Projections, by H. S. Houthakker; Lester D. Taylor. Journal of the Royal Statistical Society. Series A (General), 130(1):131–132, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344067.

Fisk:1967:BRB

[982] P. R. Fisk. Book review: An Econometric Study of International Trade Flows, by Hans Linnemann. Journal of the Royal Statistical Society. Series A (General), 130(1):132, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344068.

Oliver:1967:BRBa

[983] F. R. Oliver. Book review: Consumer Credit Finance Charges, by Wallace P. Mors. Journal of the Royal Statistical Society. Series A (General), 130(1): 132–133, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344069.

Thatcher:1967:BRB

[984] A. R. Thatcher. Book review: Output, Employment and Productivity in the United States after 1800, by National Bureau of Economic Research. Journal of the Royal Statistical Society. Series A (General), 130(1):133–134, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344070.

Lewis:1967:BRB

[985] J. Parry Lewis. Book review: A Scheme of Economic Theory, by G. L. S. Shackle. Journal of the Royal Statistical Society. Series A (General), 130 (1):134–135, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344071.

Brown:1967:BRB

[986] A. J. Brown. Book review: Papers on Regional Development, by Thomas Wilson. Journal of the Royal Statistical Society. Series A (General), 130 (1):135–136, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344072.

Nevin:1967:BRB

[987] Edward Nevin. Book review: Monetary Equilibrium and Economic Development with Special Reference to the Experience of Greece 1950–1963, by Xenophon Zolotas. Journal of the Royal Statistical Society. Series A (General), 130(1):136–137, ???? 1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344073.

Sorrie:1967:BRB

[988] G. S. Sorrie. Book review: Social and Economic Factors Affecting Mortality, by B. Benjamin. Journal of the Royal Statistical Society. Series A (General), 130(1):137, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344074.

Anonymous:1967:CNa

[989] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 130(1):138–143, ???? 1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344075.

Anonymous:1967:LAa

[990] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 130(1):144–146, ???? 1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344076.

Anonymous:1967:BMa

[991] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 130(1):i–vii, ???? 1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344077.

Anonymous:1967:FMb

[992] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 130(2):??, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343398.

Irwin:1967:WAW

[993] J. O. Irwin. William Allen Whitworth and a hundred years of probability. Journal of the Royal Statistical Society. Series A (General), 130(2):147– 176, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343399.

Barnett:1967:SLT

[994] V. D. Barnett and T. Lewis. A study of low-temperature probabilities in the context of an industrial problem. Journal of the Royal Statistical Society. Series A (General), 130(2):177–206, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343400.

Bates:1967:DCS

[995] James Bates and S. J. Henderson. The determinants of corporate saving in small private companies in Britain, 1954–56. Journal of the Royal Statistical Society. Series A (General), 130 (2):207–224, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343401.

Selmer:1967:RNN

[996] Ernst S. Selmer. Registration numbers in Norway: Some applied number theory and psychology. Journal of the Royal Statistical Society. Series A (General), 130(2):225–231, ????

1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343402.

Ord:1967:GMC

[997] J. K. Ord. Graphical methods for a class of discrete distributions. Journal of the Royal Statistical Society. Series A (General), 130(2):232–238, ???? 1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343403.

Manly:1967:AMC

[998] B. F. J. Manly. An approximate method for calculating the average sample number in certain sequential tests. Journal of the Royal Statistical Society. Series A (General), 130(2): 239–243, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343404.

Anonymous:1967:WP

[999] Anonymous. Wholesale prices in 1966. Journal of the Royal Statistical Society. Series A (General), 130(2):244– 249, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343405.

Healy:1967:BRBc

[1000] M. J. R. Healy. Book review: The Design of Experiments, by Ronald A. Fisher. Journal of the Royal Statistical Society. Series A (General), 130 (2):250, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343406.

Kerridge:1967:BRB

[1001] D. F. Kerridge. Book review: The Logic of Decision, by R. C. Jeffrey. Journal of the Royal Statistical Society. Series A (General), 130(2):250–251, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343407.

Doney:1967:BRB

[1002] R. A. Doney. Book review: Introduction to Measure and Probability, by J. F. C. Kingman; S. J. Taylor. Journal of the Royal Statistical Society. Series A (General), 130(2):251, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343408.

Lee:1967:BRB

[1003] P. M. Lee. Book review: Information Theory, by Robert B. Ash. Journal of the Royal Statistical Society. Series A (General), 130(2):252, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343409.

Kemp:1967:BRB

[1004] K. W. Kemp. Book review: Sequential Methods in Statistics, by G. Barrie Wetherill. Journal of the Royal Statistical Society. Series A (General), 130 (2):252–253, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343410.

Constable:1967:BRB

[1005] R. L. Constable. Book review: Biostatistics, by Alvin E. Lewis. Journal of the Royal Statistical Society. Series A (General), 130(2):253, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343411.

Chatfield:1967:BRB

[1006] C. Chatfield. Book review: Mathematics and Statistics for Scientists and Engineers, by P. MacDonald. Journal of the Royal Statistical Society. Series A (General), 130(2):253–254, ???? 1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343412.

Downton:1967:BRB

[1007] F. Downton. Book review: Queues and Inventories: A Study of their Basic Stochastic Processes, by N. U. Prabhu. Journal of the Royal Statistical Society. Series A (General), 130(2):254– 255, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343413.

Snell:1967:BRB

[1008] E. J. Snell. Book review: Handbook of Tables for Probability and Statistics, by William H. Beyer. Journal of the Royal Statistical Society. Series A (General), 130(2):255–256, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343414.

Rahman:1967:BRB

[1009] N. A. Rahman. Book review: Symmetric Function and Allied Tables, by F. N. David; M. G. Kendall; D. E. Barton. Journal of the Royal Statistical Society. Series A (General), 130 (2):256–257, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343415.

Welch:1967:BRB

[1010] B. L. Welch. Book review: Biometrika Tables for Statisticians, by E. S. Pearson; H. O. Hartley. Journal of the Royal Statistical Society. Series A (General), 130(2):257–258, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343416.

Lawson:1967:BRB

[1011] G. H. Lawson. Book review: Growth Rate Tables, by H. T. Burley. Journal of the Royal Statistical Society. Series A (General), 130(2):258, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343417.

Brillinger:1967:BRB

[1012] D. R. Brillinger. Book review: A Guide to FORTRAN IV Programming, by Daniel D. McCracken. Journal of the Royal Statistical Society. Series A (General), 130(2):258–259, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343418.

Oliver:1967:BRBb

[1013] F. R. Oliver. Book review: Introduction to Statistical Ideas for Social Scientists, by G. Kalton. Journal of the Royal Statistical Society. Series A (General), 130 (2):259–260, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343419.

Cooper:1967:BRB

[1014] R. A. Cooper. Book review: The Use of Economic Statistics, by Conrad Blyth. Journal of the Royal Statistical Society. Series A (General), 130(2):260– 261, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343420.

Osborn:1967:BRB

[1015] J. F. Osborn. Book review: Replication: An Approach to the Analysis of Data from Complex Surveys; Development and Evaluation of a Replication Technique for Estimating Variance. Journal of the Royal Statistical Society. Series A (General), 130(2): 261–262, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343421.

Mills:1967:BRB

[1016] G. Mills. Book review: Parking Space for Cars: Assessing the Demand, by G. J. Roth. Journal of the Royal Statistical Society. Series A (General), 130 (2):262, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343422.

Hillmann:1967:BRB

[1017] H. C. Hillmann. Book review: Foreign Tax Policies and Economic Growth: A Conference Report of the National Bureau of Economic Research and the Brookings Institution, by National Bureau of Economic Research; Brookings Institution. Journal of the Royal Statistical Society. Series A (General), 130 (2):263, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343423.

Hughes:1967:BRB

[1018] John Hughes. Book review: Productivity in Distribution, by K. D. George. Journal of the Royal Statistical Society. Series A (General), 130(2):264– 265, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343424.

Grebenik:1967:BRB

[1019] E. Grebenik. Book review: The Travel Diaries of Thomas Robert Malthus, by Patricia James. Journal of the Royal Statistical Society. Series A (General), 130(2):265–266, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343425.

Anonymous:1967:CNb

[1020] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 130(2):267–273, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343426.

Kinross:1967:RHL

[1021] John B. Kinross and Arnold Plant. Rt.

Hon. The Lord Piercy, C.B.E. Journal of the Royal Statistical Society. Series A (General), 130(2):274–276, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343427.

Jones:1967:SAC

[1022] D. Caradog Jones and T. H. Marshall.

Sir Alexander Carr-Saunders, K.B.E., F.B.A. Journal of the Royal Statistical Society. Series A (General), 130 (2):276–279, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343428.

Cochran:1967:LJR

[1023] W. G. Cochran. Lowell Jacob Reed (1886–1966). Journal of the Royal Statistical Society. Series A (General), 130 (2):279–281, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343429.

Fleming:1967:CLT

[1024] M. C. Fleming. Correction: The long-term measurement of construction costs in the United Kingdom. Journal of the Royal Statistical Society. Series A (General), 130(2):282, ???? 1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343430. See [918].

Anonymous:1967:LAb

[1025] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 130(2):283–284, ???? 1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343431.

Anonymous:1967:BMb

[1026] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 130(2):i–vii, ???? 1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343432.

Anonymous:1967:FMc

[1027] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 130(3):??, ???? 1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344264.

Neyman:1967:EWC

[1028] Jerzy Neyman. Experimentation with weather control. Journal of the Royal Statistical Society. Series A (General), 130(3):285–326, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344265.

Thomson:1967:ETP

[1029] J. M. Thomson. An evaluation of two proposals for traffic restraint in central London. Journal of the Royal Statistical Society. Series A (General), 130 (3):327–377, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344266.

Masson:1967:ECR

[1030] Francis Masson and James Theberge.

External capital requirements and economic development: The case of Argentina. Journal of the Royal Statistical Society. Series A (General), 130 (3):378–407, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344267.

Knight:1967:COS

[1031] Rose Knight. Changes in the occupational structure of the working population. Journal of the Royal Statistical Society. Series A (General), 130(3): 408–422, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344268.

Redish:1967:CJB

[1032] K. A. Redish and Alan S. C. Ross. The chance of the jackpot being opened at poker. Journal of the Royal Statistical Society. Series A (General), 130(3): 423–426, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344269.

Morgan:1967:BRB

[1033] Roger W. Morgan. Book review: Elements of Probability Theory, by J. Bass. Journal of the Royal Statistical Society. Series A (General), 130(3):427– 428, ???? 1967. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2344270.

David:1967:BRB

[1034] F. N. David. Book review: Elementary Probability, by Edward O. Thorp. Journal of the Royal Statistical Society. Series A (General), 130(3):428, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344271.

Ruben:1967:BRB

[1035] H. Ruben. Book review: An Introduction to Stochastic Processes: With Special Reference to Methods and Applications, by M. S. Bartlett. Journal of the Royal Statistical Society. Series A (General), 130(3):429–430, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344272.

Barnett:1967:BRB

[1036] V. D. Barnett. Book review: Classical and Contagious Discrete Distributions, by Ganapati P. Patil. Journal of the Royal Statistical Society. Series A (General), 130(3):430–431, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344273.

Silvey:1967:BRB

[1037] S. D. Silvey. Book review: Hypothesis-Testing Problems with the Alternative Restricted by a Number of Inequalities, by W. Schaafsma. Journal of the Royal Statistical Society. Series A (General), 130(3):431–432, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344274.

Beaumont:1967:BRB

[1038] G. P. Beaumont. Book review: Elements of Mathematical Statistics, by J. F. Ractliffe. Journal of the Royal Statistical Society. Series A (General), 130(3):432, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344275.

vanZwet:1967:BRB

[1039] W. R. van Zwet. Book review:

Einführung in die Mathematische Statistik, by Leopold Schmetterer. Journal of the Royal Statistical Society. Series A (General), 130(3):432– 433, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344276.

Conway:1967:BRB

[1040] Freda Conway. Book review: Basic Statistics for Students of Economics, Commerce, Accountancy and the Biological Sciences, by Atindra M. Goon; Milan K. Gupta; Bhagabat Dasgupta. Journal of the Royal Statistical Society. Series A (General), 130(3):433– 434, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344277.

Heyde:1967:BRB

[1041] C. C. Heyde. Book review: Proceedings of the Symposium on Congestion Theory, by Walter L. Smith; William E. Wilkinson. Journal of the Royal Statistical Society. Series A (General), 130 (3):434–435, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344278.

Patterson:1967:BRB

[1042] H. D. Patterson. Book review: Variance Analysis of Complete Designs: Some Practical Aspects, by Paul Seeger. Journal of the Royal Statistical Society. Series A (General), 130 (3):435, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344279.

Ray:1967:BRB

[1043] W. D. Ray. Book review: Applied Methods of the Theory of Random Functions, by A. A. Sveshnikov; J. Berry; L. Haller. Journal of the Royal Statistical Society. Series A (General), 130(3):435–436, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344280.

Rowe:1967:BRB

[1044] B. C. Rowe. Book review: Data Processing: How to Program and Operate Punching, Sorting, Accounting, and Electronic Statistical Machines, by Harry P. Hartkemeier. Journal of the Royal Statistical Society. Series A (General), 130(3):436–437, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344281.

Wallis:1967:BRB

[1045] Kenneth F. Wallis. Book review: Statistical Methods of Econometrics, by E. Malinvaud; A. Silvey. Journal of the Royal Statistical Society. Series A (General), 130(3):437–438, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344282.

Ball:1967:BRB

[1046] R. J. Ball. Book review: Econometric Techniques and Problems, by C. E. V. Leser. Journal of the Royal Statistical Society. Series A (General), 130 (3):438–439, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344283.

Brothwell:1967:BRB

[1047] J. F. Brothwell. Book review: Income Distribution in the United States, by Herman P. Miller. Journal of the Royal Statistical Society. Series A (General), 130(3):439–440, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344284.

Alexander:1967:BRB

[1048] K. J. W. Alexander. Book review:

Wage Trends, Wage Policies, and Collective Bargaining: The Problems for Underdeveloped Countries, by H. A. Turner. Journal of the Royal Statistical Society. Series A (General), 130 (3):440–441, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344285.

Leser:1967:BRBb

[1049] C. E. V. Leser. Book review: The Measurement of Consumers’ Expenditure and Behaviour in the United Kingdom 1920–1938, by Richard Stone; D. A. Rowe. Journal of the Royal Statistical Society. Series A (General), 130 (3):441–442, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344286.

Bellamy:1967:BRB

[1050] R. Bellamy. Book review: The Owners of Quoted Ordinary Shares: A Survey for 1963, by Jack Revell; John Moyle. Journal of the Royal Statistical Society. Series A (General), 130 (3):442–443, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344287.

Fleming:1967:BRB

[1051] M. C. Fleming. Book review: Source Book of Statistics Relating to Construction, by Robert E. Lipsey; Doris Preston. Journal of the Royal Statistical Society. Series A (General), 130(3): 443–444, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344288.

Stone:1967:BRB

[1052] P. A. Stone. Book review: Interdependence and Uncertainty: A Study of the Building Industry, by Tavistock Institute. Journal of the Royal Statistical Society. Series A (General), 130 (3):444–445, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344289.

Healy:1967:BRBd

[1053] M. J. R. Healy. Book review: Problems and Progress in Medical Care: Essays on Current Research (Second Series), by Gordon McLachlan. Journal of the Royal Statistical Society. Series A (General), 130(3):445, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344290.

Kelsall:1967:BRB

[1054] R. K. Kelsall. Book review: Graduate School: A Study of Graduate Work at the London School of Economics, by Howard Glennerster; Arthur Bennett; Christine Farrell. Journal of the Royal Statistical Society. Series A (General), 130(3):445–446, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344291.

Anonymous:1967:CNc

[1055] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 130(3):447–453, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344292.

Anonymous:1967:LAc

[1056] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 130(3):454–456, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344293.

Anonymous:1967:BMc

[1057] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 130(3):i–iv, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344294.

Anonymous:1967:VIb

[1058] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series A (General), 130(4):601– 604, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982517.

Anonymous:1967:FMd

[1059] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 130(4):??, ????

1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982518.

Bartlett:1967:ISP

[1060] M. S. Bartlett. Inference and stochastic processes. Journal of the Royal Statistical Society. Series A (General), 130 (4):457–478, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982519.

Douglas:1967:AAS

[1061] J. W. B. Douglas, J. M. Ross, and H. R. Simpson. The ability and attainment of short-sighted pupils. Journal of the Royal Statistical Society. Series A (General), 130(4):479–504, ???? 1967. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982520. See corrigenda

[1131].

Dietz:1967:ERS

[1062] Klaus Dietz. Epidemics and rumours: A survey. Journal of the Royal Statistical Society. Series A (General), 130 (4):505–528, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982521.

Nicholson:1967:RC

[1063] M. B. Nicholson. The resolution of conflict. Journal of the Royal Statistical Society. Series A (General), 130 (4):529–540, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982522.

Goodman:1967:RMT

[1064] Leo A. Goodman. On the reconciliation of mathematical theories of population growth. Journal of the Royal Statistical Society. Series A (General), 130(4):541–553, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982523.

Anonymous:1967:ARC

[1065] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 130 (4):554–573, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982524.

Anonymous:1967:POH

[1066] Anonymous. Proceedings of the one hundred and thirty-third Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 130 (4):574, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982525.

Farlie:1967:BRB

[1067] D. J. G. Farlie. Book review: Exercises in Probability and Statistics for Mathematics Undergraduates: With Answers and Hints on Solutions, by N. A. Rahman. Journal of the Royal Statistical Society. Series A (General), 130(4): 575–576, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982526.

Chambers:1967:BRBb

[1068] John M. Chambers. Book review: Multivariate Statistical Methods, by Donald F. Morrison. Journal of the Royal Statistical Society. Series A (General), 130(4):576, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982527.

Boyd:1967:BRB

[1069] D. A. Boyd. Book review: Statistics of Crop Responses to Fertilizers. Journal of the Royal Statistical Society. Series A (General), 130(4):577, ???? 1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982528.

Smith:1967:BRBb

[1070] Cedric A. B. Smith. Book review: The Mathematical Approach to Biology and Medicine, by Norman T. J. Bailey. Journal of the Royal Statistical Society. Series A (General), 130(4):577– 578, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982529.

Siskind:1967:BRB

[1071] V. Siskind. Book review: Stochastic Models for Bacteriophage, by J. Gani. Journal of the Royal Statistical Society. Series A (General), 130(4):578– 579, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982530.

Gales:1967:BRB

[1072] K. E. Gales. Book review: Sampling: An Introduction for Social Scientists, by Freda Conway. Journal of the Royal Statistical Society. Series A (General), 130(4):579, ???? 1967. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982531.

Bather:1967:BRBb

[1073] J. A. Bather. Book review: An Introduction to Dynamic Programming: The Theory of Multistage Decision Processes, by O. L. R. Jacobs. Journal of the Royal Statistical Society. Series A (General), 130(4):579–580, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982532.

Hiorns:1967:BRB

[1074] R. W. Hiorns. Book review: Statistical Association Methods for Mechanized Documentation, by Mary E. Stevens; Vincent E. Giuliano; Laurence B. Heilperin. Journal of the Royal Statistical Society. Series A (General), 130 (4):580, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982533.

Wilkinson:1967:BRBb

[1075] R. Wilkinson. Book review: The Quality of Statistical Data, by S. S. Zarkovich. Journal of the Royal Statistical Society. Series A (General), 130 (4):581, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982534.

Oliver:1967:BRBc

[1076] F. R. Oliver. Book review: A Statistical Model for the Greek Economy 1949–1959, by P. Pavlopoulos. Journal of the Royal Statistical Society. Series A (General), 130(4):581–582, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982535.

Nove:1967:BRB

[1077] A. Nove. Book review: Basic Principles and Experience of Industrial Development Planning in the Soviet Union. Journal of the Royal Statistical Society. Series A (General), 130(4):582– 583, ???? 1967. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982536.

McCarthy:1967:BRB

[1078] M. J. McCarthy. Book review: Measuring Transactions Between World Areas, by Herbert B. Woolley. Journal of the Royal Statistical Society. Series A (General), 130(4):583–584, ???? 1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982537.

Bates:1967:BRB

[1079] R. W. Bates. Book review: The Demand for Natural Gas in the United States: A Dynamic Approach for the Residential and Commercial Market, by Pietro Balestra. Journal of the Royal Statistical Society. Series A (General), 130(4):584, ???? 1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982538.

Leser:1967:BRBc

[1080] C. E. V. Leser. Book review: Die Messung des Technischen Fortschritts im Rahmen des Gesamtwirtschaftlichen Wachstumsprozesses, by Florian H. Fleck. Journal of the Royal Statistical Society. Series A (General), 130(4): 585, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982539.

Ord:1967:BRB

[1081] Keith Ord. Book review: Household Capital Formation and Financing 1897–1962, by F. Thomas Juster. Journal of the Royal Statistical Society. Series A (General), 130(4):585–586, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982540.

Nicholson:1967:BRB

[1082] R. J. Nicholson. Book review: Reports of the Commonwealth Economic Committee: Forty-First Report–Fish, by Commonwealth Economic Committee. Journal of the Royal Statistical Society. Series A (General), 130(4):586– 587, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982541.

Brass:1967:BRBb

[1083] W. Brass. Book review: Public Health and Population Change, by Mindel C. Sheps; Jeanne Clare Ridley. Journal of the Royal Statistical Society. Series A (General), 130(4):587–589, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982542.

Kalmus:1967:BRB

[1084] H. Kalmus. Book review: Inheritance of Drinking Behavior: A Study on Intelligence, Personality, and Use of Alcohol of Adult Twins, by Juha Partanen; Kettil Bruun; Touko Markkanen. Journal of the Royal Statistical Society. Series A (General), 130(4):589, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982543.

Anonymous:1967:CNd

[1085] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 130(4):590–592, ???? 1967. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982544.

Tippett:1967:WS

[1086] L. H. C. Tippett. Walter A. Shewhart. Journal of the Royal Statistical Society. Series A (General), 130(4):593– 594, ???? 1967. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2982545.

Heyworth:1967:SHB

[1087] Heyworth. Sir Hugh Beaver, K.B.E. Journal of the Royal Statistical Society. Series A (General), 130(4):594, ???? 1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982546.

Anonymous:1967:LAd

[1088] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 130(4):595–600, ???? 1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982547.

Anonymous:1967:BMd

[1089] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 130(4):i–v, ???? 1967. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982548.

Anonymous:1968:VIa

[1090] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 131(1):??, ???? 1968. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344078.

Anonymous:1968:FMa

[1091] Anonymous. Front matter. Journal of the Royal Statistical Society.

Series A (General), 131(1):??, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344079.

Anonymous:1968:I

[1092] Anonymous. [introduction]. Journal of the Royal Statistical Society. Series A (General), 131(1):1–8, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344080.

Acheson:1968:SRC

[1093] E. D. Acheson. Some remarks on contemporary British medical statistics. Journal of the Royal Statistical Society. Series A (General), 131(1):9–12, ???? 1968. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344081.

Davies:1968:FLG

[1094] Bleddyn Davies. The future of local government statistics. Journal of the Royal Statistical Society. Series A (General), 131(1):13–19, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344082.

Fry:1968:SCB

[1095] R. E. Fry. Statistical collaboration between the government and the Greater London Council. Journal of the Royal Statistical Society. Series A (General), 131(1):20–34, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344083.

Pyatt:1968:OES

[1096] Graham Pyatt. On official economic statistics. Journal of the Royal Statistical Society. Series A (General), 131 (1):35–44, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344084.

Armstrong:1968:SMT

[1097] A. G. Armstrong and J. R. C. Lecomber. Statistics for mediumterm economic planning. Journal of the Royal Statistical Society. Series A (General), 131(1):45–60, ???? 1968. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344085.

Zinkin:1968:SBP

[1098] M. Zinkin and T. G. Booer. Statistics for business planning. Journal of the Royal Statistical Society. Series A (General), 131(1):61–74, ???? 1968. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344086.

Rainnie:1968:BRB

[1099] G. F. Rainnie. Book review: Traffic in Towns, by Steering Group; Working Group. Journal of the Royal Statistical Society. Series A (General), 131(1):75– 82, ???? 1968. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2344087.

Granger:1968:SAS

[1100] C. W. J. Granger and A. O. Hughes. Spectral analysis of short series — a simulation study. Journal of the Royal Statistical Society. Series A (General), 131(1):83–99, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344088.

Welch:1968:BRB

[1101] B. L. Welch. Book review: The Selected Papers of E. S. Pearson, by Biometrika Trustees. Journal of the Royal Statistical Society. Series A (General), 131 (1):100–101, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344089.

Snell:1968:BRB

[1102] E. J. Snell. Book review: Statistical Methods and Formulae, by C. G. Lambe. Journal of the Royal Statistical Society. Series A (General), 131 (1):101–102, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344090.

Smith:1968:BRB

[1103] T. M. F. Smith. Book review: Statistical Methods in Economics, by Harold W. Guthrie. Journal of the Royal Statistical Society. Series A (General), 131 (1):102–103, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344091.

Nisbet:1968:BRB

[1104] J. Nisbet. Book review: Statistics for the Teacher, by Douglas M. McIntosh. Journal of the Royal Statistical Society. Series A (General), 131(1):103, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344092.

Rahman:1968:BRB

[1105] N. A. Rahman. Book review: On Some Problems Associated with D^{2}-Statistics and p-Statistics, by P. K. Bose; S. B. Chaudhuri. Journal of the Royal Statistical Society. Series A (General), 131 (1):103–104, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344093.

Barton:1968:BRB

[1106] D. E. Barton. Book review: Structural Models: An Introduction in the Theory of Directed Graphs, by Frank Harary; Robert Z. Norman; Dorwin Cartwright. Journal of the Royal Statistical Society. Series A (General), 131 (1):104–105, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344094.

Lee:1968:BRB

[1107] P. M. Lee. Book review: Recent Results in Information Theory, by S. Kotz. Journal of the Royal Statistical Society. Series A (General), 131(1):105– 106, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344095.

Brillinger:1968:BRB

[1108] David R. Brillinger. Book review: Stationary and Related Stochastic Processes: Sample Function Properties and Their Applications, by Harald Cramér; M. R. Leadbetter. Journal of the Royal Statistical Society. Series A (General), 131(1):106–107, ???? 1968. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344096.

Alderson:1968:BRB

[1109] M. R. Alderson. Book review: Medical Record Linkage, by E. D. Acheson. Journal of the Royal Statistical Society. Series A (General), 131(1):107– 108, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344097.

Vaizey:1968:BRB

[1110] J. Vaizey. Book review: Census 1961, England and Wales. Journal of the Royal Statistical Society. Series A (General), 131(1):108–109, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344098.

Mills:1968:BRB

[1111] G. Mills. Book review: Operational Research and the Social Sciences, by J. R. Lawrence. Journal of the Royal Statistical Society. Series A (General), 131 (1):109–110, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344099.

Fessey:1968:BRB

[1112] M. C. Fessey. Book review: Statistics of the British Economy, by F. M. M. Lewes. Journal of the Royal Statistical Society. Series A (General), 131 (1):110–111, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344100.

Holmans:1968:BRB

[1113] A. E. Holmans. Book review: Economic Activity, by G. C. Harcourt; P. H. Karmel; R. H. Wallace. Journal of the Royal Statistical Society. Series A (General), 131(1):111–112, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344101.

Champernowne:1968:BRB

[1114] D. G. Champernowne. Book review:

Studies in the Theory of Business Cycles 1933–1939, by Michal Kalecki. Journal of the Royal Statistical Society. Series A (General), 131(1):112– 113, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344102.

Cooper:1968:BRB

[1115] R. A. Cooper. Book review: The Financing of Economic Development, by U.N. Department of Economic and Social Affairs. Journal of the Royal Statistical Society. Series A (General), 131 (1):113–114, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344103.

Worswick:1968:BRB

[1116] G. D. N. Worswick. Book review: National Economic Planning: A Conference of the Universities-National Bureau Committee for Economic Research, by Max F. Millikan. Journal of the Royal Statistical Society. Series A (General), 131(1):114, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344104.

McCrone:1968:BRB

[1117] Gavin McCrone. Book review: Methods of Selecting Industries for Depressed Areas: An Introduction to Feasibility Studies, by Leo H. Klaassen. Journal of the Royal Statistical Society. Series A (General), 131(1):115, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344105.

Livingstone:1968:BRB

[1118] Ian Livingstone. Book review: Economics of Development in Village India, by M. R. Haswell. Journal of the Royal Statistical Society. Series A (General), 131(1):115–116, ????

1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344106.

Hillmann:1968:BRB

[1119] H. C. Hillmann. Book review: Some Aspects of Economic Planning in Western European Industrial Countries, by Gisele Podbielski. Journal of the Royal Statistical Society. Series A (General), 131(1):117, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344107.

Wilkinson:1968:BRBa

[1120] R. Wilkinson. Book review: Annual Bulletin of Housing and Building Statistics for Europe, 1965. Journal of the Royal Statistical Society. Series A (General), 131(1):117, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344108.

Perkins:1968:BRB

[1121] R. J. Perkins. Book review: Economic Organization and Results of the Main Types of Welsh Farm: 1960–1965 and 1964–1965, by Michael B. Jawetz. Journal of the Royal Statistical Society. Series A (General), 131(1):117– 118, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344109.

Anonymous:1968:CNa

[1122] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 131(1):119–125, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344110.

Lomnicki:1968:JDN

[1123] Z. A. Lomnicki. John Dudley North, C.B.E. Journal of the Royal Statistical Society. Series A (General), 131 (1):126–127, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344111.

Anonymous:1968:LAa

[1124] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 131(1):128–131, ???? 1968. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344112.

Anonymous:1968:BMa

[1125] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 131(1):i–x, ???? 1968. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344113.

Anonymous:1968:FMb

[1126] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 131(2):??, ???? 1968. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343839.

Thatcher:1968:DEE

[1127] A. R. Thatcher. The distribution of earnings of employees in Great Britain. Journal of the Royal Statistical Society. Series A (General), 131(2):133– 181, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343840.

Kendall:1968:FSS

[1128] M. G. Kendall. On the future of statistics — a second look. Journal of the Royal Statistical Society. Series A (General), 131(2):182–204, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343841.

Irwin:1968:GWD

[1129] J. O. Irwin. The generalized Waring distribution applied to accident theory. Journal of the Royal Statistical Society. Series A (General), 131(2):205– 225, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343842.

Joseph:1968:CMC

[1130] A. W. Joseph. A criticism of the Monte Carlo method as applied to mathematical computations. Journal of the Royal Statistical Society. Series A (General), 131(2):226–228, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343843.

Douglas:1968:CAA

[1131] J. W. B. Douglas, J. M. Ross, and H. R. Simpson. Corrigenda: The ability and attainment of short-sighted pupils. Journal of the Royal Statistical Society. Series A (General), 131(2):229, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343844. See [1061].

Ridler-Rowe:1968:BRB

[1132] C. J. Ridler-Rowe. Book review: A Graduate Course in Probability, by H. G. Tucker. Journal of the Royal Statistical Society. Series A (General), 131 (2):230–231, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343845.

Wilkinson:1968:BRBb

[1133] R. Wilkinson. Book review: Modern Elementary Statistics, by John E. Freund. Journal of the Royal Statistical Society. Series A (General), 131(2): 231–232, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343846.

Stone:1968:BRB

[1134] M. Stone. Book review: Mathematical Statistics: A Decision Theoretic Approach, by Thomas S. Ferguson. Journal of the Royal Statistical Society. Series A (General), 131(2):232, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343847.

Sprent:1968:BRB

[1135] P. Sprent. Book review: Laws and Chance Variations: Basic Concepts of Statistical Inference.: Vol. I: Elementary Introduction, by Erling Sverdrup; Laws and Chance Variations: Basic Concepts of Statistical Inference.: Vol. II: More Advanced Treatment, by Erling Sverdrup. Journal of the Royal Statistical Society. Series A (General), 131(2):232–233, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343848.

Preece:1968:BRB

[1136] D. A. Preece. Book review: Patterns and Configurations in Finite Spaces, by S. Vajda; The Mathematics of Experimental Design; Incomplete Block Designs and Latin Squares, by S. Vajda.

Journal of the Royal Statistical Society. Series A (General), 131(2):233– 234, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343849.

Bassett:1968:BRB

[1137] E. E. Bassett. Book review: Handbook of the Poisson Distribution, by F. A. Haight. Journal of the Royal Statistical Society. Series A (General), 131 (2):234–235, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343850.

Thomas:1968:BRBa

[1138] J. J. Thomas. Book review: Statistical Techniques in Business and Economics, by R. D. Mason. Journal of the Royal Statistical Society. Series A (General), 131(2):235–236, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343851.

Campling:1968:BRB

[1139] G. E. G. Campling. Book review:

Elementary Statistics: Data Analysis for the Behavioral Sciences, by Paul A. Games; George R. Klare. Journal of the Royal Statistical Society. Series A (General), 131(2):236–237, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343852.

Winch:1968:BRB

[1140] Donald Winch. Book review: Essays on Economics and Society, by John Stuart Mill; Lord Robbins; J. M. Robson. Journal of the Royal Statistical Society. Series A (General), 131(2):237– 238, ???? 1968. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2343853.

Pollard:1968:BRB

[1141] Sidney Pollard. Book review: British Economic Growth 1688–1959: Trends and Structure, by Phyllis Deane; W. A. Cole. Journal of the Royal Statistical Society. Series A (General), 131 (2):238–239, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343854.

Smyth:1968:BRB

[1142] D. J. Smyth. Book review: Higgledy Piggledy Growth Again, by A. C. Rayner; J. M. D. Little. Journal of the Royal Statistical Society. Series A (General), 131(2):239–240, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343855.

Davenport:1968:BRB

[1143] M. Davenport. Book review: The Theory of Quantitative Economic Policy with Applications to Economic Growth and Stabilization, by Karl A. Fox; Jati K. Sengupta; Erik Thorbecke. Journal of the Royal Statistical Society. Series A (General), 131(2):240–243, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343856.

Williamson:1968:BRB

[1144] John Williamson. Book review: The Monetary Mechanism of International Adjustment, by Robert A. Mundell. Journal of the Royal Statistical Society. Series A (General), 131(2):243– 244, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343857.

Sargan:1968:BRB

[1145] J. D. Sargan. Book review: Stochastically Dependent Equations: An Introductory Text for Econometricians, by P. R. Fisk. Journal of the Royal Statistical Society. Series A (General), 131 (2):244–245, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343858.

Rutland:1968:BRB

[1146] E. H. Rutland. Book review: Principles and Recommendations for the 1970 Housing Censuses. Journal of the Royal Statistical Society. Series A (General), 131(2):245–246, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343859.

Finney:1968:BRBa

[1147] D. J. Finney. Book review: Statistics for Biologists, by R. C. Campbell. Journal of the Royal Statistical Society. Series A (General), 131(2):246– 247, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343860.

Thomas:1968:BRBb

[1148] W. J. Thomas. Book review: British Agriculture and the Balance of Payments: Agriculture’s Contribution to Import Saving, by D. Friedman. Journal of the Royal Statistical Society. Series A (General), 131(2):247, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343861.

Haley:1968:BRB

[1149] K. B. Haley. Book review: Linear and Non-Linear Programming in Industry, by N. Williams. Journal of the Royal Statistical Society. Series A (General), 131(2):248, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343862.

Beale:1968:BRB

[1150] E. M. L. Beale. Book review: Nonlinear Programming, by Hans Paul Künzi; Wilhelm Krelle; Werner Oettli; Frank Levin. Journal of the Royal Statistical Society. Series A (General), 131(2): 248–249, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343863.

Priestley:1968:BRB

[1151] M. B. Priestley. Book review: Statistical Dynamics of Automatic Control Systems, by J. Bená. Journal of the Royal Statistical Society. Series A (General), 131(2):249–250, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343864.

Kmietowicz:1968:BRB

[1152] Z. W. Kmietowicz. Book review: Introduction to Operations Research, by Frederick S. Hillier; Gerald I. Lieberman. Journal of the Royal Statistical Society. Series A (General), 131(2): 250–251, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343865.

Anonymous:1968:CNb

[1153] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 131(2):252–258, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343866.

Silverstone:1968:ACA

[1154] H. Silverstone. Alexander Craig Aitken, M.A., D.Sc., Ll.D., F.R.S.E., F.R.S. (1895–1967). Journal of the Royal Statistical Society. Series A (General), 131(2):259–261, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343867.

Anonymous:1968:LAb

[1155] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 131(2):262–264, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343868.

Anonymous:1968:BMb

[1156] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 131(2):i–vi, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343869.

Anonymous:1968:FMc

[1157] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 131(3):??, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343522.

Cox:1968:NSA

[1158] D. R. Cox. Notes on some aspects of regression analysis. Journal of the Royal Statistical Society. Series A (General), 131(3):265–279, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343523.

Ehrenberg:1968:ELR

[1159] A. S. C. Ehrenberg. The elements of lawlike relationships. Journal of the Royal Statistical Society. Series A (General), 131(3):280–302, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343524.

Nelder:1968:RMB

[1160] J. A. Nelder. Regression, modelbuilding and invariance. Journal of the Royal Statistical Society. Series A (General), 131(3):303–329, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343525.

Peare:1968:RES

[1161] C. W. D. Peare and H. Thomas. Regional economic statistics. Journal of the Royal Statistical Society. Series A (General), 131(3):330–339, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343526.

Hollingsworth:1968:IMS

[1162] T. H. Hollingsworth. Internal migration statistics from the Central Register for Scotland of the National Health Service. Journal of the Royal Statistical Society. Series A (General), 131 (3):340–383, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343527.

Harbury:1968:IUU

[1163] C. D. Harbury and R. Szreter. The influence upon university performance of the study of economics at school. Journal of the Royal Statistical Society. Series A (General), 131(3):384–409, ????

1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343528.

Barnett:1968:SNR

[1164] V. D. Barnett, R. L. Holder, and T. Lewis. Some new results on the association between students’ ages and their degree results. Journal of the Royal Statistical Society. Series A (General), 131(3):410–433, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343529.

Good:1968:CGS

[1165] I. J. Good and T. N. Gover. Corrigendum: The generalized serial test and the binary expansion of \sqrt{2}. Journal of the Royal Statistical Society. Series A (General), 131(3):434, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343530. See [954].

Moore:1968:BRB

[1166] P. G. Moore. Book review: An Introduction to Management Statistics, by International Tutor Machines Ltd. Journal of the Royal Statistical Society. Series A (General), 131(3):435– 436, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343531.

Constable:1968:BRB

[1167] R. L. Constable. Book review: Elementary Statistical Methods, by G. B. Wetherill. Journal of the Royal Statistical Society. Series A (General), 131 (3):436–437, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343532.

Herzberg:1968:BRB

[1168] Agnes M. Herzberg. Book review: Elementary Statistical Procedures, by C. I. Chase. Journal of the Royal Statistical Society. Series A (General), 131 (3):437, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343533.

Gilchrist:1968:BRB

[1169] W. G. Gilchrist. Book review: Introduction to Probability and Statistical Decision Theory, by G. Hadley. Journal of the Royal Statistical Society. Series A (General), 131(3):437–438, ???? 1968. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343534.

Lindley:1968:BRBa

[1170] D. V. Lindley. Book review: Statistical Analysis for Business Decisions, by William A. Spurr; Charles P. Bonini. Journal of the Royal Statistical Society. Series A (General), 131(3):438– 439, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343535.

Hiorns:1968:BRB

[1171] R. W. Hiorns. Book review: Teach Yourself Operational Research. Problems, Techniques and Exercises, by M. S. Makower; E. Williamson. Journal of the Royal Statistical Society. Series A (General), 131(3):439–440, ???? 1968. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343536.

Mallya:1968:BRB

[1172] R. V. Mallya. Book review: Problems of Stocks and Storage, by A. J. H. Morrell. Journal of the Royal Statistical Society. Series A (General), 131(3):440, ???? 1968. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343537.

Novick:1968:BRB

[1173] M. R. Novick. Book review: Models, Methods and Analytical Procedures in Educational Research, by J. E. Hill; A. Kerber. Journal of the Royal Statistical Society. Series A (General), 131 (3):440–441, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343538.

Vandome:1968:BRB

[1174] P. Vandome. Book review: Fortran Programs for Economists, by Lucy Joan Slater. Journal of the Royal Statistical Society. Series A (General), 131 (3):441–442, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343539.

Wallis:1968:BRB

[1175] Kenneth F. Wallis. Book review:

The Construction and Use of Economic Models, by A. R. Bergstrom. Journal of the Royal Statistical Society. Series A (General), 131(3):442–443, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343540.

Shaw:1968:BRB

[1176] G. K. Shaw. Book review: The Years of High Theory: Invention and Tradition in Economic Thought 1926–1939, by G. L. S. Shackle. Journal of the Royal Statistical Society. Series A (General), 131 (3):443–444, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343541.

Brothwell:1968:BRB

[1177] J. F. Brothwell. Book review: On the Nature of Profit, by G. L. S. Shackle. Journal of the Royal Statistical Society. Series A (General), 131(3):444, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343542.

Wilkinson:1968:BRBc

[1178] R. K. Wilkinson. Book review: Incomes Policy and the Individual, by Hilde Behrend; Harriet Lynth; Howard Thomas; Jean Davies. Journal of the Royal Statistical Society. Series A (General), 131(3):445–446, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343543.

McCormick:1968:BRBa

[1179] B. J. McCormick. Book review: Wage Drift: Origins, Measurement and Behaviour, by Judith Marquand. Journal of the Royal Statistical Society. Series A (General), 131(3):446–447, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343544.

Preston:1968:BRB

[1180] M. H. Preston. Book review: National Economic Planning: A Conference of the Universities-National Bureau Committee for Economic Research, by Max F. Millikan. Journal of the Royal Statistical Society. Series A (General), 131 (3):447–448, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343545.

Meek:1968:BRB

[1181] Ronald L. Meek. Book review: Papers on Capitalism, Development and Planning, by Maurice Dobb. Journal of the Royal Statistical Society. Series A (General), 131(3):448–449, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343546.

Oliver:1968:BRB

[1182] F. R. Oliver. Book review: The Quality of Consumer Instalment Credit, by Geoffrey H. Moore; Phillip A. Klein. Journal of the Royal Statistical Society. Series A (General), 131(3):449, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343547.

Peston:1968:BRB

[1183] Maurice Peston. Book review: Betterment Levy, by K. Buckley Edwards. Journal of the Royal Statistical Society. Series A (General), 131(3):449– 450, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343548.

Howe:1968:BRB

[1184] M. Howe. Book review: Technical Cooperation for the Development of Small Scale Industries, by U. N. Department of Economic and Social Affairs. Journal of the Royal Statistical Society. Series A (General), 131(3):450–451, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343549.

McCormick:1968:BRBb

[1185] B. J. McCormick. Book review: The Utilization of Educated Manpower in Industry: A Preliminary Report, by M. Blaug; M. Peston; A. Ziderman. Journal of the Royal Statistical Society. Series A (General), 131(3):451–452, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343550.

Hollingsworth:1968:BRB

[1186] T. H. Hollingsworth. Book review: Return Migration to Puerto Rico, by A. J. Hernández. Journal of the Royal Statistical Society. Series A (General), 131 (3):452–453, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343551.

Adelstein:1968:BRB

[1187] A. M. Adelstein. Book review: World Health Statistics Annual 1964: Vol. 1. Vital Statistics and Causes of Death, by World Health Organization. Journal of the Royal Statistical Society. Series A (General), 131(3):453–454, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343552.

Thomas:1968:BRBc

[1188] C. J. Thomas. Book review: World Population Conference, 1965. Volume II: Fertility, Family Planning Mortality, by United Nations Organization; World Population Conference, 1965. Volume III: Projections, Measurement of Population Trends, by United Nations Organization. Journal of the Royal Statistical Society. Series A (General), 131(3):454–455, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343553.

Anonymous:1968:CNc

[1189] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 131(3):456–458, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343554.

Cairncross:1968:PED

[1190] Alec Cairncross. Professor Ely Devons.

Journal of the Royal Statistical Society. Series A (General), 131(3):459– 460, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343555.

Anonymous:1968:LAc

[1191] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 131(3):461–462, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343556.

Anonymous:1968:BMc

[1192] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 131(3):i–v, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343557.

Anonymous:1968:VIb

[1193] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series A (General), 131(4):636– 639, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343717.

Anonymous:1968:FMd

[1194] Anonymous. Front matter. Journal of the Royal Statistical Society.

Series A (General), 131(4):??, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343718.

Yates:1968:TPS

[1195] F. Yates. Theory and practice in statistics. Journal of the Royal Statistical Society. Series A (General), 131(4):463– 477, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343719.

Yates:1968:IRR

[1196] F. Yates. Interim report of the R.S.S. Committee on the Teaching of Statistics in Schools. Journal of the Royal Statistical Society. Series A (General), 131(4):478–499, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343720.

Downton:1968:SLR

[1197] F. Downton. Statistics at A-level — a reappraisal. Journal of the Royal Statistical Society. Series A (General), 131 (4):500–529, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343721.

Leser:1968:SE

[1198] C. E. V. Leser. A survey of econometrics. Journal of the Royal Statistical Society. Series A (General), 131 (4):530–566, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343722.

Healy:1968:LS

[1199] M. J. R. Healy. The lengths of surnames. Journal of the Royal Statistical Society. Series A (General), 131 (4):567–568, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343723.

Pollard:1968:NAS

[1200] J. H. Pollard. A note on the age structures of learned societies. Journal of the Royal Statistical Society. Series A (General), 131(4):569–578, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343724.

Bartlett:1968:FNN

[1201] M. S. Bartlett. A further note on nearest neighbour models. Journal of the Royal Statistical Society. Series A (General), 131(4):579–580, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343725.

Reep:1968:SCA

[1202] C. Reep and B. Benjamin. Skill and chance in Association Football. Journal of the Royal Statistical Society. Series A (General), 131(4):581–585, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343726.

Anonymous:1968:ARC

[1203] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 131 (4):586–608, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343727.

Anonymous:1968:POH

[1204] Anonymous. Proceedings of the one hundred and thirty-fourth Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 131 (4):609, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343728.

Cormack:1968:BRB

[1205] R. M. Cormack. Book review: Statistics in Biology: Statistical Methods for Research in the Natural Sciences, by C. I. Bliss. Journal of the Royal Statistical Society. Series A (General), 131 (4):610–611, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343729.

Dugard:1968:BRB

[1206] P. Dugard. Book review: An Introduction to Statistical Mathematics, by A. M. Mathai. Journal of the Royal Statistical Society. Series A (General), 131(4):611, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343730.

Bather:1968:BRB

[1207] J. A. Bather. Book review: Optimization of Stochastic Systems: Topics in Discrete-Time Systems, by Masanao Aoki. Journal of the Royal Statistical Society. Series A (General), 131(4): 611–612, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343731.

Draper:1968:BRB

[1208] N. R. Draper. Book review: Prediction Analysis, by John R. Wolberg. Journal of the Royal Statistical Society. Series A (General), 131(4):612–613, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343732.

David:1968:BRB

[1209] F. N. David. Book review: Cours de Calcul des Probabilités, by G. Calot. Journal of the Royal Statistical Society. Series A (General), 131(4):613– 614, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343733.

Holgate:1968:BRB

[1210] P. Holgate. Book review: Probability, by J. R. Gray. Journal of the Royal Statistical Society. Series A (General), 131(4):614–615, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343734.

Loynes:1968:BRB

[1211] R. M. Loynes. Book review: Theory of Rank Tests, by Jaroslav Hájek; Zbynék Šidák. Journal of the Royal Statistical Society. Series A (General), 131(4): 615, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343735.

Maxwell:1968:BRB

[1212] A. E. Maxwell. Book review: Modern Factor Analysis, by Harry H. Harman. Journal of the Royal Statistical Society. Series A (General), 131(4): 615–616, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343736.

Lindley:1968:BRBb

[1213] D. V. Lindley. Book review: Formulae and Tables for Statistical Work, by C. R. Rao; S. K. Mitra; A. Matthai. Journal of the Royal Statistical Society. Series A (General), 131(4):616–617, ????

1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343737.

Healy:1968:BRBa

[1214] M. J. R. Healy. Book review: Multichannel Time Series Analysis with Digital Computer Programs, by E. A. Robinson. Journal of the Royal Statistical Society. Series A (General), 131 (4):617–618, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343738.

Nicholson:1968:BRBa

[1215] R. J. Nicholson. Book review: La Mesure de la Productivité, by André L. A. Vincent. Journal of the Royal Statistical Society. Series A (General), 131 (4):618–619, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343739.

Conway:1968:BRB

[1216] Freda Conway. Book review: Charting the British Economy, by Barbara Marlow. Journal of the Royal Statistical Society. Series A (General), 131 (4):619–620, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343740.

Coppock:1968:BRB

[1217] D. J. Coppock. Book review: Assessing Economic Performance, by Michael Lipton. Journal of the Royal Statistical Society. Series A (General), 131 (4):620–622, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343741.

Curtin:1968:BRB

[1218] T. R. C. Curtin. Book review: Classical Macroeconomics and the Case for Colonies, by R. N. Ghosh. Journal of the Royal Statistical Society. Series A (General), 131(4):622–623, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343742.

Finney:1968:BRBb

[1219] D. J. Finney. Book review: Quantitative Research in Human Biology and Medicine, by Sigismund Peller. Journal of the Royal Statistical Society. Series A (General), 131(4):623–624, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343743.

Barr:1968:BRB

[1220] A. Barr. Book review: Custom and Practice in Medical Care: A Comparative Study of Two Hospitals in Arbroath, Scotland, U.K. and Waterville, Maine, U.S.A, by J. Simpson; A. Mair; R. G. Thomas; H. N. Willard; H. J. Bakst. Journal of the Royal Statistical Society. Series A (General), 131(4): 624, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343744.

Healy:1968:BRBb

[1221] M. J. R. Healy. Book review: Problems and Progress in Medical Care, by G. McLachlan. Journal of the Royal Statistical Society. Series A (General), 131(4):625, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343745.

Nicholson:1968:BRBb

[1222] Michael Nicholson. Book review: Issues in Defense Economics, by R. N. McKean. Journal of the Royal Statistical Society. Series A (General), 131 (4):625–626, ???? 1968. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343746.

Saul:1968:BRB

[1223] S. B. Saul. Book review: The Technical Development of Modern Aviation, by Ronald Miller; David Sawers. Journal of the Royal Statistical Society. Series A (General), 131(4):626–627, ???? 1968. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343747.

Anonymous:1968:AR

[1224] Anonymous. Addendum: Reviews. Journal of the Royal Statistical Society. Series A (General), 131(4):627, ???? 1968. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343748.

Anonymous:1968:CNd

[1225] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 131(4):628–632, ???? 1968. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343749.

Anonymous:1968:LAd

[1226] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 131(4):633–635, ???? 1968. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343750.

Anonymous:1968:BMd

[1227] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 131(4):i–iv, ???? 1968. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343751.

Anonymous:1969:VIa

[1228] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series A (General), 132(1):??, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343752.

Anonymous:1969:FMa

[1229] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 132(1):??, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343753.

Benjamin:1969:STP

[1230] B. Benjamin. Statistics in Town planning. Journal of the Royal Statistical Society. Series A (General), 132(1):1– 28, ???? 1969. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343754.

Herzberg:1969:RWD

[1231] Agnes M. Herzberg and D. R. Cox.

Recent work on the design of experiments: A bibliography and a review. Journal of the Royal Statistical Society. Series A (General), 132(1):29–67, ???? 1969. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343755. See correction

Hawkes:1969:AAE

[1232] A. G. Hawkes. An approach to the analysis of electoral swing. Journal of the Royal Statistical Society. Series A (General), 132(1):68–79, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343756.

Pollard:1969:CTD

[1233] J. H. Pollard. Continuous-time and discrete-time models of population growth. Journal of the Royal Statistical Society. Series A (General), 132(1):80– 88, ???? 1969. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2343757.

Thomas:1969:SPM

[1234] V. J. Thomas. A stochastic population model related to human populations. Journal of the Royal Statistical Society. Series A (General), 132(1):89–104, ???? 1969. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2343758.

Silvey:1969:BRB

[1235] S. D. Silvey. Book review: Collected Papers. Contributions to Mathematical Statistics, by S. S. Wilks; T. W. Anderson. Journal of the Royal Statistical Society. Series A (General), 132 (1):105–106, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343759.

Kingman:1969:BRB

[1236] J. F. C. Kingman. Book review: An Introduction to Probability Theory, by P. A. P. Moran. Journal of the Royal Statistical Society. Series A (General), 132(1):106, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343760.

White:1969:BRB

[1237] D. J. White. Book review: Bayesian Decision Problems and Markov Chains, by J. J. Martin. Journal of the Royal Statistical Society. Series A (General), 132(1):106–107, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343761.

Lindley:1969:BRBa

[1238] D. V. Lindley. Book review: Statistical Analysis for Business Decisions, by William S. Peters; George W. Summers. Journal of the Royal Statistical Society. Series A (General), 132(1): 107–108, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343762.

Shutler:1969:BRBa

[1239] M. F. Shutler. Book review: Concepts of Operational Research, by Patrick Rivett. Journal of the Royal Statistical Society. Series A (General), 132 (1):108–109, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343763.

Vajda:1969:BRBa

[1240] S. Vajda. Book review: Mathematical Programming in Practice, by E. M. L. Beale. Journal of the Royal Statistical Society. Series A (General), 132 (1):109, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343764.

Wishart:1969:BRBa

[1241] D. M. G. Wishart. Book review: Introduction to the Mathematical Theory of Control Processes. Volume 1 — Linear Equations and Quadratic Criteria, by R. E. Bellman. Journal of the Royal Statistical Society. Series A (General), 132(1):109–110, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343765.

Wishart:1969:BRBb

[1242] D. M. G. Wishart. Book review: Foundations of Optimal Control Theory, by E. B. Lee; L. Markus. Journal of the Royal Statistical Society. Series A (General), 132(1):110, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343766.

Scott:1969:BRB

[1243] J. F. Scott. Book review: Forest Biometrics, by Michail Prodan; Sabine H. Gardiner. Journal of the Royal Statistical Society. Series A (General), 132 (1):110–111, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343767.

Turvey:1969:BRB

[1244] Ralph Turvey. Book review: Economics of Scale and Technological Change in Thermal Power Generation, by Malcolm Galatin. Journal of the Royal Statistical Society. Series A (General), 132(1):111, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343768.

Buckland:1969:BRBa

[1245] W. R. Buckland. Book review:

Forecasting on a Scientific Basis– Proceedings of an International Summer Institute in Curia, Portugal. Journal of the Royal Statistical Society. Series A (General), 132(1):111–113, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343769.

Yannoulis:1969:BRB

[1246] Y. Yannoulis. Book review: Étude Comparée des Tableaux d’Entrées et de Sorties des Communautés Européennes. Journal of the Royal Statistical Society. Series A (General), 132 (1):113–114, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343770.

Oliver:1969:BRBa

[1247] F. R. Oliver. Book review: Public Policies and Private Investment, by John F. Helliwell. Journal of the Royal Statistical Society. Series A (General), 132 (1):115, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343771.

Taylor:1969:BRBa

[1248] Wallis Taylor. Book review: Proceedings of the World Population Conference, Belgrade, 30 August–10 September 1965. Volume 1. Summary Report. Journal of the Royal Statistical Society. Series A (General), 132(1):116– 117, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343772.

Brass:1969:BRB

[1249] W. Brass. Book review: Corrected Age Data of the 1931 Indian Census, by S. N. Agarwala. Journal of the Royal Statistical Society. Series A (General), 132 (1):117, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343773.

Osborn:1969:BRBa

[1250] J. F. Osborn. Book review: Health and Vital Statistics, by B. Benjamin. Journal of the Royal Statistical Society. Series A (General), 132(1):118, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343774.

Spicer:1969:BRB

[1251] C. C. Spicer. Book review: Manual of the International Statistical Classification of Diseases Injuries, and Causes of Death–Volume 1, by U. N. World Health Organization. Journal of the Royal Statistical Society. Series A (General), 132(1):118–119, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343775.

Anonymous:1969:CNa

[1252] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 132(1):120–121, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343776.

Anonymous:1969:RBS

[1253] Anonymous. Recommended basic statistical sources for community use. Journal of the Royal Statistical Society. Series A (General), 132(1):122– 128, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343777.

Benjamin:1969:FM

[1254] Bernard Benjamin. F. A. A. Menzler.

Journal of the Royal Statistical Society. Series A (General), 132(1):129– 130, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343778.

Anonymous:1969:LAa

[1255] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 132(1):131–132, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343779.

Anonymous:1969:BMa

[1256] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 132(1):i–viii, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343780.

Anonymous:1969:FMb

[1257] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 132(2):??, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343781.

Coen:1969:LRE

[1258] P. J. Coen, E. D. Gomme, and M. G. Kendall. Lagged relationships in economic forecasting. Journal of the Royal Statistical Society. Series A (General), 132(2):133–163, ???? 1969. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2343782. See comments

[1345].

Horsnell:1969:TCB

[1259] Gareth Horsnell. A theory of consumer behaviour derived from repeat paired preference testing. Journal of the Royal Statistical Society. Series A (General), 132(2):164–193, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343783.

Glennerster:1969:PR

[1260] Howard Glennerster. The Plowden research. Journal of the Royal Statistical Society. Series A (General), 132 (2):194–204, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343784.

Phillips:1969:SSP

[1261] M. J. Phillips. A survey of sampling procedures for continuous production. Journal of the Royal Statistical Society. Series A (General), 132(2):205– 228, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343785.

Subramanian:1969:SWO

[1262] S. Subramanian. Statistical wisdom– old or new? Journal of the Royal Statistical Society. Series A (General), 132 (2):229–234, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343786.

Armitage:1969:RST

[1263] P. Armitage, C. K. McPherson, and B. C. Rowe. Repeated significance tests on accumulating data. Journal of the Royal Statistical Society. Series A (General), 132(2):235–244, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343787.

Wabe:1969:LFP

[1264] J. S. Wabe. Labour force participation rates in the London metropolitan region. Journal of the Royal Statistical Society. Series A (General), 132 (2):245–264, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343788.

Heien:1969:IPL

[1265] Dale M. Heien. Income and price lags in consumer-demand analysis. Journal of the Royal Statistical Society. Series A (General), 132(2):265–271, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343789.

Maxwell:1969:BRB

[1266] A. E. Maxwell. Book review: Simplified Statistics for Education and Psychology: A Work-Text with Feedback, by B. L. Turney; G. P. Robb. Journal of the Royal Statistical Society. Series A (General), 132(2):272, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343790.

Gray:1969:BRB

[1267] P. G. Gray. Book review: Survey Sampling, by Leslie Kish. Journal of the Royal Statistical Society. Series A (General), 132(2):272–274, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343791.

Williams:1969:BRB

[1268] Alan Williams. Book review: Public Sector Economics, by A. R. Prest. Journal of the Royal Statistical Society. Series A (General), 132(2):274– 275, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343792.

Reddaway:1969:BRB

[1269] W. B. Reddaway. Book review: The Wealth of the Nation: The National Balance Sheet of the U.K., 1957–1961, by Jack Revell. Journal of the Royal Statistical Society. Series A (General), 132(2):276, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343793.

Taylor:1969:BRBb

[1270] C. T. Taylor. Book review: R & D and Small-Company Growth: A Statistical Review and Company Case Studies, by Wm. J. J. Smith; D. Creamer. Journal of the Royal Statistical Society. Series A (General), 132(2):276–277, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343794.

Cox:1969:BRBa

[1271] P. R. Cox. Book review: Introduction to Demography, by Mortimer Spiegelman. Journal of the Royal Statistical Society. Series A (General), 132(2): 278–279, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343795.

Cox:1969:BRBb

[1272] P. R. Cox. Book review: Demographic Analysis, by Bernard Benjamin. Journal of the Royal Statistical Society. Series A (General), 132(2):279–280, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343796.

Smith:1969:BRB

[1273] T. E. Smith. Book review: World Views of Population Problems, by E. Szabady. Journal of the Royal Statistical Society. Series A (General), 132 (2):280–281, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343797.

Barrett:1969:BRB

[1274] J. C. Barrett. Book review: World Population: An Analysis of Vital Data, by Nathan Keyfitz; Wilhelm Flieger. Journal of the Royal Statistical Society. Series A (General), 132(2):281, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343798.

Walker:1969:BRB

[1275] Monica A. Walker. Book review:

Accidents and Homicide, by Albert P. Iskrant; Paul V. Joliet. Journal of the Royal Statistical Society. Series A (General), 132(2):281–282, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343799.

Vaizey:1969:BRB

[1276] John Vaizey. Book review: Educational Finance, by Alan Peacock; Howard Glennerster; Robert Lavers. Journal of the Royal Statistical Society. Series A (General), 132(2):282–283, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343800.

Stuart:1969:BRB

[1277] Alan Stuart. Book review: Normal Centroids, Medians and Scores for Ordinal Data, by F. N. David; D. E. Barton; S. Ganeshalingam; H. L. Harter; P. J. Kim; M. Merrigton; D. Walley. Journal of the Royal Statistical Society. Series A (General), 132(2):283, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343801.

Anonymous:1969:CNb

[1278] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 132(2):284–286, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343802.

Anonymous:1969:RPa

[1279] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 132(2):287–289, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343803.

Anonymous:1969:LAb

[1280] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 132(2):290–291, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343804.

Anonymous:1969:BMb

[1281] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 132(2):i–vi, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343805.

Anonymous:1969:FMc

[1282] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 132(3):??, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344114.

Wishart:1969:SCT

[1283] David M. G. Wishart. A survey of control theory. Journal of the Royal Statistical Society. Series A (General), 132 (3):293–319, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344115.

Whittle:1969:VSC

[1284] P. Whittle. A view of stochastic control theory. Journal of the Royal Statistical Society. Series A (General), 132 (3):320–334, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344116.

Bather:1969:DMS

[1285] J. A. Bather. Diffusion models in stochastic control theory. Journal of the Royal Statistical Society. Series A (General), 132(3):335–352, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344117.

Kramer:1969:SMD

[1286] Morton Kramer. Statistics of mental disorders in the United States: Current status, some urgent needs and suggested solutions. Journal of the Royal Statistical Society. Series A (General), 132(3):353–407, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2344118.

Smith:1969:BBE

[1287] C. Selby Smith. Benefits to British employers from post-secondary education. Journal of the Royal Statistical Society. Series A (General), 132(3):408– 417, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344119.

Rao:1969:TDC

[1288] V. M. Rao. Two decompositions of concentration ratio. Journal of the Royal Statistical Society. Series A (General), 132(3):418–425, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344120.

Smith:1969:NRE

[1289] T. M. F. Smith. A note on ratio estimates in multi-stage sampling. Journal of the Royal Statistical Society. Series A (General), 132(3):426–430, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344121.

Kpedekpo:1969:WLT

[1290] G. M. K. Kpedekpo. On working life tables in Ghana with particular reference to the female working population. Journal of the Royal Statistical Society. Series A (General), 132(3):431– 441, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344122.

Herzberg:1969:BRB

[1291] Agnes M. Herzberg. Book review: Statistical Methods, by G. W. Snedecor; W. G. Cochran. Journal of the Royal Statistical Society. Series A (General), 132(3):442, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344123.

Rosenhead:1969:BRB

[1292] Jonathan Rosenhead. Book review:

Basic Operational Research, by P. G. Moore. Journal of the Royal Statistical Society. Series A (General), 132 (3):443, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344124.

Lewis:1969:BRB

[1293] J. A. Lewis. Book review: Handbook of Methods of Applied Statistics, by I. M. Chakravarti; R. G. Laha; J. Roy. Journal of the Royal Statistical Society. Series A (General), 132(3):443–444, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344125.

Mitchell:1969:BRB

[1294] Ann F. S. Mitchell. Book review:

Statistics for Mathematicians — an Introduction, by D. J. Finney. Journal of the Royal Statistical Society. Series A (General), 132(3):444–445, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344126.

Lindley:1969:BRBb

[1295] D. V. Lindley. Book review: Statistical Theory, by B. W. Lindgren. Journal of the Royal Statistical Society. Series A (General), 132(3):445, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344127.

Dunsmore:1969:BRB

[1296] I. R. Dunsmore. Book review: An Introduction to Bayesian Statistical Decision Processes, by Bruce W. Morgan. Journal of the Royal Statistical Society. Series A (General), 132(3):445– 446, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344128.

Andrews:1969:BRB

[1297] David F. Andrews and M. Stone. Book review: The Structure of Inference, by D. A. S. Fraser. Journal of the Royal Statistical Society. Series A (General), 132(3):446–449, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344129.

Mardia:1969:BRB

[1298] K. V. Mardia. Book review: Handbook of Nonparametric Statistics III: Analysis of Variance, by John E. Walsh. Journal of the Royal Statistical Society. Series A (General), 132(3):450, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344130.

Vajda:1969:BRBb

[1299] S. Vajda. Book review: Einführung in die Unternehmensforschung, by R. Henn; H. P. Künzi. Journal of the Royal Statistical Society. Series A (General), 132(3):450, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344131.

Brillinger:1969:BRB

[1300] David R. Brillinger. Book review: Multivariate Statistical Methods, by D. F.

Morrison. Journal of the Royal Statistical Society. Series A (General), 132 (3):450–451, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344132.

Brown:1969:BRB

[1301] S. Brown. Book review: Statistics in the Computer Age, by J. M. Craddock. Journal of the Royal Statistical Society. Series A (General), 132(3):451– 452, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344133.

Buckland:1969:BRBb

[1302] W. R. Buckland. Book review: Handbook of Tables for Probability and Statistics, by William H. Beyer. Journal of the Royal Statistical Society. Series A (General), 132(3):452, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344134.

Tippett:1969:BRB

[1303] L. H. C. Tippett. Book review: Random Times for Activity Sampling, by I. D. Hill; Peter A. Hill. Journal of the Royal Statistical Society. Series A (General), 132(3):453, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344135.

Nobay:1969:BRB

[1304] A. R. Nobay. Book review: Economic Statistics and Economic Problems, by R. Nicholson. Journal of the Royal Statistical Society. Series A (General), 132 (3):453–454, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344136.

Thomas:1969:BRBa

[1305] C. J. Thomas. Book review: Population, Labor Force and Long Swings in Economic Growth: The American Experience, by R. A. Easterlin. Journal of the Royal Statistical Society. Series A (General), 132(3):454–455, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344137.

Phillips:1969:BRB

[1306] Celia Phillips. Book review: Education and Manpower: Theoretical Models and Empirical Applications, by Tore Thonstad. Journal of the Royal Statistical Society. Series A (General), 132 (3):455–456, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344138.

Wilkinson:1969:BRB

[1307] R. K. Wilkinson. Book review: Pay and Profits, by E. H. Phelps-Brown. Journal of the Royal Statistical Society. Series A (General), 132(3):456–457, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344139.

Thatcher:1969:BRB

[1308] A. R. Thatcher. Book review: The Structure of Earnings, by Harold Lydall. Journal of the Royal Statistical Society. Series A (General), 132(3): 457–458, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344140.

Lowe:1969:BRB

[1309] E. A. Lowe. Book review: Planned Organizational Change, by Garth N.

Jones. Journal of the Royal Statistical Society. Series A (General), 132(3): 458–459, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344141.

Anonymous:1969:CNc

[1310] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 132(3):460–463, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344142.

Anonymous:1969:RPb

[1311] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 132(3):464–469, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344143.

Anonymous:1969:LAc

[1312] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 132(3):470–471, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344144.

Anonymous:1969:BMc

[1313] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 132(3):i–iv, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344145.

Anonymous:1969:VIb

[1314] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series A (General), 132(4):615– 617, ???? 1969. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2344146.

Anonymous:1969:FMd

[1315] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 132(4):??, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344147.

Cockfield:1969:RE

[1316] F. A. Cockfield. The reality of experience. Journal of the Royal Statistical Society. Series A (General), 132 (4):473–483, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344148.

Moran:1969:SMP

[1317] P. A. P. Moran. Statistical methods in psychiatric research. Journal of the Royal Statistical Society. Series A (General), 132(4):484–525, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344149.

Rees:1969:INH

[1318] M. S. Rees. The inflation of National Health Service registers of patients and its effect on the remuneration of general practitioners. Journal of the Royal Statistical Society. Series A (General), 132(4):526–542, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2344150.

Downton:1969:NBS

[1319] F. Downton. A note on betting strategies in compound games. Journal of the Royal Statistical Society. Series A (General), 132(4):543–547, ????

1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344151.

Mantel:1969:PPO

[1320] Nathan Mantel. Permutational probabilities for opening the jackpot. Journal of the Royal Statistical Society. Series A (General), 132(4):548–550, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344152.

Anonymous:1969:POH

[1321] Anonymous. Proceedings of the one hundred and thirty-fifth Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 132 (4):551, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344153.

Anonymous:1969:ARC

[1322] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 132 (4):552–575, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344154.

Garside:1969:BRB

[1323] M. J. Garside. Book review: Methods of Multivariate Analysis, by Keith Hope. Journal of the Royal Statistical Society. Series A (General), 132(4): 576–577, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344155.

Ilersic:1969:BRB

[1324] A. R. Ilersic. Book review: Mathematics for Introductory Statistics: A Programmed Review, by Andrew R. Baggaley. Journal of the Royal Statistical Society. Series A (General), 132(4):577, ???? 1969. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2344156.

Thomas:1969:BRBb

[1325] J. J. Thomas. Book review: Introduction to Business Statistics, by G. Hadley. Journal of the Royal Statistical Society. Series A (General), 132 (4):578–579, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344157.

Oliver:1969:BRBb

[1326] F. R. Oliver. Book review: Elementary Matrices for Economics, by Maurice H. Peston. Journal of the Royal Statistical Society. Series A (General), 132(4):579, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344158.

Shutler:1969:BRBb

[1327] M. F. Shutler. Book review: Operations Research for Management Decisions, by Samuel B. Richmond. Journal of the Royal Statistical Society. Series A (General), 132(4):579–580, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344159.

Wishart:1969:BRBc

[1328] D. M. G. Wishart. Book review: Selected Statistical Papers. Vols. 1 & 2. Journal of the Royal Statistical Society. Series A (General), 132(4):580– 581, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344160.

Mills:1969:BRB

[1329] G. Mills. Book review: Information, Expectations and Inventory Fluctuations: A Study of Materials Stock on Hand and on Order, by Ruth P. Mack. Journal of the Royal Statistical Society. Series A (General), 132(4):581, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344161.

Pluck:1969:BRB

[1330] R. A. Pluck. Book review: Toward a Social Report. Journal of the Royal Statistical Society. Series A (General), 132 (4):581–584, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344162.

Worswick:1969:BRB

[1331] G. D. N. Worswick. Book review: The Service Economy, by Victor R. Fuchs. Journal of the Royal Statistical Society. Series A (General), 132(4):584– 585, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344163.

Nicholson:1969:BRB

[1332] R. J. Nicholson. Book review: Six Papers on the Size Distribution of Wealth and Income, by Lee Soltow. Journal of the Royal Statistical Society. Series A (General), 132(4):585–586, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344164.

Osborn:1969:BRBb

[1333] W. T. Osborn. Book review: Economic Aspects of Pensions: A Summary Report 1968, by R. F. Murray. Journal of the Royal Statistical Society. Series A (General), 132(4):586–587, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344165.

Lawson:1969:BRB

[1334] G. H. Lawson. Book review: Investment Economics, by J. L. Carr. Journal of the Royal Statistical Society. Series A (General), 132(4):587–588, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344166.

Darlington:1969:BRB

[1335] D. Darlington. Book review: Measurement in Medicine: The Interpretation of Numerical Data, by P. D. Oldham. Journal of the Royal Statistical Society. Series A (General), 132(4):588– 589, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344167.

Thapar:1969:BRBa

[1336] Savitri Thapar. Book review: Infant, Perinatal, Maternal and Childhood Mortality in the United States, by Sam Shapiro; Edward R. Schlesinger; Robert E. L. Nesbitt, Jr. Journal of the Royal Statistical Society. Series A (General), 132(4):589–590, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344168.

Cox:1969:BRBc

[1337] P. R. Cox. Book review: Size and Sex Composition of Population in India, 1901–1961, by P. B. Desai. Journal of the Royal Statistical Society. Series A (General), 132(4):590–591, ????

1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344169.

Thapar:1969:BRBb

[1338] Savitri Thapar. Book review: Principles of Demography, by Donald J. Bogue. Journal of the Royal Statistical Society. Series A (General), 132 (4):591–592, ???? 1969. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2344170.

Anonymous:1969:CNd

[1339] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 132(4):593–595, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344171.

Campion:1969:SR

[1340] H. Campion. Stuart A. Rice. Journal of the Royal Statistical Society. Series A (General), 132(4):596, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344172.

Stringer:1969:WNJ

[1341] John Stringer. William Neil Jessop.

Journal of the Royal Statistical Society. Series A (General), 132(4):596– 597, ???? 1969. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2344173.

Anonymous:1969:RPc

[1342] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 132(4):598–611, ???? 1969. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2344174.

Anonymous:1969:LAd

[1343] Anonymous. Library accessions. Journal of the Royal Statistical Society. Series A (General), 132(4):612–614, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344175.

Anonymous:1969:BMd

[1344] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 132(4):i–vi, ???? 1969. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2344176.

Box:1971:SCP

[1345] George E. P. Box and Paul Newbold. Some comments on a paper of Coen, Gomme and Kendall. Journal of the Royal Statistical Society. Series A (General), 134(2):229–240, ???? 1971. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343873. See [1258].

Herzberg:1978:CRW

[1346] A. M. Herzberg and D. R. Cox. Correction: Recent work on the design of experiments: A bibliography and a review. Journal of the Royal Statistical Society. Series A (General), 141 (2):286, ???? 1978. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2344479. See

[1231].